### v1

# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

Abstract
Context. To date, three nuclear transients have been associated with high-energy neutrino events. These transients are generally thought to be powered by tidal disruptions of stars (TDEs) by massive black holes. However, AT2019aalc, hosted in a Seyfert-1 galaxy, was not yet classified due to a lack of multiwavelength observations. Interestingly, the source has re-brightened 4 years after its discovery.
Aims. We aim to classify the transient and explain the mechanism responsible for its second optical flare.
Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023.
Results. The observations revealed a uniquely bright UV counterpart and multiple X-ray flares during the second optical flaring episode of the transient. The second flare, similarly to the first one, is also accompanied by IR dust echo emission. A long-term radio flare is found with an inverted spectrum. Optical spectroscopic observations reveal the presence of Bowen Fluorescence lines and strong high-ionization coronal lines indicating an extreme level of ionization in the system.
Conclusions. The results suggest that the transient can be classified as a Bowen Fluorescence Flare (BFF), a relatively new sub-class of flaring active galactic nuclei (AGN). AT2019aalc can be also classified as an extreme coronal line emitter (ECLE). We found that, in addition to AT2019aalc, another BFF AT2021loi is spatially coincident with a high-energy neutrino event. The multi-wavelength properties of these transients suggest a possible connection between ECLEs, BFFs and TDEs in AGN.
Key Words.:
galaxies: active – galaxies: Seyfert – quasars: emission lines – individual: AT2019aalc – Neutrinos1 Introduction
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. Despite most of these transient sources have been discovered by optical telescopes, the origin of their optical/UV emission is poorly understood. One model is colliding stellar debris stream shocks (Piran et al. 2015). Another scenario is reprocessed soft X-ray/far UV emission (originating from the debris accretion onto the black hole) by an extended elliptical disk, or an envelope, or an outflow surrounding the central debris disk (for a recent review see Bu et al. 2022). TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying optical emission. Several TDEs were detected in X-rays with a very soft X-ray spectrum typically explained by late-time accretion disk formation (Saxton et al. 2020). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (Horesh et al. 2021; Cendes et al. 2023), or only in instances by newly launched relativistic on-axis jets (Berger et al. 2012; Brown et al. 2015; Cenko et al. 2012; De Colle & Lu 2020; Yao et al. 2023a). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. As the photosphere shrinks over time, the flux of the ionized lines such as He ii tend to increase while the He ii/H ratio increases (e.g., Hung et al. 2017; Charalampopoulos et al. 2022a) and, in a few cases, Bowen Fluorescence (BF; Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al. 2019; Charalampopoulos et al. 2022a).
To date, roughly promising candidate TDEs are known. However, since most of the above-mentioned observed properties do not exclusively characterize TDEs, spectroscopic coverage is required to distinguish TDEs from ‘TDE impostors’ such as conventionally accreting active galactic nuclei (AGN), changing-look AGN (CLAGN), supernovae (SNe) or even stellar collisions (Zabludoff et al. 2021).
TDEs generally result in one optical flare, however, in a few cases a re-brightening episode (or episodes) was detected even years after the initial flare. These might be explained in the following way. Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in X-ray re-brightening explained by renewed accretion. The first systematically identified repeating partial TDE, AT2020vdq, has been classified after simultaneous multi-wavelength monitoring (Somalwar et al. 2023). Only a few other partial TDEs or candidates are known to date, these are ASASSN 14ko (Payne et al. 2021); eRASSt J045650.3-203750 (Liu et al. 2023); RX J133157.6-324319.7 (Hampel et al. 2022); AT2018fyk (Wevers et al. 2019a, 2023); AT2019avd (Chen et al. 2022) and AT2022dbl (Lin et al. 2024).
Based on the observed properties of TDE host galaxies, French et al. (2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al. 2023b). Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al. 2019; Ryu et al. 2024). The multi-wavelength characteristics of such a TDE should clearly differ from the ones taken place in quiescent galaxies, as the AGN is surrounded by a pre-existing accretion disk. Due to the extreme difference in density between the two, neither the star nor the accretion disk are affected after the first passage through the disk. However, the confined debris stream has such a low density after the star is tidally disturbed that when it interacts with the disk the collision between the stream and the disk can potentially dissipate much of the kinetic energy possessed by the stream and the disk gas near the impact point, and it can be radiated away (Chan et al. 2019). These transients are challenging to distinguish from AGN variability. Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a Narrow-Line Seyfert-1 (NLSy1) galaxy (Blanchard et al. 2017) or the multiple soft X-ray flares in IC 3599 (Campana et al. 2015) and GSN 069 (Shu et al. 2018)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al. 2021). In addition to TDEs, flaring NLSy1s, CLAGN and different types of SNe were classified as typical nuclear transients. Recently, Trakhtenbrot et al. (2019) identified a new class of flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray loud transient AT2019avd (Trakhtenbrot et al. 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (Makrygianni et al. 2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al. (2019). Those show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission. These AGN are characterized by significantly and longer-term intensified accretion and might be powered by a binary TDE or an interaction between an outflow and the broad-line region (BLR) or binary SMBHs (Trakhtenbrot et al. 2019). Notably, these transients take place in different types of AGN; AT2017bgt and OGLE17aaj in narrow-line AGN, AT2021loi in a broad-line AGN, F01004-2237 in an ultra-luminous infrared galaxy (ULIRG) while AT2019avd occurred in an earlier quiescent galaxy. Hereafter, we will refer to this new class of AGN flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al. (2023). Using the Zwicky Transient Facility (ZTF, Masci et al. 2019) Public Survey data, Dgany et al. (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Other candidates, namely VT J1548 (Somalwar et al. 2022) and AT2022fpx, (Koljonen et al. 2024) were found recently.
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al. 2023). The TDE AT2019dsg is likely associated with the IceCube neutrino, IC191001A (Stein et al. 2021) and a further promising association, between the candidate TDE AT2019fdr and the IceCube neutrino IC200530A (Reusch et al. 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were found to have unusually luminous mid-infrared emission, and a search for similar flares (van Velzen et al. 2024) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A (see Sect. 2. for explanation).
In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc especially on its significant optical re-brightening that has started years after its initial flare (Veres et al. 2023).
We conducted multi-wavelength observations (from radio to X-rays) during the re-brightening and compare it to the initial flare associated with the high-energy neutrino, with the aim to characterize the multi-wavelength properties of this unusual transient event and to classify it.
The paper is organized as follows. We explain the discovery of the transient event and give an insight about its host galaxy in Sect. 2, we summarize our observing campaign and the data reduction in Sect. 3, we present the results of the observations in Sect. 4 and discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al. 2012). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (Oke 1974).
2 The transient event: AT2019aalc
Studying the infrared properties of AT2019dsg and AT2019fdr have revealed a strong reverberation signal i.e. the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Dou et al. 2016; Lu et al. 2016; van Velzen et al. 2021b, 2024). In addition to the exceptionally high infrared luminosities (due to the infrared echo) of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al. 2021; Reusch et al. 2022). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al. 2024). This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC-191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al. 2013). Based on its Sloan Digital Sky Survey (SDSS; Abazajian et al. 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al. 2013; Liu et al. 2019). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST; Condon et al. 1998). Fig. 1 shows the Pan-STARRS (Chambers et al. 2016) i-band optical image of the host galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in . The host galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
In addition to the unusually significant infrared echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transient events: a soft X-ray spectrum and transient radio emission (van Velzen et al. 2024). Moreover, similar to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc, due to its high estimated SMBH mass (, van Velzen et al. 2024) and low redshift (, Ahn et al. 2012).
Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention or being directly monitored at the time when the neutrino was emitted. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al. 2022; van Velzen et al. 2024) suggesting a very unusual AGN case. The TDE-characteristic evolution of the blackbody light curve including a large and rapid optical flux increase is compatible with a TDE-related scenario. However, due to the host galaxy classification, extreme AGN variability provides an alternative explanation for the transient event.
Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al. 2023), illustrated by the long-term ZTF difference light curve shown in Fig. 2. Although double-peaked optical light curves could be a quasi-universal signature of TDEs around massive () black holes (Wevers et al. 2019b), such a long double peaked light curve with two comparable peak magnitudes would be very untypical for a regular TDE. Throughout this paper we will discuss the results of our observations taken around and after the peak of the second optical flare in the light of different TDE,- and AGN-related scenarios.
3 Observations and data reduction
3.1 Optical/UV
3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al. (2023). We corrected for extinction with adopting a galactic extinction of (Schlafly & Finkbeiner 2011). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band these were obtained in between March and May, 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
3.1.2 UVOT
The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972 and 19188, Proposers: Reusch, Veres& Sniegowska) on AT2019aalc from 2023-06-23 to 2024-05-27 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, Roming et al. 2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperature of arcseconds (which capture nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the host galaxy. To estimate the UV flux before the flare, we use GALEX (Martin et al. 2005) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al. 2016) to extract the flux. Over the 4 year baseline the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following van Velzen et al. (2021a), we combine the mean GALEX flux with the SDSS (York et al. 2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al. 2009) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. Finally, we corrected for galactic dust extinction for each filter.
The baseline flux in the UVM2 band is 17.2, implying a flux increase of 1.8 mag relative to the peak of the UVM2 light curve. This large flux increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.
3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al. 2003) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (Bertin et al. 2002). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes. Finally, we corrected for galactic dust extinction.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al. 2010), which is continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al. 2011), at (W2). It visits any point in the sky roughly every six months for about ten times. Using timewise (Necker & Mechbal 2024; Necker et al. 2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July .
3.3 X-ray
The Swift’s X-ray Telescope (XRT, Burrows et al. 2005) observations were obtained simultaneously with the UVOT observations. The observations were performed in photon counting (PC) mode. For each of the epochs, we performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way. First, cleaned event-files were produced using the xrtpipeline task. The exposure map, used to correct for the loss of flux caused by some of the CCD pixels not being used to collect data, was created with the xrtexpomap tool. The ximage tool allowed us to determine the source detection level and the source position. In case of non-detections, upper limits for the count rate were derived and converted to flux using the WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl tool and assuming the mean power-law index of the observations yielded in detections. When our target source was detected above the level, we ran the swxrtdas tool in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. In addition, Ancillary Response Files (ARFs) i.e. effective response files for the extracted spectra and Response Matrices Files (RMF), that are used to link the instrumental channel scale with the physical wavelength scale, were automatically created. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra and to create a readable spectrum for spectral fitting tools. The relevant ARF and RMF files were applied during this step. We rebinned each spectrum to have a minimum of counts per spectral bin. Channels below keV were ignored.
We fitted the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b) in order to be able to study the X-ray flux evolution of AT2019aalc. First, we have determined the needed Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al. 2016). Then, we loaded the source spectrum and converted the instrumental Channel scale to ’physical’ wavelength scale. The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as most of the AGN are dominated by a simple power-law and the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023). We performed renormalization and ignored bad data points. When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We than added a blackbody model component, to take into account the soft excess. The blackbody temperature and the normalization of the new component were also free to vary and we fitted again which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. This test runs a given number of simulations of datasets based on the two models provided (which are the simple power-law and the more complex power-law+blackbody models in our case) and calculates the likelihood ratio for the more complex model relative to the simpler one. After running simulations, the test resulted only in smaller likelihood differences than the observed one () clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit than a normal power-law fit to model the stacked spectrum. The source luminosity of each dataset was determined with adding the clumin component in front of the model components. We fixed the normalization in the additive models (power-law and blackbody) and performed the fit again. Finally, we derived the confidence level () error of the luminosity and the power-law index and, when it was relevant, the blackbody temperature. When fitting the individual datasets, a simple power-law model was enough to adequately describe the data and the soft excess becomes significant only in the stacked spectrum. Only one of the individual datasets is better fit with a blackbody model than with a power-law+blackbody or a simple power-law models.
3.4 Radio
3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al. 2004). Our very long baseline interferomety (VLBI) maps enable us to study the pc-scale radio structure of AT2019aalc. These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). Each of the eight intermediate-frequency channels (IFs) were divided into thirty-two kHz spectral channels, resulting in a total bandwidth of MHz. The whole observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The following antennas participated in the observations: Cm: Cambridge, UK; Da: Darnhall, UK; De: Defford, UK, Ef - Effelsberg, Germany; Hh - Hartebeesthoek, South Africa; Jb - Jodrell Bank Mk2, UK; Kn - Knockin, UK; MC - Medicina, Italy; Pi - Pickmere, UK; O8 - Onsala, Sweden; T6 - Tianma, China; Tr - Toruń, Poland; Ur - Urumqi, China; Wb - Westerbork, The Netherlands. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (Keimpema et al. 2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator. Earlier EVN observations of this calibrator source taken at GHz resolved two lobes, with a hint of a jet-like emission (Xiang et al. 2002). In addition, the following radio sources were observed as fringe-finders: J14301043, J15500527 and J17510939. The bright CSO J14072827 (a.k.a. OQ208) was observed as amplitude calibrator and in order to calibrate the Merlin antennas properly.
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures. We performed time flagging using the .uvflg ancilliary table. The amplitude calibration has been done using the measured system temperatures444Wb, T6, Ur and the e-MERLIN (Cm, Da, De, Kn, Pi) antennas had no or bad system temperature measurements, we used nominal values for amplitude calibration. and gain curves of each telescope contained in the .antab ancilliary table. Since our observations were conducted at below Ghz, we corrected for the dispersive ionospheric delays. Then, we corrected for the parallactic angle and performed global fringe-fitting for the calibrator sources. We imaged the phase-calibrator with the Difmap software (Shepherd et al. 1994), using the hybrid mapping technique with clean deconvolution (Högbom 1974). It included several iterations of cleaning and phase self-calibration before we performed an overall amplitude self-calibration to obtain the antenna-specific gain correction factors. We then applied these amplitude corrections for telescopes with deviation from the gain solutions in AIPS. Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct for the source structure since it contributes to the measured interferometric phases. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. As a final step, we exported the calibrated visibility file from AIPS to Difmap for imaging purposes described in the following.
After flagging the outlier data points, we imaged the target source using the hybrid mapping technique including the clean algorithm and phase-only self-calibration. Then, we obtained the antenna-specific gain correction factors. At the end of the procedure, several steps of amplitude and phase self-calibration were performed, starting from the total observation time, gradually for shorter and shorter solution time intervals.
A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data. This way we can describe the brightness distribution of the radio source for a quantitative analysis and to reveal any possible structure of it. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)555https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad; Sault et al. 1995). The data reduction procedure was conducted in accordance with standard protocol. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
Due to the limited ()-coverage and the low elevation of the source resulting in a very elongated beamshape, we were not able to perform self-calibration and cleaning. Fortunately, AT2019aalc was always dominating the flux within the resolution elements of our short observations. Therefore, we imaged the region around our source for each observation using an oversampling of the synthesised beam by a factor of at least 10. Under the assumption that AT2019aalc is always a point-source, we can take the highest pixel value in the image as our measured flux. Due to the uncertainties this method introduced we used a plain % flux error.
Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz | Flux Calibrator | Phase Calibrator |
28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
19.07.2023 | 210 | 6D | 2.1 | 1 | 1934-638 | 1548+056 |
09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
3.5 Spectroscopic observations
The Low Resolution Imaging Spectrograph (LRIS; Oke et al. 1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni), almost exactly at half time between the two optical flares. At that time the source was not significantly detected with ZTF as indicated by the forced photometry results. Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). These optical spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al. 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al. 2022) and PypeIt (Prochaska et al. 2020) for the DBSP. We normalized the flux of each spectrum and plotted it versus the rest-frame wavelength. Details of these observations of AT2019aalc are summarized in Table LABEL:table:sp. The obtained spectra cover a timescale of a year and allow us to study the spectroscopic evolution of the transient during its optical re-brightening. We describe the features of each spectra in Subsect. 4.5 and discuss the remarkable lines in Subsect. 5.1.
Date | Telescope | Instrument | Exp. | |
time | ||||
(d) | (s) | |||
Post-flare-1 | ||||
2021-07-06 | Keck-I | LRIS | 300 | |
Over-flare-2 | ||||
2023-06-15 | Keck-I | LRIS | 300 | |
2023-06-22 | LDT | DeVeny | 800 | |
2023-07-11 | Palomar Hale m | DBSP | 180 | |
2023-08-17 | Palomar Hale m | DBSP | 900 | |
2024-05-21 | Palomar Hale m | DBSP | 1200 |
To determine the type of AGN, we performed a stellar population synthesis fit of the host galaxy’s SDSS spectrum with (pPXF; Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates. Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al. (1989) extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al. (1989): a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10). In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .666https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
The SPS fit of the latest DBSP spectrum was conducted in a manner analogous to the one applied to the SDSS. However, additional kinematic components were required to account for the internal structure in order to achieve a satisfactory result. Each emission line with an excitation potential above is treated as a separate component. The [S ii] doublet is fitted as an additional component as well.
4 Results
4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 5. The transient was first detected by the ZTF in January with a g-band magnitude of mag. After the initial detection of the first optical flare in May (), AT2019aalc had brighten to a peak magnitude of () reached on 2019-06-18 on a timescale of roughly days. After the quickly and monotonically evolving flare, a slow decay has started, and the source was last detected with ZTF on 2021-02-18 with a magnitude of before its second flare. The re-brightening started mid-May , approximately years after the initial flare (Veres et al. 2023). The second flare has evolved very similarly to the original one. From the beginning of the second flare (the source was detected anew on 2023-05-11 at mag), a mag rise can be seen till the peak of the flare on 2023-07-13 with mag (), on a timescale of only days. Both flares evolved with constant color. However, the second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase lasted for around days. Both flares decay very slowly, on timescales of years. The first flare decayed with a power-law index of while the second has been decaying even slower with a power-law index of in -band. The decaying phases are not monotonous as bumps are tend to appear.
Prior to the discovery of AT2019aalc, the host galaxy of the transient has been monitored by several optical surveys. To study the optical variability of AT2019aalc prior to the first optical flare in , we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017) using ATLAS forced photometry server (Tonry et al. 2018; Smith et al. 2020) and the Catalina Real-Time Transient Survey (CRTS; Djorgovski et al. 2011) data. In addition, Palomar Transient Factory (PTF Law et al. 2009) magnitudes were extracted from the PTF Lightcurve Table777https://irsa.ipac.caltech.edu. No significant optical variability can be seen between April and the discovery of the transient in January by ZTF (see Fig. 4).
Both optical flares of AT2019aalc detected by ZTF rising with constant color and show reddening when decaying. The second flare has similar color evolution to the first one when evolving but with a color of around the peak.
The first optical flare was not monitored in UV. The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The UV flare evolved quickly. The U-band peak is coincident with the optical peak while the shorter wavelength bands (UM2, UW1 and UW2) peaked weeks later888We note that right after the -band peak, a short plateau phase started in optical and the UV emission in the bands UM2, UW1 and UW2 peak during this plateau. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in filter UW2. After the peak epoch, the UW2U color shows reddening which suggests increasing absorption or a drop of the blackbody temperature. Towards the end of the UVOT monitoring, the measured UM2 magnitudes decrease back to the level of the baseline GALEX-NUV magnitudes measured between and .
Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (Bianchi et al. 2011) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux. The amplitude of this flare is higher than typically seen for AGN. Only out of the AGN presented in the GALEX Time Domain Survey (Gezari et al. 2013) have larger amplitude NUV variability.
4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 5. The host galaxy of AT2019aalc is presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al. 2003) as 2MASS 152416660451191. The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 5 also shows the long-term light curve starting just before the first optical flare. There are clear signs of variability throughout the infrared light curve starting in February with the distinct first dust echo flare that peaks around six months after the first optical peak. The rise of the dust echo of the second flare becomes visible in the latest WISE data release. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al. 2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2. The first detection of the dust echo of the second flare shows a brightening of mag in both bands and can be expected to increase further in upcoming data releases. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al. 2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al. 2012). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al. 2012), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust, which is shown in Fig. 6. The change from bluer to redder colors come from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue.
Notably, AT2019aalc is not part of the Flaires sample (Necker et al. 2024), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs. This is because the broad application of the Flaires pipeline necessarily included a strict cut on extraneous variability that AT2019aalc did not pass. The authors did note however that a desirable relaxation of this criterion would include AT2019aalc. The source is, however, part of the dust echo sample of van Velzen et al. (2024) and, notably has the highest dust echo flux of all ZTF transients.
4.3 X-ray
AT2019aalc was monitored with the Swift/XRT at keV energies for weeks with weekly cadence around the peak of the optical flare and later for additional months with weekly and bi-weekly cadences. The luminosity evolution is shown in the bottom panel of Fig. 5 while the count rate variability is plotted in the top panel of Fig. 7. The first two visits show an increase of the X-ray emission. This flaring episode peaked at on 2023-06-29 (MJD) (see Fig. 5.) before it experienced a very quick decreasing. This flare peaked weeks before the optical peak. The count rate dropped by a factor of , from to , within a timescale of days. Later, the X-ray emission shows limited variability for several months, however, less significant flaring episodes appear that peaked in the end of September (MJD) and mid-February (MJD). The latter was followed by a bump in the optical and UV light curves around weeks later while no optical/UV data is available around the another X-ray increasing. In the end of April , a rapid and extreme flaring episode was detected in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by an increasing phase in the optical weeks later, according to the most recent ZTF observations999https://lasair-ztf.lsst.ac.uk/objects/ZTF19aaejtoy/. Notably, days passed between both the first and second and the third and fourth X-ray flaring episodes. While no data is available for a longer period between the second and third flares when the source was too close to the Sun.
In Fig. 8 we show the source spectrum derived from the full observing campaign. The spectrum is dominated by a power-law component but a blackbody component is also significantly presented. The spectral index of the stacked spectrum fitted with a power-law+blackbody model is while the blackbody temperature is eV. These imply an overall unabsorbed flux of F erg cm-2s-1 corresponding to a luminosity value of L erg/s. The count rate of the observations is cts s-1. The photon index variability of AT2019aalc is presented in the bottom panel of Fig. 7.
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al. 2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al. 2024). One observation with XMM-Newton Observatory (Jansen et al. 2001) has been performed in 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023). These imply that the two X-ray flares observed with the Swift/XRT peaked at much higher luminosities than the plateau or long-term flaring episode between the optical flares, however, there is no X-ray data from around the peak of the first optical flare. Furthermore, the eROSITA results confirm the presence of a blackbody component in the X-ray spectrum, which is very soft and dominating the variability of the X-ray source.
4.4 Radio
4.4.1 Temporal evolution
The host galaxy of AT2019aalc was detected in the NRAO VLA Sky Survey (NVSS, White et al. 1997) in February 1995 and in the FIRST survey in February 2000 (both at 1.4 GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability of the host galaxy prior to the discovery of the transient event.
Later, the transient’s sky location has been observed during the first and second epochs of the Very Large Array Sky Survey (VLASS). The VLASS covers the entire sky visible to the Karl G. Jansky Very Large Array (VLA) at GHz wide band offering an angular resolution of arcsec and an rms of mJy beam-1 in each epoch (Lacy et al. 2020). One of the main goals of the VLASS multi-epoch survey is to detect various types of extragalactic radio transients, such as TDEs, accretion state changing AGNs and core-collapse supernova (Lacy et al. 2020; Zhang et al. 2022). To date, the first two epochs of the survey have been fully published. Our source has peak flux density values of mJy/beam (on 2019-03-14) and mJy/beam (on 2021-11-06), corresponding to 3 GHz radio luminosities of and in the Epoch 1 and Epoch 2 images, respectively. This factor of increase in flux from the first VLASS epoch (three months before the optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al. 2024).
In addition, we found a radio source at the position of AT2019aalc in the first data release (DR1) of the Rapid ASKAP Continuum Survey (RACS; McConnell et al. 2020). The RACS DR1 includes initial observations made for RACS-low (central frequency of MHz)101010https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/racs/. Moreover, the source has been observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
Interestingly, the radio luminosity in the beginning of our ATCA monitoring is at the same level within uncertainties ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau. Our multi-frequency monitoring with the ATCA implies no substantial variability on monthly timescales at GHz while it shows slow decreasing at GHz and at GHz over time. We observed the radio source at GHz at one epoch. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 5 and focusing on the ATCA monitoring in Fig. 9.
4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities (these flux values are explained in Sect. 4.4.1.). The spectral index we derived between GHz and GHz this way is .
Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz. The spectrum of high frequency peaked radio sources are usually explained with significant absorption taken place in the system due to synchrotron self-absorption (although free-free absorption may also play a role, (e.g., Berton et al. 2020). We further discuss the features of the radio spectrum in Subsect. 5.3.
4.4.3 VLBI detection
The final naturally-weighted VLBI map of AT2019aalc is shown in Fig. 11. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of with our GHz EVN+e-MERLIN VLBI observation. We obtained the following parameters of the fitted component; a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999). The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al. (2005). This yields mas implying that the fitted Gaussian component is resolved. We note that at high SNR, can be significantly smaller than the size of the resolving beam (Bertero & de Mol 1996; Kovalev et al. 2005) due to applying a specific a priori hypothesis about the shape of the emitting region which is in our case the two-dimensional Gaussian used to fit the observed brightness distribution of the radio source.
The positional uncertainty was estimated as follows. When calculating the VLBI positional uncertainty, we took into account both the positional uncertainty of the phase calibrator source and its angular separation from our target source. The position of the phase calibrator source is listed with an error of mas in the third realization of the International Celestial Reference Frame (ICRF3; Charlot et al. 2020). The angular separation between the target and the calibrator source signifies a positional error of about mas at GHz (Chatterjee et al. 2004). The resolution of the observation depending on the thermal noise of the interferometric images (Reid et al. 1988) has been also considered. Following Reid et al. (1988), we estimated a thermal noise-limited positional accuracy of mas. The resultant positional uncertainty of mas is therefore dominated by the uncertainty arising from the angular separation between the calibrator and the target source. The radio position agrees with the source position in the Gaia DR3 (Gaia Collaboration et al. 2023) catalog within the positional uncertainties.
We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Condon et al. 1982; Ulvestad et al. 2005):
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015). It can be determined by imaging and modeling the structure of a given emitting region. The derived value is crucial to characterize the detected radio emission, which can be either thermal or non-thermal.
4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 12 with marking the notable emission lines. In the following, we highlight the remarkable line detection of the host spectrum observed before the first optical flare (pre-flare spectrum), the one observed between the two flares (post-flare-1 spectrum) and spectra that were taken during the second flare (over-flare-2 spectra).
4.5.1 Pre-flare spectrum
A spectrum of the host galaxy of AT2019aalc was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 13.
We identify strong Balmer emission lines. These lines are generally thought to be originating from the broad-line region i.e. the vicinity of the SMBH. We estimate the FWHM of the Balmer lines to FWHM(Balmer) kms-1 for the broad component of the Balmer lines (see Fig. 13 subplots) in the host spectrum, which is typical for an unobscured, broad line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al. 2019) for our source are consistent with this classification. In addition to the Balmer lines, we detect the recombination line He i which is also produced in the BLR.
The remarkable [O iii] and [O iii] forbidden lines exhibited already in the SDSS spectrum indicate that the host AGN was not completely dormant before the optical flares, which is further supported by the radio detections prior to the first flare (Subsect. 4.4.1). Other forbidden emission lines commonly seen in AGN such as [S ii] , [S ii] and [N ii] are detected as well. These lines are produced in the narrow-line region (NLR) in AGN. We note that the sulfur lines are also present in the spectra of the TDE AT2019dsg, which was hosted by a quiescent galaxy. We detect the [O ii] [O ii] line doublet, which is a potential indicator of star formation or AGN-driven outflows (e.g., Santoro et al. 2020).
4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was taken almost years after the first flare and years before the second one.
We detected an increase in the Balmer lines series and the forbidden lines [O iii] and [O iii] . The Balmer lines are also remarkably broadened.
In addition to the strong Balmer emission lines, we can identify the He ii He ii and transitions and the Bowen Fluorescence line O iii . Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. In the case of Bowen Fluorescence, the He ii–O iii and N iii coincidences lead the mechanism. The BF lines occur due to the excitation in certain states of O iii and N iii by the absorption of He ii Lyman photons. These excited states lead to a cascade of transitions that can be observed as emission lines in the optical and NUV regimes (Netzer 1990; Trakhtenbrot et al. 2019). Interestingly, the existence of BF emission lines in astronomical objects has been predicted by Netzer et al. (1985) as the required gas densities are consistent with what is expected from the extremely dense environment in some AGN and gaseous nebulae.
We detect the high ionization coronal lines [Fe vii], [Fe x] and [Fe xiv]. The very high ionozation potentials of eV and eV of [Fe xiv] and [Fe x] , respectively, point to very energetic processes at work. First Oke & Sargent (1968) studied coronal lines in detail in a Seyfert galaxy (NGC 4151). These lines are commonly detected in AGN and commonly thought to be powered by photoionization of the NLR (Cerqueira-Campos et al. 2021). In some extreme cases, a shock model i.e. collosional ionization for the coronal line region is required to explain the origin of these lines as it is the case in the Seyfert galaxy NGC 1068 (Oliva 1997).
We identify the [Ne V] emission lines at and in the spectrum. These lines are not strongly excited in the inner regions of AGNs, but are prominent emission lines in the lower density extended emission line regions (EELRs) excited by the AGN via photoionization or shocks. The forbidden line [O iii] also appears, however, bending from H is possible.
4.5.3 Over-flare-2 spectra
We performed optical spectroscopic observations of AT2019aalc during its second flare. These are characterized by a blue continuum with several lines appearing in comparison to the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017). This comparison shows similarities but also significant differences discussed below.
The line widths and strengths of the earlier detected broad Balmer emission lines (H, H and H) increased and both the line widths and strengths of the Balmer lines peaked days before the optical peak. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc we detected another BF line N iii together with the transition line He ii . This doublet has not been detected in the archival SDSS spectrum taken years before the discovery of the transient or in the post-flare-I spectrum. We detected this doublet already in the first spectrum taken during the second flare. The strengths of these lines increased significantly around the second optical flare and peaked on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we can conclude the long-term persistence of at least one BF line after the second flare as well. In addition, the second LRIS spectrum shows a clear increase in the BF line O iii and in He ii when comparing it to the post-flare-1 spectrum. Both lines increased by a factor of with respect to the post-flare-1 LRIS spectrum. We might detected O iii which is also known as a Bowen line (Selvelli et al. 2007), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which suggests its detection. O iii was detected in some metal-rich TDEs before (Leloudas et al. 2019) such as AT2019dsg (Cannizzaro et al. 2021).
We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line (Fe xi, ionization potential: eV) appears on 2023-08-17, around a month after the peak of the second optical flare. Interestingly, the line intensities of these lines are larger after the optical peak than around the peak in contrast with the BF lines and the above-mentioned ionized helium lines which experienced their maximums very close to the continuum peak. Moreover, [Fe xi and [Fe x] are not just persistent but the strongest in the late-time DBSP spectrum, taken almost a year after the optical peak of the second flare. We detect the high-ionization coronal line [S xii only in the late-time spectrum. This line is an extreme ionization line with an ionization potential of eV thus implies that unusually high level of ionization is taken place in AT2019aalc almost a year after its second optical peak. For this coronal line we estimate FWHM(S xii) km s-1.
We detected the [Ne V] emission lines at and in the second LRIS spectrum, similarly to the post-flare-1 spectrum and implying continuous ionization of the EELR. The [O iii] line appears in the late-time spectrum.
5 Discussion
5.1 Spectroscopic properties
5.1.1 Balmer lines
The line width of the Balmer lines and its limited variability are not common for regular TDEs and flaring AGN which tend to have decreasing and increasing Balmer line widths over time, respectively (Gezari 2021). However, this behavior is more compatible with the spectroscopic properties of the TDE-AGN PS16dtm (Blanchard et al. 2017). Still, the slight increasing and broadening indicate high-velocity gas moving in the BLR, close to the SMBH. In addition, a more complex structure of these lines seen in the late-time DBSP spectrum.
5.1.2 Bowen lines
The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii (Selvelli et al. 2007). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al. 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al. 1990), however, we note that due to the atmospheric limit at these lines are not easy to detect. This line is typically much weaker in normal Seyfert galaxies as in TDEs, BFFs or in our case (see Fig. 14). The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al. 2001), however, has been observed in several optical TDEs (Charalampopoulos et al. 2022b) and other SMBH-related transients; the BFFs AT2017bgt, F01004-2237, OGLE17aaj (Trakhtenbrot et al. 2019) and AT2021loi (Makrygianni et al. 2023), the peculiar transient event AT2019avd (Malyali et al. 2021) and a few flaring NLSy1s (Frederick et al. 2021). In a TDE case, the disrupted star’s material forms an accretion disk and the photosphere absorbs the EUVsoft X-ray photons originating from accretion processes and this way provide a plausible explanation for the appearance of the BF lines. BFFs have been first identified by Trakhtenbrot et al. (2019) who revealed three flares (AT2017bgt, F01004-2237 and OGLE17aaj) in AGN with the appearance of persistent Bowen lines. Later, Makrygianni et al. (2023) studied a transient, AT2021loi, classified as a BFF. The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the cases of the BFFs studied by Trakhtenbrot et al. (2019). In the case of AT2017bgt, this Bowen line was still detected days after the discovery of the transient. These lines in BFFs identified by Trakhtenbrot et al. (2019) may originate from a pre-existing BLR (a region with dense line-emitting gas) which was suddenly exposed to the intense ionizing UV emission, which also indicates enhanced accretion onto the SMBH. Alternatively, the BF lines might be related to a newly launched outflow driven by a sudden increase in accretion rate which interacts with the BLR. We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc with the spectra of the BFFs AT2017bgt and AT2021loi, the BFF candidate AT2019avd, the TDE-Bowen AT2018dyb and the repeating partial TDE AT2020vdq in Fig. 15 with focusing on the N iii He ii emission line doublet. We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 16 zooming into the O iii and He ii region.
The increasing of these lines around the optical peak indicate that these lines were enhanced by the Bowen Fluorescence mechanism. The BF process is particularly strong in some He ii, O iii and N iii transitions. These lines are obviously detected during the second optical flare, while atypical for normal Seyfert 1 galaxies (see a comparison plot in Fig. 14). Moreover, the O iii lines at and normally detected in AGN do not show up in our spectra, similarly to the case of AT2017bgt. These results suggest that the BF mechanism plays an important role and this mechanism traces significantly enhanced accretion onto the SMBH suggesting the unusual nuclear transient nature of AT2019aalc.
The He ii ionized helium line is presented in our spectra taken between the optical flares, and is also increased during the second optical flare. This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. Wang & Kron 2020). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al. 2019; Cannizzaro et al. 2021). In the spectra of AT2019aalc, He ii is present together with strong N iii and O iii BF lines and the less commonly detected He ii transition, indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of whose value significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al. (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ). Our value is, however, more comparable with the line ratio estimated for the BF transient AT2017bgt which is (Trakhtenbrot et al. 2019). The line intensities of He ii and N iii peak days before the continuum peak, suggesting extreme ionization by UV/soft X-ray photons around the peak. The He ii /He i ratio is sensitive to the change in the temperature and density as these lines belong to the same element but in two different ionization states (e.g., Ilić et al. 2009). We found the highest ratio close to the continuum peak, implying increased temperature in the BLR. The smallest ratio can be estimated from the post-flare-1 and late-time spectrum taken days after the continuum peak that is due to a drop in the temperature.
5.1.3 Coronal lines
We detected several high-ionization coronal lines. Two of the detected coronal lines present already in the post-flare-1 spectrum taken years after the first optical peak indicating a connection to the first optical flare.
Based on a study of SDSS spectra, Wang et al. (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines. Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to have coronal lines (Onori et al. 2022; Newsome et al. 2022; Short et al. 2023). The BFF AT2021loi (Makrygianni et al. 2023) and the spectacular transient event AT2019avd (Malyali et al. 2021) have significantly detected coronal lines, too. Based on the publicly available spectra of AT2017bgt on the Transient Name Server111111https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well. Wang et al. (2012) states that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later supported by Short et al. (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. Interestingly, the coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al. 2012). According to Wang et al. (2012), the ECLEs could be explained through the coronal lines arising from Fe liberated from dust grains that were destroyed by the TDE flare. The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection clearly have more luminous IR dust echo compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al. (2024) due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al. 2024). The Bowen line transient AT2019pev that shows strong high-ionization coronal lines (Frederick et al. 2021; Yu et al. 2022) is also part of the MIR flare sample of Hinkle (2024). These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs. Fig. 17 shows the [Fe x] versus the [O iii luminosities of the EECLEs studied by Wang et al. (2012), AT2019qiz (Short et al. 2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al. (2009). This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
The [Fe x] line evolution is shown in Fig. 18. The strength of this coronal line in AT2019aalc is further increased during the second flare and seem to be peaking lately with respect to the continuum peak. The coronal line TDE AT2019qiz shows a late-time increase in the this line as well (Short et al. 2023). Similarly to the case of AT2019qiz, we also found that [Fe x line is the strongest high-ionization coronal line presented in the spectra, while [Fe vii is the weakest one. The coronal line [Fe x] has been detected in several AGN before, however, the maximum line intensity ratio we calculated ([Fe x / [O iii] ) clearly overshoots the maximum ratio (, Nagao et al. (2000) found for AGN. For AT2021loi, Makrygianni et al. (2023) calculated even higher ratios, up to .The line width of [Fe x] shows a clear evolution during the monitoring of AT2019aalc as we estimated values of FWHM([Fe x]. Interestingly, the line widths we estimated for [Fe x are clearly larger than the maximum values found for coronal line emitters (around km s-1, Wang et al. 2012), however, very consistent with the width estimated for AT2021loi. For AT2021loi, Makrygianni et al. (2023) gives a line width of km s-1 which implies that the line is consistent with a BLR instead of NLR origin since the latter would require the typical values estimated for coronal line emitters ( km s-1, Wang et al. 2012). The BLR origin would naturally explain the increasing emission in these lines after the optical peak (due to the light travel time) but is also consistent with the distance between the dust responsible for the reprocessed IR emission and the SMBH ( pc, van Velzen et al. 2024). This suggests a connection between the unusually large dust echos and the extreme coronal lines in these transient sources. An alternative explanation is collisional excitation of these coronal lines. This is compatible with a detailed multi-wavelength study of AT2019qiz which suggests that outflows (powered by accretion or stream collision) are responsible for driving the optical flare of this coronal-line TDE (Nicholl et al. 2020).
The sulfur line [S xii has the second highest ionization potential among the optical lines detected in AGN before (the highest one is [Ar xiv which we do not detect in AT2019aalc). We zoom into the SPS fitted DBSP spectrum to show the detection of this line in Fig. 19. We can reasonably assume that the most recent and very strong X-ray flare, started less than a month before the appearance of the sulfur line, is responsible for the creation of this emission line. Apart from ECLEs in the sample of Wang et al. (2012), [S xii] has only been reported in Seyfert galaxies (Kraemer & Crenshaw 2000; Mazzalay et al. 2010; Cerqueira-Campos et al. 2021; Oh et al. 2022) including the neutrino emitter (IceCube Collaboration et al. 2022) NGC 1068 and in the nuclear transients VT J1548 (Somalwar et al. 2022) and AT2022fpx (Koljonen et al. 2024). VT J1548 is a radio-detected transient that was classified as a promising BFF candidate and notably with remarkable TDE-like characteristics (Somalwar et al. 2022). Similarly, AT2022fpx shares properties both with TDEs and BFFs (Koljonen et al. 2024). To our knowledge, AT2019aalc is so far only the third nuclear transient detected with a line requiring such a high creation ionization potential.
5.1.4 Forbidden lines
We find evidence for the forbidden line [O iii] in the post-flare-1 LRIS spectrum and the late-time DBSP spectrum. Its detection, however, remains ambiguous because the Balmer line series are complex in our spectra, including blending from H. Nevertheless, the detections of [O iii] years after the first flare and year after the second flare is not unexpected. After all, this emission line requires densities much lower than the UV-line emitting clouds, and consistent with the outer NLR (or simple ISM) clouds (Ji et al. 2024) but is located inner than the emitting region of [O iii ] (Nagao et al. 2001). The line ratio of [O iii] /H suggests extremely hard ionizing radiation from the ionizing source. Using James Webb Space Telescope (JWST) observations, this line is sometimes detected and studied in detail in high- AGN (used to distinguish AGN from star-forming regions), however, such a ratio was detected only in a very few cases (Übler et al. 2024; Ji et al. 2024; Mazzolari et al. 2024). The [O iii ]/[O iii] ratio of is clearly higher than calculated for AGN which are typically below (Baskin & Laor 2005; Übler et al. 2024; Binette et al. 2024)).
The highly ionized neon lines typically originate from star formation activity or photoionization of the EELR by the AGN. In some extreme cases, high-velocity AGN-driven shocks that ionize extended emission regions are expected to be responsible for the presence of the lines [Ne V] at and (Maddox 2018). The strengths of these lines in AT2019aalc are higher than the [O ii] [O ii] doublets which indicate their origin related to AGN activity since [Ne V] is not excited by star formation (Maddox 2018). These shocks are possibly, but not exclusively, related to radio jets as high shock velocities of km s-1 are required to produce these lines (Maddox 2018). We estimated slightly lower velocities, which are therefore more consistent with the EELR origin and significant AGN photoionization in this outer region. From the increasing line intensities of [O iii] and [Ne V] we can conclude that the NLR and EELR are also strongly ionized, however, in the late-time spectrum these decreased faster than the Bowen line N iii. The ionization level of the BLR is therefore more persistent than in the outer regions.
5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well separated flares. Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales. The first flare decayed for more than a year, while the second flare has been decaying for around a year.
Several ZTF monitored extragalactic transients have been found with bumps seen in their optical light curves (see Soraisam et al. (2022) for a recent review)121212We note that ZTF has not been running long enough to discover flare repeats within longer timescales e.g., a decade that actually might also be common. Most of them are SNe, however, a couple of them have been identified as TDEs (AT2020acka, AT2020nov and AT2021ehb). The latter two TDEs show prominent bumps in their light curves while decaying, while AT2020acka exhibits a long and monotonic bump. These TDEs are, however, experienced these bumps within a year from their peak. Moreover, the second flare of AT2019aalc peaked even at higher luminosities than the initial one. AT2020nov and AT2020acka experienced less significant second bumps, while AT2021ehb has two comparable peaks. The optical emission of these transients are poorly understood, and Soraisam et al. (2022) also states that their identifications as TDEs are ambiguous. Generally, TDEs are characterized by smoothly and rapidly declining light curves ( as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that has started years after the initial flare. The first systemically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al. 2023). This implies an even more significant optical re-brightening than the one experienced by AT2019aalc. The second flares of partial TDEs are expected to follow an extremely steep power-law when declining. In the case of AT2020vdq, the second flare declined much slower (Somalwar et al. 2023) than this but still clearly faster than the second flare of AT2019aalc which decayed with a power-law index of . Other candidate repeating TDEs have been found with periodic flares at different wavelengths; ASASSN 14ko with more than observed optical/UV/X-ray flares (Payne et al. 2021), eRASSt J045650.3-20375 with X-ray/UV flares (Liu et al. 2023), AT2022dbl with two optical flares (Lin et al. 2024), while AT2018fyk (Wevers et al. 2019a, 2023) and RX J133157.6-324319.7 (Hampel et al. 2022) showed a one-time X-ray/UV and X-ray re-brightening episodes, respectively. In these cases, due to its grazing orbit, the star orbiting in the vicinity of the SMBH is thought to be only partially disrupted. After another pericenter passage, the surviving core thus feeds a new accretion flare, resulting in X-ray/UV re-brightening. For the Bowen-line and coronal-line emitter transient AT2019avd with two distinct optical flares, Chen et al. (2022) discusses a partial TDE scenario.
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 20 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly seen in the binned light curve published in Trakhtenbrot et al. (2019). One of the two other BFFs originally studied by (Trakhtenbrot et al. 2019), F01004-2237, experienced a significant optical re-brightening around years after its initial flare detected in . Originally, Catalina Sky Survey (CSS) light-curve data revealed that this ULIRG had undergone a luminous continuum flare (Tadhunter et al. 2021) and later Trakhtenbrot et al. (2019) classified it as a BFF similar to AT2017bgt. The binned light curve of this source published in Makrygianni et al. (2023) shows a bump around a year after its peak. We investigated the recent activity of the source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and decayed faster. In addition to these two BFFs, AT2021loi shows a bump during its decaying, roughly a year after the first flare reached its peak. Makrygianni et al. (2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date. The first flares of the studied BFFs decay on yearly timescales, i.e. much slower than TDE flares but more similar to AT2019aalc. However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution in addition to the similarities we see. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses have been defined based on their light curve evolution (e.g., Charalampopoulos et al. 2023).
5.3 Possible explanations of the long-term radio flare
In this subsection we discuss the long-term radio variability of AT2019aalc. The source experienced a brightening of its flux density by a factor of on a timescale of less than years followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring. Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., Mundell et al. 2009; Koay et al. 2016) and therefore require further investigation. The unusual radio spectrum with a high frequency turnover is also discussed below.
The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (Condon 1992). Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see Bontempi et al. 2012 as an example how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies). This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, Gallimore et al. 1997) as the origin of the detected radio emission.
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997). Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that the advection-dominated accretion flows cannot explain the observed radio emission in our case.
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG; Green et al. 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993). Later, Behar et al. (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. Magnetized coronal winds were proposed to explain the observed radio emission of the Seyfert galaxy Mrk 590 by Koay et al. (2016), however, the authors noted that a jet contribution cannot be ruled out. These two types of outflows might even be equivalent and cannot be distinguished in Seyfert galaxies with unresolved compact core features seen in their VLBI maps (e.g., Kharb et al. 2015), like in the case of AT2019aalc. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be rapidly variable at radio wavelengths and should evolve similarly to the X-ray flux, which is clearly not the case. Although, a contribution from the mechanisms discussed so far to the total radio flux remains a possibility, we need to look for another possible origin to explain the main origin.
Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets which can be either on or off-axis relative to the observer. These features are often visible on VLBI radio maps of radio-detected Seyfert galaxies (e.g., Kozák et al. 2024) and a few TDEs (e.g., Paragi et al. 2017; Mohan et al. 2022). In our case, the high brightness temperature of the EVN-detected radio feature together with its position (in good agreement with the optical Gaia position) indicate its radio core identification, however, no other components were detected. Nevertheless, the steep () radio spectrum revealed by our ATCA observations indicates the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum. The steep radio spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s). The ATCA radio spectrum indicates a high-frequency turnover above GHz which is a possible hint of a newly ejected radio-emitting component. We further discuss the radio spectrum of AT2019aalc in the following.
The turnover suggests the multiple component nature of the spectrum. Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al. 2020). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. Since older sources are larger than young ones, the expanding relativistic jet or non-relativistic outflow loses energy as the density decreases and the spectral peak moves to lower energies over time. With other words, as the projected linear size of the radio source increases, the turnover frequency decreases. These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al. 2020). The low-frequency peak is, however, more likely to be explained with star formation activity (as NLSy1 galaxies have a significantly higher star formation rate as other AGNs) instead of AGN emission. The latter is assumed to be synchrotron self-absorbed (SSA) to explain the faintness or the non-detection of these galaxies at lower frequencies. As another possibility to explain the high-frequency excess of, Berton et al. (2020) discusses a scenario in which the AGN emission is absorbed below GHz due to free-free absorption (FFA). In that case, the circumnuclear gas might be ionized both by the AGN and by hot stars and might act as a screen for the jet at low frequencies or the passage of the jet in the interstellar medium causes the formation of a cocoon of ionized gas and this way being responsible for the FFA. Finally, coronal emission could also be responsible for the inverted spectrum.
In the case of AT2019aalc, the coronal emission scenario can be safely excluded as we should see a rapidly variable radio light curve which should evolve similarly to the X-ray emission. On the other hand, the other two scenarios might also be challenging. The large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA. Moreover, the presence of a relatively strong [O iii] line in the SDSS host spectrum implies earlier AGN activity as being responsible for the radio emission before the long-term flare. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
Nevertheless, the contribution of a newly ejected component to the overall radio spectrum remains a possibility. In the cases of the earlier discussed NLSy1 galaxies, the high frequency peak is thought to be suggesting a newly born radio source with no signs of earlier jet activity. Similarly to gigahertz-peaked sources (GPS), the high-frequency peak is thought to be shifting towards lower energies over time. This way, NLSy1 galaxies might be extremely young GPS which was first suggested by Oshlack et al. (2001). In our case, a newly ejected component can be responsible for the high-frequency excess whilst the low-frequency part is naturally explained with optically thin synchrotron emission of the AGN indicating that we do not see a newly born radio-AGN but an AGN with recently enhanced, intermittent activity. The result is a two-component radio spectrum with a relatively steep spectrum at lower frequencies and a high-frequency excess. The steepness of the spectrum between MHz and GHz might be explained with cooling of electrons originating from past AGN activity, and also a contribution from star formation can play a role here. Järvelä et al. (2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al. (2020). The broad-line AGN classification of the host galaxy of AT2019aalc is more consistent with a restarted AGN picture as this type of AGN are thought to be older than NLSy1s making a restarted activity scenario over a newly born radio source one more probable. The restarted activity results in a peaked spectrum at higher frequencies, superimposed on a spectrum from a period (or periods) of earlier activity peaking at lower frequencies. Intermittent activity is further supported by theoretical models suggesting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al. 2009). This is compatible with the accretion-induced lines seen in the post-flare-1 optical spectrum of AT2019aalc (see Subsubsect. 4.5.2 and van Velzen et al. 2024).
We note that the fitted ellipsoid model fit component to the EVN visibility data is aligned to the northwest-southeast direction (clearly differs from the position angle of the restoring beam indicating the angle-dependent resolution of the VLBI array) which potentially indicates a newly ejected radio-emitting component. Higher frequency observations with improved angular resolution are required in order to resolve the radio structure we detected at GHz. We observed AT2019aalc with the EVN and the Korean VLBI Network (KVN) jointly at GHz in March and expect to reach sub-mas-scale angular resolution, roughly an order of magnitude better than with the currently discussed observation. These results will be presented in a follow-up publication
5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al. 2015). Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess. This soft excess can be well modeled by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al. 2012).
The stacked Swift/XRT X-ray spectrum of AT2019aalc can be adequately described with an absorbed blackbody+power-law model. We note that in the cases of strongly beamed jetted sources e.g. blazars, the X-ray spectrum can be well modeled with a simple or broken power-law without adding any additional model component (e.g., Kammoun et al. 2015; Acciari et al. 2022). Although AT2019aalc shows enhanced radio activity, its brightness temperature suggests a non-beamed nature of its radio emission in agreement with the X-ray spectrum.
The X-ray spectrum is dominated by a power-law component with a soft power-law index of while the thermal component is extremely soft with a blackbody temperature of eV derived from the Swift/XRT observations. Most of the BLSy1s typically have a significantly harder X-ray spectrum with a power-law index (e.g., Pons & Watson 2014). Our value is more typical for sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al. 1996; Saxton et al. 2020).
The soft X-ray excess appears in the stacked spectrum and becomes more significant over time as we added a blackbody component in order to better describe the X-ray spectrum. Furthermore, the data of the peak of the second X-ray flare is clearly better-fit with a single blackbody model than with a power-law one (Subsect. 4.3) implying not just the need of the blackbody component to adequately fit the data but even its dominance over the power-law component at this epoch. A softening of the power-law index over time can be seen as well. These results suggest significantly enhanced accretion or disk instability. Delayed accretion flares have been observed in TDEs (Saxton et al. 2020). In our case, the post-flare-1 spectrum shows clear signs of accretion and the second flare is followed by enhanced accretion activity as well.
The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x]. The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al. 2009). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability we consider only observations taken on the same day). This implies that the expected soft X-ray luminosity is consistent with the observed values. The soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as it was suggested for ECLEs by Wang et al. (2012). The photoionization or collisional excitation of the dust grains is rather due to a current event. This further supports that the reoccurring optical/UV outbursts seen during the decaying of the flares are driven by the soft X-ray flares which lead these outbursts only with a couple of weeks. These suggest an accretion-driven origin of the reoccurring outbursts in the optical/UV during the decaying of the flare.
5.5 SED fitting
Following Reusch et al. (2022), we fitted two blackbody models in order to describe the SED of the transient using lmfit. These fitted curves to the optical/UV (’blue’ blackbody) and IR datapoints (’red’ blackbody) are shown in Fig. 21, together with a combined double blackbody fit. The red blackbody fit was introduced to account for the IR dust echo. These curves fit the SED around the optical peak of the investigated flare.
The continuum emission of TDEs are well described by a thermal blackbody model (Gezari 2021). The case of AT2019aalc appears to be more complex. An excess in the UW2 band can be seen regardless the consideration of a simple or the double blackbody fit. The UV emission have a contribution from the Bowen Fluorescence mechanism and potentially also from Comptonization explaining the poor fit towards the blue. The red wing of the optical and the IR part cannot be well fitted with a simple blackbody model suggesting a complex, multi-component dusty reprocessing region. The excess in the -band might be explained with the early appearance of some Fe lines in this range (such as [Fe x] and [Fe vii]) indicating the release of this element previously locked up by the dust grains. Since all of this seem to be started already around the optical peak, at least a fraction of the dust is located on subpc scales to the SMBH. This is further consistent with the extreme line widths of the coronal lines (see Subsect. 5.1.3). Therefore, not only the dusty tori is responsible for the IR dust echo. It is possible that the significant amount of dust in the nucleus region is the consequence of enhanced star formation sometime in the past. The Baldwin, Phillips& Terlevich (BPT, Baldwin et al. 1981) emission-line ratio diagnostics of the host SDSS spectrum are log([O iii] and log([N ii] . We plotted the BPT line-ratio of SDSS spectra with indicating the ratios of the host galaxy of AT2019aalc and the BFFs AT2021loi, AT2017bgt and F01004-2237 in Fig. 22 using astroml (Vanderplas et al. 2012; Ivezić et al. 2014). All of the investigated galaxies are in the composite region, consequently, in addition to AGN activity, star formation activity might also have a remarkable contribution for producing the narrow lines seen in the spectra. Moreover, these ratios place these two galaxies solidly in the region of the diagram where the post-starburst sample of French et al. (2015) is located. We note that to derive the ratios for AT2019aalc and F01004-2237, we used pre-flare spectra. While as no host spectra are available for AT2017bgt and AT2021loi, we considered the spectra taken as early as possible in these cases.
5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021). Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the decay of the optical emission, which is significantly slower than for any TDE identified so far. The color evolution of the flares is also atypical for TDEs.
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of strong high-ionization coronal lines.
The mechanism that powers the long-term accretion enhancement in BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al. 2017) or the AGN’s radiation field. The bumps in optical and UV during the decay of the flares of AT2019aalc led by soft X-ray flares indicate the flare’s accretion-driven origin consistently within this picture. Moreover, the flaring episodes in X-rays seem to be happen periodically which is expected from interaction between a circularizing stellar stream and the accretion disk. The origin through reprocessing is further consistent with the ionization of the BLR and so on the increase in the Bowen and the Balmer lines seen in the spectra during the second flare. The dusty region (expected from a source with IR dust echo) and the liberation of Fe and S from these dust grains due to photoionization by the soft X-ray flares provide an explanation for the high-ionization coronal lines. The increase of the W1W2 color implies significant cooling expected from cooling dust after the onset of the dust echo. This scenario provides explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. The second flare is even more luminous than the first one, similarly to the case of the partial TDE AT2020vdq (Somalwar et al. 2023). A partial TDE scenario might explain the second optical flare.
Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well. The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understood their unique optical light curves. Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs. This would explain the TDE-like properties of these sources. The most remarkable differences i.e. the slowly decaying optical emission and the continuum emission not being well described by blackbody models might appear due to a more complex environment the TDE takes place with comparing to quiescent galaxies. Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al. 2012) and the BFFs is consistent with this picture. A known potential partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al. 2015). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 17). Interestingly, IC 3599 is the only known AGN with fading coronal lines (Frederick et al. 2019).
Alternatively, a double TDE i.e. the tidal disruptions of a pair of binary stars might explain the double-peaked optical light curve. Notably, double TDEs may represent nearly of all stellar tidal disruptions (Mandel & Levin 2015).
The strength of the high-ionozation coronal lines that regular AGN activity cannot explain the unusual behaviour of this transient. It is not clear why we do not see as strong high-ionozation coronal lines in normal AGNs as in the transients with significant dust echo emission i.e. TDEs and BFFs. It might be explained with dust sublimation due to the UV radiation of the AGN (the sublimation radius is typically in the subpc range, van Velzen et al. 2021b) which clears the dust from the nucleus region. AGN-driven outflows play also a role here. The limited variability of the host galaxy of AT2019aalc prior the first flare and its location in the BPT diagram suggest weak AGN activity before the transient occurred. Therefore, the dust possibly originating from star formation activity in the nucleus could survive. Additionally, the first flare could not clear the nucleus region which suggests that it must have been way less luminous in UV than the second flare. Unfortunately, due to lack of UV data around the first flare this hypothesis cannot be further tested. Nevertheless, the detection of some Fe lines in the post-flare-1 spectrum implies that a certain fraction of the dust was destroyed following the first flare.
Interestingly, the host AGN of other BFFs also showed weaker pre-transient activity than persistently accreting AGN (Trakhtenbrot et al. 2019; Makrygianni et al. 2023). The TDE-like features suggest that TDEs might restarted these AGN. Beyond the TDE-related scenarios, it is possible that the restarted AGN activity is rather explained with the dusty environment itself.
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs available. Based on the radio observations explained in Makrygianni et al. (2023) AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz which can be compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The host galaxy of AT2017bgt was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. In addition, the BFF candidate VT J1548 was also found to exhibit a radio spectrum peaking around GHz (in contrast with non-jetted AGN that peak below GHz, e.g., Padovani et al. 2017) which was explained with an ejection of a synchrotron-emitting outflow (Somalwar et al. 2022). These findings allow us to conclude that newly ejected radio outflow/jet components play a crucial role in BFFs. These outflows ejected due to enhanced accretion might be partly responsible for the liberation of Fe and S from the dust grains located relatively close to the SMBH. The collisional ionization in the case of AT2019aalc is supported by the extreme broadness of the coronal lines especially [S xii.
5.7 BFFs and high-energy neutrinos
AT2019aalc has been associated with the high-energy neutrino event IC-191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al. 2024). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1; Abbasi et al. 2023). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC-230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Figs. 24 and 24. We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
The neutrino associated with AT2021loi is days delayed with respect to its first optical peak and days with respect to its second optical peak. Note, that the nuclear transients (including AT2019aalc) are associated with neutrinos that were delayed by days, which might be explained with delayed mass accretion, delayed jet formation or outflow ejection (van Velzen et al. 2024). In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al. 2023) is compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc. However, from the current data, we cannot tell if the radio-emitting region of AT2019aalc has been ejected before or after the neutrino. Although we do not see a clear increase in the MHz RACS data months after the neutrino arrival. However, if the radio flare is indeed related to a newly ejected, synchrotron self-absorbed component, it is not surprising that it does not result in any increasing at low frequencies at early times. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al. 2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients. These findings indicate that outflows play an important role in the neutrino production. Moreover, Winter & Lunardini (2023) estimate the highest neutrino fluence for AT2019aalc among the studied nuclear transients associated with neutrinos. A similar case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al. 2024; Hinkle 2024; see also Subsect. 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al. 2023; see also Subsect. 5.1).
6 Conclusions
We conducted a multi-wavelength campaign during the second optical flare of the nuclear transient AT2019aalc associated with a high-energy neutrino event. We found that the second optical flare is even more luminous and decays even slower than the first one. Both flares evolve rapidly and with constant color. The optical flare is accompanied by UV emission which peaks at extremely high luminosities (at erg/s). The IR emission increases, similar to the first flare when it peaked delayed with respect to the optical and was explained with a dust echo flare. We detected a very soft X-ray flare that peaked before the optical/UV emission peaks. Later, the Swift/XRT detected an almost times more luminous X-ray flare which is clearly dominated by a blackbody component, suggesting enhanced accretion or disk instability.
Our VLBI observation provides a hint of extension of the radio-emitting source and clearly favors an AGN-like origin over a star-formation scenario. The radio monitoring revealed a long-term flare and an inverted spectrum above GHz, which is most probably explained by a newly ejected, self synchrotron-absorbed outflow or jet component. As we do not see obvious signs of variability well before the first optical flare, we suggest that the transient event in renewed the activity of the host AGN. Restarted AGN tend to have inverted radio spectrum. This also naturally explains the long-term flare.
The results from optical spectroscopy help us to better understand the event responsible for restarting the AGN. In the optical spectra, the presence of persistent Bowen lines are consistent with intensified accretion over a year after the second flare. The UV bright nature of the transient is consistent with the Bowen mechanism which pumps lines mostly in this range. The line intensities of the Bowen lines and the temperature of the BLR both peak around the peak of the optical flare indicating that the optical flare is driven by the ionization of the clouds of most probably the BLR. As we find the X-ray flaring episodes prior to the rise in optical/UV, the origin of the latter is most probably reprocessed X-ray photons which are also required for the extreme ionization implied by the Bowen lines. The X-ray flares are very soft which explains the unusually strong He ii lines. These findings altogether suggest the accretion-driven origin of the flaring episode we monitored. Note, that reprocessed soft X-ray emission i.e. by an outflow surrounding the disk is a possible explanation for the origin of the optical/UV emission of TDEs (Bu et al. 2022).
We detected several persistent or even late-time increasing high-ionization coronal lines in the spectra of AT2019aalc. The detection of [S xii requires an extremely high ionization potential. To our knowledge, AT2019aalc is only the third extragalactic transient event detected with this line and only Seyfert galaxies (including the neutrino emitter NGC 1068) and ECLEs were reported with a line with such a high ionization potential. The Fe/X-ray ratios disfavor a past TDE interpretation responsible for the coronal lines, which was proposed to explain these lines seen in most of the ECLEs studied earlier. The slowly decaying optical emission might instead be explained with continuous photoionization of the dust grains (located at distances consistent with the BLR instead of the NLR as seen for ECLEs before). This is consistent with the early observation of the dust echo emission with respect to the optical peak which suggests that the dusty region is located very close to the SMBH. Based on the narrow-line ratios, we found that enhanced star formation activity is a favorable explanation for the dusty environment. Moreover, the gas-rich environment might explain the restarted activity of the AGN.
The high [Fe x / [O iii luminosity ratios of the BFFs, coronal line-detected TDEs and ECLEs suggest a possible connection between these source types. These sources tend to have large dust echo-like flares, which implies that the liberation of Fe from the dust grains can be explained via their TDE(-like) flaring episodes. The significant photoionization is supported by the detection of strongly ionized neon lines in the EELR. A delayed outflow/jet component and collisional excitation provides an alternative explanation for the coronal lines and their late-time increasing or appearance. The BFF and coronal-line emitter AT2021loi has an inverted radio spectrum and based on survey data we found a flaring radio source at the position of AT2017bgt. These findings suggest that newly ejected radio-emitting components play an important role not only in BFFs, but might be connected to the coronal line emission. A detailed radio study of other BFFs and coronal line emitters is crucial to better understand the appearance of these highly ionized lines.
Altogether, our results are consistent with the Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter. A connection between the ECLEs, some TDEs and BFFs suggested by our spectroscopy results makes a non-standard TDE-related scenario more probable to explain the optical flares and i.e. the restarted activity of the host AGN. Moreover, due to the pre-existing BLR and therefore the high densities, TDE-AGN are more capable for the Bowen Flurosence mechanism and the high-ionization coronal lines. Only a few TDEs were detected with coronal lines; these are also known as TDEs with extreme dust echo emission.
Continuous interactions of a pair of binary star and the SMBH favored by Trakhtenbrot et al. (2019) might explain the re-brightened and slowly decaying optical emission of BFFs. An alternative scenario for the two distinct optical flares is a partial disruption of a single star in an AGN. Here, the slowly decaying reprocessed optical/UV emission originates from soft X-ray flares powered by continuous tidal interaction between the stellar debris and the pre-existing accretion disk. The double-peaked optical light curve is explained with a second disruption due to the surviving core after the first passage. However, more BFFs are needed to study the mechanism responsible for these transients. TDE-related scenarios nevertheless provide an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the ionized gas. Future observations will help us to better understand the nature of the transient; a binary TDE would not produce a third flare, a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC-230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al. (2024). Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionozation coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode/bumps in optical. In both cases the host galaxies are classified as broad-line AGN with remarkable star formation activity.
Acknowledgements
We thank Benny Trakhtenbrot for useful discussions. PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal. EH acknowledges support by NASA under award number 80GSFC21M0002.
The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: EV027. e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. We acknowledge the use of public data from the Swift data archive. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. The Australia Telescope Compact Array is part of the Australia Telescope National Facility https://ror.org/05qajvd42 which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the Observatory site. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. These results made use of the Lowell Discovery Telescope (LDT) at Lowell Observatory. Lowell is a private, non-profit institution dedicated to astrophysical research and public appreciation of astronomy and operates the LDT in partnership with Boston University, the University of Maryland, the University of Toledo, Northern Arizona University and Yale University. The upgrade of the DeVeny optical spectrograph has been funded by a generous grant from John and Ginger Giovale and by a grant from the Mt. Cuba Astronomical Foundation. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.

### v2

@@
 Aims. We aim to classify the transient and explain the mechanism responsible for its second optical flare.
-Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023.
+Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023/2024.
 Results.
@@
 About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum.
-Despite most of these transient sources have been discovered by optical telescopes, the origin of their optical/UV emission is poorly understood.
-One model is colliding stellar debris stream shocks (Piran et al.
-2015).
-Another scenario is reprocessed soft X-ray/far UV emission (originating from the debris accretion onto the black hole) by an extended elliptical disk, or an envelope, or an outflow surrounding the central debris disk (for a recent review see Bu et al.
-2022).
-TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying optical emission.
-Several TDEs were detected in X-rays with a very soft X-ray spectrum typically explained by late-time accretion disk formation (Saxton et al.
+TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying emission.
+Several TDEs were detected in X-rays with a very soft spectrum typically explained by late-time accretion disk formation (Saxton et al.
 2020).
@@
 Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H.
-As the photosphere shrinks over time, the flux of the ionized lines such as He ii tend to increase while the He ii/H ratio increases (e.g., Hung et al.
-2017; Charalampopoulos et al.
-2022a) and, in a few cases, Bowen Fluorescence (BF; Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al.
+In a few cases, Bowen Fluorescence (BF: Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al.
 2019; Charalampopoulos et al.
 2022a).
-To date, roughly promising candidate TDEs are known.
-However, since most of the above-mentioned observed properties do not exclusively characterize TDEs, spectroscopic coverage is required to distinguish TDEs from ‘TDE impostors’ such as conventionally accreting active galactic nuclei (AGN), changing-look AGN (CLAGN), supernovae (SNe) or even stellar collisions (Zabludoff et al.
-2021).
-TDEs generally result in one optical flare, however, in a few cases a re-brightening episode (or episodes) was detected even years after the initial flare.
+TDEs generally result in one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare.
 These might be explained in the following way.
 Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH.
-The secondary disruption of a surviving core after it reaches the pericenter again results in X-ray re-brightening explained by renewed accretion.
-The first systematically identified repeating partial TDE, AT2020vdq, has been classified after simultaneous multi-wavelength monitoring (Somalwar et al.
-2023).
-Only a few other partial TDEs or candidates are known to date, these are ASASSN 14ko (Payne et al.
-2021); eRASSt J045650.3-203750 (Liu et al.
-2023); RX J133157.6-324319.7 (Hampel et al.
-2022); AT2018fyk (Wevers et al.
-2019a, 2023); AT2019avd (Chen et al.
-2022) and AT2022dbl (Lin et al.
-2024).
+The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening.
+Only a few partial TDEs or candidates are known to date.
+These are AT2020vdq (Somalwar et al.
+2023), ASASSN 14ko (Payne et al.
+2021), eRASSt J045650.3-203750 (Liu et al.
+2023), RX J133157.6-324319.7 (Hampel et al.
+2022), AT2018fyk (Wevers et al.
+2019, 2023), AT2019avd (Chen et al.
+2022), AT2022dbl (Lin et al.
+2024) and AT2021aeuk (Sun et al.
+2025).
 Based on the observed properties of TDE host galaxies, French et al.
@@
 2024).
-The multi-wavelength characteristics of such a TDE should clearly differ from the ones taken place in quiescent galaxies, as the AGN is surrounded by a pre-existing accretion disk.
-Due to the extreme difference in density between the two, neither the star nor the accretion disk are affected after the first passage through the disk.
-However, the confined debris stream has such a low density after the star is tidally disturbed that when it interacts with the disk the collision between the stream and the disk can potentially dissipate much of the kinetic energy possessed by the stream and the disk gas near the impact point, and it can be radiated away (Chan et al.
-2019).
-These transients are challenging to distinguish from AGN variability.
+These transients are challenging to distinguish from regular AGN variability.
 Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (Blanchard et al.
@@
 2021).
-In addition to TDEs, flaring NLSy1s, CLAGN and different types of SNe were classified as typical nuclear transients.
+In addition to TDEs, flaring NLSy1s, changing-look AGN (CLAGN) and different types of SNe were classified as typical nuclear transients.
 Recently, Trakhtenbrot et al.
-(2019) identified a new class of flares from accreting SMBHs.
+(2019) identified a new class of spectroscopically unique flares from accreting SMBHs.
 In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features.
-Later, the peculiar X-ray loud transient AT2019avd (Trakhtenbrot et al.
+Later, the peculiar X-ray transient AT2019avd (Trakhtenbrot et al.
 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear.
@@
 (2019).
-Those show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission.
-These AGN are characterized by significantly and longer-term intensified accretion and might be powered by a binary TDE or an interaction between an outflow and the broad-line region (BLR) or binary SMBHs (Trakhtenbrot et al.
-2019).
-Notably, these transients take place in different types of AGN; AT2017bgt and OGLE17aaj in narrow-line AGN, AT2021loi in a broad-line AGN, F01004-2237 in an ultra-luminous infrared galaxy (ULIRG) while AT2019avd occurred in an earlier quiescent galaxy.
-Hereafter, we will refer to this new class of AGN flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al.
+These events show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission.
+These AGN are characterized by significantly and longer-term intensified accretion.
+Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al.
 (2023).
@@
 (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare.
-Other candidates, namely VT J1548 (Somalwar et al.
-2022) and AT2022fpx, (Koljonen et al.
-2024) were found recently.
+Two other candidates, namely AT2022fpx, (Koljonen et al.
+2024) and AT2020afhd (Arcavi et al.
+2024) were found later.
 In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al.
 2023).
-The TDE AT2019dsg is likely associated with the IceCube neutrino, IC191001A (Stein et al.
-2021) and a further promising association, between the candidate TDE AT2019fdr and the IceCube neutrino IC200530A (Reusch et al.
+The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Stein et al.
+2021) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Reusch et al.
 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos.
-Both sources were found to have unusually luminous mid-infrared emission, and a search for similar flares (van Velzen et al.
+Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (van Velzen et al.
 2024) led to the identification of the transient event AT2019aalc (a.k.a.
-ZTF19aaejtoy) with the neutrino alert IC191119A (see Sect.
-2.
-for explanation).
+ZTF19aaejtoy) with the neutrino alert IC191119A.
+In addition, two candidate obscured TDEs (Jiang et al.
+2023) and the extremely energetic candidate TDE AT2021lwx (Yuan et al.
+2024) were associated with high-energy neutrinos.
+A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources play a crucial rule in producing neutrinos (Li et al.
+2024).
 In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc especially on its significant optical re-brightening that has started years after its initial flare (Veres et al.
 2023).
-We conducted multi-wavelength observations (from radio to X-rays) during the re-brightening and compare it to the initial flare associated with the high-energy neutrino, with the aim to characterize the multi-wavelength properties of this unusual transient event and to classify it.
 The paper is organized as follows.
 We explain the discovery of the transient event and give an insight about its host galaxy in Sect.
-2, we summarize our observing campaign and the data reduction in Sect.
-3, we present the results of the observations in Sect.
+2, summarize our observing campaign and the data reduction in Sect.
+3, present the results of the observations in Sect.
 4 and discuss them in Sect.
@@
 2 The transient event: AT2019aalc
-Studying the infrared properties of AT2019dsg and AT2019fdr have revealed a strong reverberation signal i.e.
-the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Dou et al.
+An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal i.e.
+the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Jiang et al.
+2016; Dou et al.
 2016; Lu et al.
 2016; van Velzen et al.
-2021b, 2024).
-In addition to the exceptionally high infrared luminosities (due to the infrared echo) of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al.
+2021b; Jiang et al.
+2021; van Velzen et al.
+2024).
+In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al.
 2021; Reusch et al.
@@
 2024).
-This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC-191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
+This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
 The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al.
 2013).
-Based on its Sloan Digital Sky Survey (SDSS; Abazajian et al.
+Based on its Sloan Digital Sky Survey (SDSS: Abazajian et al.
 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al.
@@
 2019).
-The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST; Condon et al.
+The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: Condon et al.
 1998).
@@
 1 shows the Pan-STARRS (Chambers et al.
-2016) i-band optical image of the host galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in .
-The host galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
-In addition to the unusually significant infrared echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transient events: a soft X-ray spectrum and transient radio emission (van Velzen et al.
+2016) i-band optical image of the galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in .
+The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
+In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (van Velzen et al.
 2024).
-Moreover, similar to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event.
-Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc, due to its high estimated SMBH mass (, van Velzen et al.
-2024) and low redshift (, Ahn et al.
-2012).
-Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention or being directly monitored at the time when the neutrino was emitted.
+Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event.
+Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, van Velzen et al.
+2024) and low redshift.
+Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention.
 Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event.
@@
 2022; van Velzen et al.
-2024) suggesting a very unusual AGN case.
-The TDE-characteristic evolution of the blackbody light curve including a large and rapid optical flux increase is compatible with a TDE-related scenario.
-However, due to the host galaxy classification, extreme AGN variability provides an alternative explanation for the transient event.
+2024) suggesting a very unusual transient.
 Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al.
@@
 2.
-Although double-peaked optical light curves could be a quasi-universal signature of TDEs around massive () black holes (Wevers et al.
-2019b), such a long double peaked light curve with two comparable peak magnitudes would be very untypical for a regular TDE.
-Throughout this paper we will discuss the results of our observations taken around and after the peak of the second optical flare in the light of different TDE,- and AGN-related scenarios.
 3 Observations and data reduction
@@
 (2023).
-We corrected for extinction with adopting a galactic extinction of (Schlafly & Finkbeiner 2011).
+We corrected for extinction by adopting a galactic extinction of (Schlafly & Finkbeiner 2011).
 As shown in Fig.
@@
 This implies that during these epochs, the flux decreased below the mean flux of the reference images.
-For the -band these were obtained in between March and May, 2018.
+For the -band these were obtained between March and May, 2018.
 While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
 3.1.2 UVOT
-The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972 and 19188, Proposers: Reusch, Veres& Sniegowska) on AT2019aalc from 2023-06-23 to 2024-05-27 with a total exposure time of ks.
+The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks.
 The Ultraviolet and Optical Telescope (UVOT, Roming et al.
@@
 We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package.
-We measured the flux using uvotsource, applying a circular aperature of arcseconds (which capture nearly all the flux of the host galaxy).
-The background was measured using four 6 arcsecond regions in four quadrants away from the host galaxy.
+We measured the flux using uvotsource, applying a circular aperature of arcseconds (which captures nearly all the flux of the host galaxy).
+The background was measured using four 6 arcsecond regions in four quadrants away from the target source.
 To estimate the UV flux before the flare, we use GALEX (Martin et al.
@@
 2016) to extract the flux.
-Over the 4 year baseline the GALEX observations, the NUV magnitudes change by about 0.3 mag.
+Over the 4 year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag.
 Following van Velzen et al.
@@
 From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux.
-Finally, we corrected for galactic dust extinction for each filter.
-The baseline flux in the UVM2 band is 17.2, implying a flux increase of 1.8 mag relative to the peak of the UVM2 light curve.
-This large flux increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux.
+The baseline flux in the UVM2 band is 17.2 mag, implying an increase of 1.8 mag relative to the peak of the UVM2 light curve.
+This large increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux.
 This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.
@@
 The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
-Finally, we corrected for galactic dust extinction.
 We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al.
-2010), which is continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al.
+2010), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al.
 2011), at (W2).
@@
 All selected single-exposure photometry data points are within arcsecond of the nucleus of the host.
-The resulting light curve spans almost years, starting in February until July .
+The resulting light curve spans almost years, starting in February until July 2024.
 3.3 X-ray
-The Swift’s X-ray Telescope (XRT, Burrows et al.
+The X-ray Telescope (XRT, Burrows et al.
 2005) observations were obtained simultaneously with the UVOT observations.
-The observations were performed in photon counting (PC) mode.
-For each of the epochs, we performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way.
-First, cleaned event-files were produced using the xrtpipeline task.
-The exposure map, used to correct for the loss of flux caused by some of the CCD pixels not being used to collect data, was created with the xrtexpomap tool.
-The ximage tool allowed us to determine the source detection level and the source position.
-In case of non-detections, upper limits for the count rate were derived and converted to flux using the WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl tool and assuming the mean power-law index of the observations yielded in detections.
-When our target source was detected above the level, we ran the swxrtdas tool in order to extract the source and the background spectra.
+The observations were performed in photon counting mode.
+We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way.
+First, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks.
+The ximage tool allowed us to determine the source detection level and its position.
+In case of non-detections, upper limits for the count rate were derived and converted to flux using WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming the mean power-law index of the observations that resulted in detections.
+When our target source was detected above the level, we ran swxrtdas in order to extract the source and the background spectra.
 Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius.
-In addition, Ancillary Response Files (ARFs) i.e.
-effective response files for the extracted spectra and Response Matrices Files (RMF), that are used to link the instrumental channel scale with the physical wavelength scale, were automatically created.
-Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra and to create a readable spectrum for spectral fitting tools.
-The relevant ARF and RMF files were applied during this step.
-We rebinned each spectrum to have a minimum of counts per spectral bin.
-Channels below keV were ignored.
-We fitted the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b) in order to be able to study the X-ray flux evolution of AT2019aalc.
-First, we have determined the needed Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al.
+Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra.
+We binned each spectrum to have a minimum of counts per spectral bin.
+We fit the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b).
+First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al.
 2016).
-Then, we loaded the source spectrum and converted the instrumental Channel scale to ’physical’ wavelength scale.
 The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976).
-To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as most of the AGN are dominated by a simple power-law and the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023).
-We performed renormalization and ignored bad data points.
+To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023).
 When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary.
 The resultant fit quality is C-Stat for degrees of freedom.
-We than added a blackbody model component, to take into account the soft excess.
-The blackbody temperature and the normalization of the new component were also free to vary and we fitted again which resulted in a better fit with C-Stat for degrees of freedom.
+We than added a blackbody model component, to take into account a possible soft excess.
+The blackbody temperature and the normalization of the new component were also free to vary and we fit again which resulted in a better fit with C-Stat for degrees of freedom.
 We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component.
-This test runs a given number of simulations of datasets based on the two models provided (which are the simple power-law and the more complex power-law+blackbody models in our case) and calculates the likelihood ratio for the more complex model relative to the simpler one.
-After running simulations, the test resulted only in smaller likelihood differences than the observed one () clearly implying that the difference is unlikely to be due to random fluctuations.
-Therefore, accounting for the soft X-ray excess results in a significantly better fit than a normal power-law fit to model the stacked spectrum.
-The source luminosity of each dataset was determined with adding the clumin component in front of the model components.
-We fixed the normalization in the additive models (power-law and blackbody) and performed the fit again.
-Finally, we derived the confidence level () error of the luminosity and the power-law index and, when it was relevant, the blackbody temperature.
-When fitting the individual datasets, a simple power-law model was enough to adequately describe the data and the soft excess becomes significant only in the stacked spectrum.
-Only one of the individual datasets is better fit with a blackbody model than with a power-law+blackbody or a simple power-law models.
+After running simulations, the test resulted only in smaller likelihood differences than the observed one clearly implying that the difference is unlikely to be due to random fluctuations.
+Therefore, accounting for the soft X-ray excess results in a significantly better fit.
 3.4 Radio
@@
 2004).
-Our very long baseline interferomety (VLBI) maps enable us to study the pc-scale radio structure of AT2019aalc.
 These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres).
-Each of the eight intermediate-frequency channels (IFs) were divided into thirty-two kHz spectral channels, resulting in a total bandwidth of MHz.
-The whole observation time was hr.
+The observation time was hr.
 The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing.
-The following antennas participated in the observations: Cm: Cambridge, UK; Da: Darnhall, UK; De: Defford, UK, Ef - Effelsberg, Germany; Hh - Hartebeesthoek, South Africa; Jb - Jodrell Bank Mk2, UK; Kn - Knockin, UK; MC - Medicina, Italy; Pi - Pickmere, UK; O8 - Onsala, Sweden; T6 - Tianma, China; Tr - Toruń, Poland; Ur - Urumqi, China; Wb - Westerbork, The Netherlands.
 The data were recorded in left and right circular polarizations with a data rate of Gbit s-1.
@@
 4C 04.51) at an angular distance of was observed as phase-reference calibrator.
-Earlier EVN observations of this calibrator source taken at GHz resolved two lobes, with a hint of a jet-like emission (Xiang et al.
-2002).
-In addition, the following radio sources were observed as fringe-finders: J14301043, J15500527 and J17510939.
-The bright CSO J14072827 (a.k.a.
-OQ208) was observed as amplitude calibrator and in order to calibrate the Merlin antennas properly.
 The data were calibrated with the U.S.
-National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures.
-We performed time flagging using the .uvflg ancilliary table.
-The amplitude calibration has been done using the measured system temperatures444Wb, T6, Ur and the e-MERLIN (Cm, Da, De, Kn, Pi) antennas had no or bad system temperature measurements, we used nominal values for amplitude calibration.
-and gain curves of each telescope contained in the .antab ancilliary table.
-Since our observations were conducted at below Ghz, we corrected for the dispersive ionospheric delays.
-Then, we corrected for the parallactic angle and performed global fringe-fitting for the calibrator sources.
+National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources.
 We imaged the phase-calibrator with the Difmap software (Shepherd et al.
-1994), using the hybrid mapping technique with clean deconvolution (Högbom 1974).
-It included several iterations of cleaning and phase self-calibration before we performed an overall amplitude self-calibration to obtain the antenna-specific gain correction factors.
-We then applied these amplitude corrections for telescopes with deviation from the gain solutions in AIPS.
-Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct for the source structure since it contributes to the measured interferometric phases.
+1994), using the hybrid mapping technique (Högbom 1974).
+Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure.
 Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source.
-As a final step, we exported the calibrated visibility file from AIPS to Difmap for imaging purposes described in the following.
-After flagging the outlier data points, we imaged the target source using the hybrid mapping technique including the clean algorithm and phase-only self-calibration.
-Then, we obtained the antenna-specific gain correction factors.
-At the end of the procedure, several steps of amplitude and phase self-calibration were performed, starting from the total observation time, gradually for shorter and shorter solution time intervals.
+We imaged the target source using the hybrid mapping technique in Difmap.
+At the end of the procedure, several steps of amplitude and phase self-calibration were performed.
 A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data.
-This way we can describe the brightness distribution of the radio source for a quantitative analysis and to reveal any possible structure of it.
 Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
 3.4.2 ATCA monitoring
-The Australia Telescope Compact Array (ATCA)555https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc.
+The Australia Telescope Compact Array (ATCA)444https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc.
 The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration.
@@
 The observations are summarized in Table LABEL:table:atca.
-The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad; Sault et al.
+The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: Sault et al.
 1995).
@@
 Due to the uncertainties this method introduced we used a plain % flux error.
-Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz | Flux Calibrator | Phase Calibrator |
-28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
-30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
-08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
-19.07.2023 | 210 | 6D | 2.1 | 1 | 1934-638 | 1548+056 |
-09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
-19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
-09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
-29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
+Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz |
+28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
+30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
+08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 |
+19.07.2023 | 210 | 6D | 2.1 | 1 |
+09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 |
+19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 |
+09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 |
+29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 |
 3.5 Spectroscopic observations
-The Low Resolution Imaging Spectrograph (LRIS; Oke et al.
-1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni), almost exactly at half time between the two optical flares.
-At that time the source was not significantly detected with ZTF as indicated by the forced photometry results.
+The Low Resolution Imaging Spectrograph (LRIS: Oke et al.
+1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni).
 Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal).
-These optical spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al.
+The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al.
 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al.
@@
 2020) for the DBSP.
-We normalized the flux of each spectrum and plotted it versus the rest-frame wavelength.
-Details of these observations of AT2019aalc are summarized in Table LABEL:table:sp.
-The obtained spectra cover a timescale of a year and allow us to study the spectroscopic evolution of the transient during its optical re-brightening.
-We describe the features of each spectra in Subsect.
-4.5 and discuss the remarkable lines in Subsect.
-5.1.
-Date | Telescope | Instrument | Exp. | |
-time | ||||
-(d) | (s) | |||
-Post-flare-1 | ||||
-2021-07-06 | Keck-I | LRIS | 300 | |
-Over-flare-2 | ||||
-2023-06-15 | Keck-I | LRIS | 300 | |
-2023-06-22 | LDT | DeVeny | 800 | |
-2023-07-11 | Palomar Hale m | DBSP | 180 | |
-2023-08-17 | Palomar Hale m | DBSP | 900 | |
-2024-05-21 | Palomar Hale m | DBSP | 1200 |
-To determine the type of AGN, we performed a stellar population synthesis fit of the host galaxy’s SDSS spectrum with (pPXF; Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates.
+Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig.
+17.
+To determine the type of AGN prior the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates (see Fig.
+18).
 Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al.
@@
 However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line.
-For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10).
-In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .666https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
-The SPS fit of the latest DBSP spectrum was conducted in a manner analogous to the one applied to the SDSS.
-However, additional kinematic components were required to account for the internal structure in order to achieve a satisfactory result.
-Each emission line with an excitation potential above is treated as a separate component.
-The [S ii] doublet is fitted as an additional component as well.
+For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10) to model the AGN continuum.
+In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
+The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (Guo et al.
+2018; Shen et al.
+2019; Ren et al.
+2024).
+The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex.
+Emission lines in the H and H complexes were modeled using a combination of broad and narrow Gaussian profiles to represent different kinematic components.
+The high-ionization coronal lines of interest were fitted with additional Gaussian profiles.
+The H complex shows a strong double-peaked feature around .
+This region is known as Bowen Fluorescence complex.
+It was fitted with a set of broad Gaussian components.
+We computed parameter uncertainties using MC resampling.
+Examples of the fitted spectra and the inferred emission line parameters are given in Fig.
+19 and in Table 3, respectively.
 4 Results
@@
 The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig.
-5.
-The transient was first detected by the ZTF in January with a g-band magnitude of mag.
-After the initial detection of the first optical flare in May (), AT2019aalc had brighten to a peak magnitude of () reached on 2019-06-18 on a timescale of roughly days.
-After the quickly and monotonically evolving flare, a slow decay has started, and the source was last detected with ZTF on 2021-02-18 with a magnitude of before its second flare.
-The re-brightening started mid-May , approximately years after the initial flare (Veres et al.
+3.
+The transient was first detected by the ZTF in January with an -band magnitude of mag.
+AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days.
+The second flare started mid-May , approximately years after the initial flare (Veres et al.
 2023).
-The second flare has evolved very similarly to the original one.
-From the beginning of the second flare (the source was detected anew on 2023-05-11 at mag), a mag rise can be seen till the peak of the flare on 2023-07-13 with mag (), on a timescale of only days.
-Both flares evolved with constant color.
-However, the second flare peaked at roughly times the peak luminosity of the first flare.
-The peak of the second flare was followed by a short plateau phase lasted for around days.
-Both flares decay very slowly, on timescales of years.
-The first flare decayed with a power-law index of while the second has been decaying even slower with a power-law index of in -band.
-The decaying phases are not monotonous as bumps are tend to appear.
-Prior to the discovery of AT2019aalc, the host galaxy of the transient has been monitored by several optical surveys.
-To study the optical variability of AT2019aalc prior to the first optical flare in , we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al.
+The second flare evolved very similarly to the original one.
+From the beginning of the second flare, a mag rise was seen up to the peak of the flare on 2023-07-13 with mag (), on a timescale of days.
+The second flare peaked at roughly times the peak luminosity of the first flare.
+The peak of the second flare was followed by a short plateau phase that lasted for around days.
+Both flares declined very slowly, on timescales of years.
+The first flare decayed with a power-law index of while the second with in the -band.
+The decaying phases are not monotonic, as bumps tend to appear.
+During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3).
+Both optical flares brightened with constant color () and showed reddening when decaying.
+Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys.
+To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: Shappee et al.
 2014; Kochanek et al.
-2017) using ATLAS forced photometry server (Tonry et al.
-2018; Smith et al.
-2020) and the Catalina Real-Time Transient Survey (CRTS; Djorgovski et al.
-2011) data.
+2017) and the Catalina Real-Time Transient Survey (CRTS: Djorgovski et al.
+2011).
 In addition, Palomar Transient Factory (PTF Law et al.
 2009) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu.
-No significant optical variability can be seen between April and the discovery of the transient in January by ZTF (see Fig.
+No significant optical variability is visible between April and the discovery of the transient in January by ZTF (see Fig.
 4).
-Both optical flares of AT2019aalc detected by ZTF rising with constant color and show reddening when decaying.
-The second flare has similar color evolution to the first one when evolving but with a color of around the peak.
-The first optical flare was not monitored in UV.
 The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare.
-The UV flare evolved quickly.
-The U-band peak is coincident with the optical peak while the shorter wavelength bands (UM2, UW1 and UW2) peaked weeks later888We note that right after the -band peak, a short plateau phase started in optical and the UV emission in the bands UM2, UW1 and UW2 peak during this plateau.
+The U-band peak is coincident with the optical peak while the shorter wavelength bands peaked weeks later.
+We note that right after the -band peak, a short plateau phase started in optical and UV.
 The transient is most luminous at the shortest wavelength ranges covered by UVOT.
-The UV emission peaks around erg/s in filter UW2.
-After the peak epoch, the UW2U color shows reddening which suggests increasing absorption or a drop of the blackbody temperature.
-Towards the end of the UVOT monitoring, the measured UM2 magnitudes decrease back to the level of the baseline GALEX-NUV magnitudes measured between and .
+The UV emission peaks around erg/s in the filter UW2.
+After the peak, the UW2U color shows reddening.
+Towards the end of the UVOT monitoring, the measured UM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
 Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (Bianchi et al.
@@
 The observed magnitudes are plotted in the top panel of Fig.
-5.
-The host galaxy of AT2019aalc is presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al.
-2003) as 2MASS 152416660451191.
+3.
+We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al.
+2003).
 The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
 The top panel of Fig.
-5 also shows the long-term light curve starting just before the first optical flare.
-There are clear signs of variability throughout the infrared light curve starting in February with the distinct first dust echo flare that peaks around six months after the first optical peak.
-The rise of the dust echo of the second flare becomes visible in the latest WISE data release.
+3 shows the light curve starting just before the first optical flare.
+The distinct first dust echo flare that peaks around six months after the first optical peak.
+The rise of the dust echo of the second flare becomes also visible in the latest WISE data releases.
 The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al.
 2004).
-In contrast, the dust echo presents an increase of mag in W1 and mag in W2.
-The first detection of the dust echo of the second flare shows a brightening of mag in both bands and can be expected to increase further in upcoming data releases.
+In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare.
+Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying.
 Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al.
@@
 2012), the color of AT2019aalc suggests a reliability of only %.
-With the onset of the dust echo, the IR emission becomes dominated by the heated dust, which is shown in Fig.
-6.
-The change from bluer to redder colors come from cooling of the dust.
-The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue.
+With the onset of the dust echo, the IR emission becomes dominated by the heated dust.
+The change from bluer to redder colors comes from cooling of the dust.
+The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue shown in Fig.5.
 Notably, AT2019aalc is not part of the Flaires sample (Necker et al.
@@
 4.3 X-ray
-AT2019aalc was monitored with the Swift/XRT at keV energies for weeks with weekly cadence around the peak of the optical flare and later for additional months with weekly and bi-weekly cadences.
-The luminosity evolution is shown in the bottom panel of Fig.
+The X-ray luminosity evolution is shown in the bottom panel of Fig.
 3 while the count rate variability is plotted in the top panel of Fig.
-7.
-The first two visits show an increase of the X-ray emission.
-This flaring episode peaked at on 2023-06-29 (MJD) (see Fig.
-5.) before it experienced a very quick decreasing.
-This flare peaked weeks before the optical peak.
-The count rate dropped by a factor of , from to , within a timescale of days.
-Later, the X-ray emission shows limited variability for several months, however, less significant flaring episodes appear that peaked in the end of September (MJD) and mid-February (MJD).
-The latter was followed by a bump in the optical and UV light curves around weeks later while no optical/UV data is available around the another X-ray increasing.
-In the end of April , a rapid and extreme flaring episode was detected in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode.
-Interestingly, this X-ray flare was also accompanied by an increasing phase in the optical weeks later, according to the most recent ZTF observations999https://lasair-ztf.lsst.ac.uk/objects/ZTF19aaejtoy/.
-Notably, days passed between both the first and second and the third and fourth X-ray flaring episodes.
-While no data is available for a longer period between the second and third flares when the source was too close to the Sun.
+6.
+The first flaring episode peaked at in the end of June (see Fig.
+3.) This flare peaked weeks before the optical peak.
+The count rate dropped by a factor of within a timescale of days.
+Later, the X-ray emission shows limited variability for several months.
+A hint of increased emission was detected in mid-February which was also followed by a bump in the optical and UV light curves around weeks later.
+In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode.
+Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later.
+A similarly luminous and rapid flare was detected in August and once more, accompanied by an optical bump weeks later.
+Around days passed between the second and third and the third and fourth X-ray peaks.
 In Fig.
-8 we show the source spectrum derived from the full observing campaign.
-The spectrum is dominated by a power-law component but a blackbody component is also significantly presented.
-The spectral index of the stacked spectrum fitted with a power-law+blackbody model is while the blackbody temperature is eV.
-These imply an overall unabsorbed flux of F erg cm-2s-1 corresponding to a luminosity value of L erg/s.
-The count rate of the observations is cts s-1.
+7 we show the source spectrum derived from the observing campaign.
+The spectral index of the stacked spectrum fitted with a power-lawblackbody model is while the blackbody temperature is eV.
+These imply an overall unabsorbed luminosity value of L erg/s.
+The mean count rate of the observations is cts s-1.
 The photon index variability of AT2019aalc is presented in the bottom panel of Fig.
-7.
+6.
 AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al.
-2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half year after the peak of the first optical flare) with a months cadence.
+2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence.
 The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV.
@@
 One observation with XMM-Newton Observatory (Jansen et al.
-2001) has been performed in 2021-02-21 during a period of optical quiescence.
+2001) was performed on 2021-02-21 during a period of optical quiescence.
 The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023).
-These imply that the two X-ray flares observed with the Swift/XRT peaked at much higher luminosities than the plateau or long-term flaring episode between the optical flares, however, there is no X-ray data from around the peak of the first optical flare.
-Furthermore, the eROSITA results confirm the presence of a blackbody component in the X-ray spectrum, which is very soft and dominating the variability of the X-ray source.
 4.4 Radio
 4.4.1 Temporal evolution
-The host galaxy of AT2019aalc was detected in the NRAO VLA Sky Survey (NVSS, White et al.
-1997) in February 1995 and in the FIRST survey in February 2000 (both at 1.4 GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability of the host galaxy prior to the discovery of the transient event.
-Later, the transient’s sky location has been observed during the first and second epochs of the Very Large Array Sky Survey (VLASS).
-The VLASS covers the entire sky visible to the Karl G.
-Jansky Very Large Array (VLA) at GHz wide band offering an angular resolution of arcsec and an rms of mJy beam-1 in each epoch (Lacy et al.
+The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, White et al.
+1997) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability.
+These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz which is typical for Seyfert galaxies (Meurs & Wilson 1984).
+Later, the transient’s sky location was observed in the Karl G.
+Jansky Very Large Array Sky Survey (VLASS, Lacy et al.
 2020).
-One of the main goals of the VLASS multi-epoch survey is to detect various types of extragalactic radio transients, such as TDEs, accretion state changing AGNs and core-collapse supernova (Lacy et al.
-2020; Zhang et al.
-2022).
-To date, the first two epochs of the survey have been fully published.
-Our source has peak flux density values of mJy/beam (on 2019-03-14) and mJy/beam (on 2021-11-06), corresponding to 3 GHz radio luminosities of and in the Epoch 1 and Epoch 2 images, respectively.
-This factor of increase in flux from the first VLASS epoch (three months before the optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al.
+The VLASS observes the entire sky visible to the VLA at GHz wide band three times.
+Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively.
+This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al.
 2024).
-In addition, we found a radio source at the position of AT2019aalc in the first data release (DR1) of the Rapid ASKAP Continuum Survey (RACS; McConnell et al.
+In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: McConnell et al.
 2020).
-The RACS DR1 includes initial observations made for RACS-low (central frequency of MHz)101010https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/racs/.
-Moreover, the source has been observed in two different fields and, consequently, at two different epochs.
+The source was observed in two different fields and, consequently, at two different epochs.
 The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
-Interestingly, the radio luminosity in the beginning of our ATCA monitoring is at the same level within uncertainties ( ) when the source was observed for the second time in the VLASS in April 2021.
+The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021.
 Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau.
-Our multi-frequency monitoring with the ATCA implies no substantial variability on monthly timescales at GHz while it shows slow decreasing at GHz and at GHz over time.
-We observed the radio source at GHz at one epoch.
 We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig.
-5 and focusing on the ATCA monitoring in Fig.
-9.
+3.
 4.4.2 The radio spectrum
@@
 We give an average spectral index of .
-We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities (these flux values are explained in Sect.
-4.4.1.).
+We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities .
 The spectral index we derived between GHz and GHz this way is .
 Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz.
-The spectrum of high frequency peaked radio sources are usually explained with significant absorption taken place in the system due to synchrotron self-absorption (although free-free absorption may also play a role, (e.g., Berton et al.
-2020).
-We further discuss the features of the radio spectrum in Subsect.
-5.3.
 4.4.3 VLBI detection
-The final naturally-weighted VLBI map of AT2019aalc is shown in Fig.
-11.
-We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of with our GHz EVN+e-MERLIN VLBI observation.
-We obtained the following parameters of the fitted component; a total flux density of mJy and a size of mas at a position angle of .
+The naturally-weighted VLBI map of AT2019aalc is presented in Fig.
+9.
+We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of .
+We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of .
 We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999).
@@
 This yields mas implying that the fitted Gaussian component is resolved.
-We note that at high SNR, can be significantly smaller than the size of the resolving beam (Bertero & de Mol 1996; Kovalev et al.
-2005) due to applying a specific a priori hypothesis about the shape of the emitting region which is in our case the two-dimensional Gaussian used to fit the observed brightness distribution of the radio source.
-The positional uncertainty was estimated as follows.
-When calculating the VLBI positional uncertainty, we took into account both the positional uncertainty of the phase calibrator source and its angular separation from our target source.
-The position of the phase calibrator source is listed with an error of mas in the third realization of the International Celestial Reference Frame (ICRF3; Charlot et al.
-2020).
-The angular separation between the target and the calibrator source signifies a positional error of about mas at GHz (Chatterjee et al.
-2004).
-The resolution of the observation depending on the thermal noise of the interferometric images (Reid et al.
-1988) has been also considered.
-Following Reid et al.
-(1988), we estimated a thermal noise-limited positional accuracy of mas.
-The resultant positional uncertainty of mas is therefore dominated by the uncertainty arising from the angular separation between the calibrator and the target source.
-The radio position agrees with the source position in the Gaia DR3 (Gaia Collaboration et al.
-2023) catalog within the positional uncertainties.
 We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Condon et al. 1982; Ulvestad et al. 2005):
@@
 The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015).
-It can be determined by imaging and modeling the structure of a given emitting region.
-The derived value is crucial to characterize the detected radio emission, which can be either thermal or non-thermal.
 4.5 Spectroscopic results
@@
 17 with marking the notable emission lines.
-In the following, we highlight the remarkable line detection of the host spectrum observed before the first optical flare (pre-flare spectrum), the one observed between the two flares (post-flare-1 spectrum) and spectra that were taken during the second flare (over-flare-2 spectra).
+In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and five spectra that were taken during the second flare (over-flare-2 spectra).
 4.5.1 Pre-flare spectrum
-A spectrum of the host galaxy of AT2019aalc was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 13.
+A spectrum of the pre-flare system was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
 We identify strong Balmer emission lines.
-These lines are generally thought to be originating from the broad-line region i.e.
-the vicinity of the SMBH.
-We estimate the FWHM of the Balmer lines to FWHM(Balmer) kms-1 for the broad component of the Balmer lines (see Fig.
-13 subplots) in the host spectrum, which is typical for an unobscured, broad line AGN.
+We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN.
 The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al.
 2019) for our source are consistent with this classification.
-In addition to the Balmer lines, we detect the recombination line He i which is also produced in the BLR.
-The remarkable [O iii] and [O iii] forbidden lines exhibited already in the SDSS spectrum indicate that the host AGN was not completely dormant before the optical flares, which is further supported by the radio detections prior to the first flare (Subsect.
-4.4.1).
-Other forbidden emission lines commonly seen in AGN such as [S ii] , [S ii] and [N ii] are detected as well.
+Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected.
 These lines are produced in the narrow-line region (NLR) in AGN.
-We note that the sulfur lines are also present in the spectra of the TDE AT2019dsg, which was hosted by a quiescent galaxy.
-We detect the [O ii] [O ii] line doublet, which is a potential indicator of star formation or AGN-driven outflows (e.g., Santoro et al.
+We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., Santoro et al.
 2020).
+The pre-flare spectrum also shows the detections of two high-ionization coronal lines ([Fe vii] and [Fe xiv]) which are further discussed below.
 4.5.2 Post-flare-1 spectrum
-The first LRIS spectrum was taken almost years after the first flare and years before the second one.
-We detected an increase in the Balmer lines series and the forbidden lines [O iii] and [O iii] . The Balmer lines are also remarkably broadened.
-In addition to the strong Balmer emission lines, we can identify the He ii He ii and transitions and the Bowen Fluorescence line O iii .
-Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation.
-In the case of Bowen Fluorescence, the He ii–O iii and N iii coincidences lead the mechanism.
-The BF lines occur due to the excitation in certain states of O iii and N iii by the absorption of He ii Lyman photons.
-These excited states lead to a cascade of transitions that can be observed as emission lines in the optical and NUV regimes (Netzer 1990; Trakhtenbrot et al.
-2019).
-Interestingly, the existence of BF emission lines in astronomical objects has been predicted by Netzer et al.
-(1985) as the required gas densities are consistent with what is expected from the extremely dense environment in some AGN and gaseous nebulae.
-We detect the high ionization coronal lines [Fe vii], [Fe x] and [Fe xiv].
-The very high ionozation potentials of eV and eV of [Fe xiv] and [Fe x] , respectively, point to very energetic processes at work.
-First Oke & Sargent (1968) studied coronal lines in detail in a Seyfert galaxy (NGC 4151).
-These lines are commonly detected in AGN and commonly thought to be powered by photoionization of the NLR (Cerqueira-Campos et al.
-2021).
-In some extreme cases, a shock model i.e.
-collosional ionization for the coronal line region is required to explain the origin of these lines as it is the case in the Seyfert galaxy NGC 1068 (Oliva 1997).
+The first LRIS spectrum was taken almost years after the first optical flare and years before the second one.
+In this spectrum, prominent Balmer lines were detected, similarly to the SDSS spectrum. The spectrum exhibits the aforementioned forbidden lines detected in the archival spectrum.
+We identify the Bowen Fluorescence line O iii and He ii .
+The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected.
 We identify the [Ne V] emission lines at and in the spectrum.
-These lines are not strongly excited in the inner regions of AGNs, but are prominent emission lines in the lower density extended emission line regions (EELRs) excited by the AGN via photoionization or shocks.
-The forbidden line [O iii] also appears, however, bending from H is possible.
+Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here.
+These lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks and are sometimes considered as coronal lines (e.g., Cann et al.
+2020).
 4.5.3 Over-flare-2 spectra
-We performed optical spectroscopic observations of AT2019aalc during its second flare.
-These are characterized by a blue continuum with several lines appearing in comparison to the spectra discussed above.
-We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017).
+We performed optical spectroscopic observations of AT2019aalc during the second flare.
+These are characterized by a blue continuum with several emission lines appearing in comparison to the spectra discussed above.
+We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017) ,in Fig.
+10.
 This comparison shows similarities but also significant differences discussed below.
-The line widths and strengths of the earlier detected broad Balmer emission lines (H, H and H) increased and both the line widths and strengths of the Balmer lines peaked days before the optical peak.
-Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
-Over the second optical flare of AT2019aalc we detected another BF line N iii together with the transition line He ii .
-This doublet has not been detected in the archival SDSS spectrum taken years before the discovery of the transient or in the post-flare-I spectrum.
-We detected this doublet already in the first spectrum taken during the second flare.
-The strengths of these lines increased significantly around the second optical flare and peaked on 2023-07-11, very close to the optical continuum peak (2023-07-13).
-As we detected the Bowen line N iii even days after the optical peak of the second flare, we can conclude the long-term persistence of at least one BF line after the second flare as well.
-In addition, the second LRIS spectrum shows a clear increase in the BF line O iii and in He ii when comparing it to the post-flare-1 spectrum.
-Both lines increased by a factor of with respect to the post-flare-1 LRIS spectrum.
-We might detected O iii which is also known as a Bowen line (Selvelli et al.
+Over the second optical flare, strong Balmer lines were detected. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
+Over the second optical flare of AT2019aalc, we detected the Bowen Fluorescence line complex N iii +He ii .
+This complex was not detected in the earlier spectra.
+The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13).
+As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales.
+We also find hints of O iii which is also known as a Bowen line (Selvelli et al.
 2007), however, due to a possible blending from Fe vii, its detection is questionable.
-Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which suggests its detection.
+Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which supports the hint of detection.
 O iii was detected in some metal-rich TDEs before (Leloudas et al.
@@
 We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum.
-Furthermore, a new coronal line (Fe xi, ionization potential: eV) appears on 2023-08-17, around a month after the peak of the second optical flare.
-Interestingly, the line intensities of these lines are larger after the optical peak than around the peak in contrast with the BF lines and the above-mentioned ionized helium lines which experienced their maximums very close to the continuum peak.
-Moreover, [Fe xi and [Fe x] are not just persistent but the strongest in the late-time DBSP spectrum, taken almost a year after the optical peak of the second flare.
-We detect the high-ionization coronal line [S xii only in the late-time spectrum.
-This line is an extreme ionization line with an ionization potential of eV thus implies that unusually high level of ionization is taken place in AT2019aalc almost a year after its second optical peak.
-For this coronal line we estimate FWHM(S xii) km s-1.
-We detected the [Ne V] emission lines at and in the second LRIS spectrum, similarly to the post-flare-1 spectrum and implying continuous ionization of the EELR.
-The [O iii] line appears in the late-time spectrum.
+Furthermore, a new coronal line appears: (Fe xi (ionization potential: eV).
+We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
 5 Discussion
@@
 5.1.1 Balmer lines
-The line width of the Balmer lines and its limited variability are not common for regular TDEs and flaring AGN which tend to have decreasing and increasing Balmer line widths over time, respectively (Gezari 2021).
-However, this behavior is more compatible with the spectroscopic properties of the TDE-AGN PS16dtm (Blanchard et al.
-2017).
-Still, the slight increasing and broadening indicate high-velocity gas moving in the BLR, close to the SMBH.
-In addition, a more complex structure of these lines seen in the late-time DBSP spectrum.
-5.1.2 Bowen lines
-The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii (Selvelli et al.
-2007).
+AT2019aalc exhibits prominent Balmer lines in each spectrum.
+In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum.
+The Balmer series lines here exhibit a redward wing which, however, in some cases, might result from blending with other lines, such as H and the BF line N,iii .
+Interestingly, the post-flare-I spectrum shows similar structure of the Balmer line series.
+5.1.2 Bowen fluorescence lines
+Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation.
+The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (Netzer et al.
+1985; Netzer 1990).
 This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc.
@@
 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al.
-1990), however, we note that due to the atmospheric limit at these lines are not easy to detect.
-This line is typically much weaker in normal Seyfert galaxies as in TDEs, BFFs or in our case (see Fig.
-14).
+1990).
+However, we note that due to the atmospheric cutoff at these lines are difficult to detect.
+This line is typically weaker in regular Seyfert galaxies than in the case of AT2019aalc (see Fig.
+10).
 The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al.
-2001), however, has been observed in several optical TDEs (Charalampopoulos et al.
-2022b) and other SMBH-related transients; the BFFs AT2017bgt, F01004-2237, OGLE17aaj (Trakhtenbrot et al.
-2019) and AT2021loi (Makrygianni et al.
+2001), however, it has been observed in several optical TDEs (Charalampopoulos et al.
+2022b) and other SMBH-related transients: BFFs (Trakhtenbrot et al.
+2019; Makrygianni et al.
 2023), the peculiar transient event AT2019avd (Malyali et al.
@@
 2021).
-In a TDE case, the disrupted star’s material forms an accretion disk and the photosphere absorbs the EUVsoft X-ray photons originating from accretion processes and this way provide a plausible explanation for the appearance of the BF lines.
-BFFs have been first identified by Trakhtenbrot et al.
-(2019) who revealed three flares (AT2017bgt, F01004-2237 and OGLE17aaj) in AGN with the appearance of persistent Bowen lines.
-Later, Makrygianni et al.
-(2023) studied a transient, AT2021loi, classified as a BFF.
-The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the cases of the BFFs studied by Trakhtenbrot et al.
+The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot et al.
 (2019).
-In the case of AT2017bgt, this Bowen line was still detected days after the discovery of the transient.
-These lines in BFFs identified by Trakhtenbrot et al.
-(2019) may originate from a pre-existing BLR (a region with dense line-emitting gas) which was suddenly exposed to the intense ionizing UV emission, which also indicates enhanced accretion onto the SMBH.
-Alternatively, the BF lines might be related to a newly launched outflow driven by a sudden increase in accretion rate which interacts with the BLR.
-We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc with the spectra of the BFFs AT2017bgt and AT2021loi, the BFF candidate AT2019avd, the TDE-Bowen AT2018dyb and the repeating partial TDE AT2020vdq in Fig.
-15 with focusing on the N iii He ii emission line doublet.
+In the case of AT2017bgt, this line was still detected days after the discovery of the transient.
+The Bowen features in F01004-2237 were detected even years after the initial flare (Tadhunter et al.
+2021).
+We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig.
+11 around the N iii He ii complex.
 We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig.
-16 zooming into the O iii and He ii region.
-The increasing of these lines around the optical peak indicate that these lines were enhanced by the Bowen Fluorescence mechanism.
-The BF process is particularly strong in some He ii, O iii and N iii transitions.
-These lines are obviously detected during the second optical flare, while atypical for normal Seyfert 1 galaxies (see a comparison plot in Fig.
-14).
-Moreover, the O iii lines at and normally detected in AGN do not show up in our spectra, similarly to the case of AT2017bgt.
-These results suggest that the BF mechanism plays an important role and this mechanism traces significantly enhanced accretion onto the SMBH suggesting the unusual nuclear transient nature of AT2019aalc.
-The He ii ionized helium line is presented in our spectra taken between the optical flares, and is also increased during the second optical flare.
+12 zooming into the O iii He ii region.
+The He ii ionized helium line is presented in the spectrum taken between the optical flares..
 This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g.
@@
 2021).
-In the spectra of AT2019aalc, He ii is present together with strong N iii and O iii BF lines and the less commonly detected He ii transition, indicating its relation to the Bowen Fluorescence mechanism.
-Furthermore, we calculated a maximum line intensity ratio of whose value significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al.
+In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism.
+Furthermore, we calculated a maximum line intensity ratio of which significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al.
 (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ).
-Our value is, however, more comparable with the line ratio estimated for the BF transient AT2017bgt which is (Trakhtenbrot et al.
+Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot et al.
 2019).
-The line intensities of He ii and N iii peak days before the continuum peak, suggesting extreme ionization by UV/soft X-ray photons around the peak.
-The He ii /He i ratio is sensitive to the change in the temperature and density as these lines belong to the same element but in two different ionization states (e.g., Ilić et al.
-2009).
-We found the highest ratio close to the continuum peak, implying increased temperature in the BLR.
-The smallest ratio can be estimated from the post-flare-1 and late-time spectrum taken days after the continuum peak that is due to a drop in the temperature.
 5.1.3 Coronal lines
-We detected several high-ionization coronal lines.
-Two of the detected coronal lines present already in the post-flare-1 spectrum taken years after the first optical peak indicating a connection to the first optical flare.
 Based on a study of SDSS spectra, Wang et al.
 (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines.
-Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to have coronal lines (Onori et al.
+Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to exhibit coronal lines (Onori et al.
 2022; Newsome et al.
@@
 2023) and the spectacular transient event AT2019avd (Malyali et al.
-2021) have significantly detected coronal lines, too.
-Based on the publicly available spectra of AT2017bgt on the Transient Name Server111111https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
+2021) also exhibit strong coronal lines.
+The BFF F01004-2237 showed variable coronal lines on a timescale of years.
+Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
 Wang et al.
-(2012) states that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later supported by Short et al.
+(2012) state that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later confirmed by Short et al.
 (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN.
-Interestingly, the coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario.
+The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario.
 Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al.
 2012).
+Later, Callow et al.
+(2024) discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations.
+They found that the ECLE rates are below the TDE rates, however, can still be consistent with assuming that only to per cent of all TDEs produce variable coronal lines.
 According to Wang et al.
-(2012), the ECLEs could be explained through the coronal lines arising from Fe liberated from dust grains that were destroyed by the TDE flare.
+(2012), the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the TDE flare.
+A follow-up spectroscopic analysis of the original ECLE sample of Wang et al.
+(2012) revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (Clark et al.
+2024).
 The large dust echo of AT2019aalc is consistent with this picture.
-Furthermore, the TDEs with coronal line detection clearly have more luminous IR dust echo compared to other optical TDEs.
+Furthermore, the TDEs with coronal line detection have more luminous IR dust echo compared to other optical TDEs.
 AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al.
@@
 2024).
-The Bowen line transient AT2019pev that shows strong high-ionization coronal lines (Frederick et al.
+The Bowen line transient AT2019pev shows strong high-ionization coronal lines (Frederick et al.
 2021; Yu et al.
-2022) is also part of the MIR flare sample of Hinkle (2024).
+2022) and is also part of the MIR flare sample of Hinkle (2024).
 These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs.
+We performed modeling of the publicly available LRIS spectrum of AT2017bgt taken on 2017-02-23 using PyQSOFit, following the procedure described in Subsect.
+3.5.
 Fig.
-17 shows the [Fe x] versus the [O iii luminosities of the EECLEs studied by Wang et al.
+13 shows the [Fe x] versus the [O iii luminosities of the ECLEs studied by Wang et al.
 (2012), AT2019qiz (Short et al.
@@
 This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
-The [Fe x] line evolution is shown in Fig.
-18.
-The strength of this coronal line in AT2019aalc is further increased during the second flare and seem to be peaking lately with respect to the continuum peak.
-The coronal line TDE AT2019qiz shows a late-time increase in the this line as well (Short et al.
+5.2 Optical re-brightening
+The long-term ZTF light curve of AT2019aalc shows two well-separated flares.
+Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales.
+Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption.
+AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began years after the initial flare.
+The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al.
 2023).
-Similarly to the case of AT2019qiz, we also found that [Fe x line is the strongest high-ionization coronal line presented in the spectra, while [Fe vii is the weakest one.
-The coronal line [Fe x] has been detected in several AGN before, however, the maximum line intensity ratio we calculated ([Fe x / [O iii] ) clearly overshoots the maximum ratio (, Nagao et al.
-(2000) found for AGN.
-For AT2021loi, Makrygianni et al.
-(2023) calculated even higher ratios, up to .The line width of [Fe x] shows a clear evolution during the monitoring of AT2019aalc as we estimated values of FWHM([Fe x].
-Interestingly, the line widths we estimated for [Fe x are clearly larger than the maximum values found for coronal line emitters (around km s-1, Wang et al.
-2012), however, very consistent with the width estimated for AT2021loi.
-For AT2021loi, Makrygianni et al.
-(2023) gives a line width of km s-1 which implies that the line is consistent with a BLR instead of NLR origin since the latter would require the typical values estimated for coronal line emitters ( km s-1, Wang et al.
-2012).
-The BLR origin would naturally explain the increasing emission in these lines after the optical peak (due to the light travel time) but is also consistent with the distance between the dust responsible for the reprocessed IR emission and the SMBH ( pc, van Velzen et al.
-2024).
-This suggests a connection between the unusually large dust echos and the extreme coronal lines in these transient sources.
-An alternative explanation is collisional excitation of these coronal lines.
-This is compatible with a detailed multi-wavelength study of AT2019qiz which suggests that outflows (powered by accretion or stream collision) are responsible for driving the optical flare of this coronal-line TDE (Nicholl et al.
-2020).
-The sulfur line [S xii has the second highest ionization potential among the optical lines detected in AGN before (the highest one is [Ar xiv which we do not detect in AT2019aalc).
-We zoom into the SPS fitted DBSP spectrum to show the detection of this line in Fig.
-19.
-We can reasonably assume that the most recent and very strong X-ray flare, started less than a month before the appearance of the sulfur line, is responsible for the creation of this emission line.
-Apart from ECLEs in the sample of Wang et al.
-(2012), [S xii] has only been reported in Seyfert galaxies (Kraemer & Crenshaw 2000; Mazzalay et al.
-2010; Cerqueira-Campos et al.
-2021; Oh et al.
-2022) including the neutrino emitter (IceCube Collaboration et al.
-2022) NGC 1068 and in the nuclear transients VT J1548 (Somalwar et al.
-2022) and AT2022fpx (Koljonen et al.
-2024).
-VT J1548 is a radio-detected transient that was classified as a promising BFF candidate and notably with remarkable TDE-like characteristics (Somalwar et al.
-2022).
-Similarly, AT2022fpx shares properties both with TDEs and BFFs (Koljonen et al.
-2024).
-To our knowledge, AT2019aalc is so far only the third nuclear transient detected with a line requiring such a high creation ionization potential.
-5.1.4 Forbidden lines
-We find evidence for the forbidden line [O iii] in the post-flare-1 LRIS spectrum and the late-time DBSP spectrum.
-Its detection, however, remains ambiguous because the Balmer line series are complex in our spectra, including blending from H.
-Nevertheless, the detections of [O iii] years after the first flare and year after the second flare is not unexpected.
-After all, this emission line requires densities much lower than the UV-line emitting clouds, and consistent with the outer NLR (or simple ISM) clouds (Ji et al.
-2024) but is located inner than the emitting region of [O iii ] (Nagao et al.
-2001).
-The line ratio of [O iii] /H suggests extremely hard ionizing radiation from the ionizing source.
-Using James Webb Space Telescope (JWST) observations, this line is sometimes detected and studied in detail in high- AGN (used to distinguish AGN from star-forming regions), however, such a ratio was detected only in a very few cases (Übler et al.
-2024; Ji et al.
-2024; Mazzolari et al.
-2024).
-The [O iii ]/[O iii] ratio of is clearly higher than calculated for AGN which are typically below (Baskin & Laor 2005; Übler et al.
-2024; Binette et al.
-2024)).
-The highly ionized neon lines typically originate from star formation activity or photoionization of the EELR by the AGN.
-In some extreme cases, high-velocity AGN-driven shocks that ionize extended emission regions are expected to be responsible for the presence of the lines [Ne V] at and (Maddox 2018).
-The strengths of these lines in AT2019aalc are higher than the [O ii] [O ii] doublets which indicate their origin related to AGN activity since [Ne V] is not excited by star formation (Maddox 2018).
-These shocks are possibly, but not exclusively, related to radio jets as high shock velocities of km s-1 are required to produce these lines (Maddox 2018).
-We estimated slightly lower velocities, which are therefore more consistent with the EELR origin and significant AGN photoionization in this outer region.
-From the increasing line intensities of [O iii] and [Ne V] we can conclude that the NLR and EELR are also strongly ionized, however, in the late-time spectrum these decreased faster than the Bowen line N iii.
-The ionization level of the BLR is therefore more persistent than in the outer regions.
-5.2 Optical re-brightening
-The long-term ZTF light curve of AT2019aalc shows two well separated flares.
-Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales.
-The first flare decayed for more than a year, while the second flare has been decaying for around a year.
-Several ZTF monitored extragalactic transients have been found with bumps seen in their optical light curves (see Soraisam et al.
-(2022) for a recent review)121212We note that ZTF has not been running long enough to discover flare repeats within longer timescales e.g., a decade that actually might also be common.
-Most of them are SNe, however, a couple of them have been identified as TDEs (AT2020acka, AT2020nov and AT2021ehb).
-The latter two TDEs show prominent bumps in their light curves while decaying, while AT2020acka exhibits a long and monotonic bump.
-These TDEs are, however, experienced these bumps within a year from their peak.
-Moreover, the second flare of AT2019aalc peaked even at higher luminosities than the initial one.
-AT2020nov and AT2020acka experienced less significant second bumps, while AT2021ehb has two comparable peaks.
-The optical emission of these transients are poorly understood, and Soraisam et al.
-(2022) also states that their identifications as TDEs are ambiguous.
-Generally, TDEs are characterized by smoothly and rapidly declining light curves ( as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption.
-AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that has started years after the initial flare.
-The first systemically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al.
-2023).
-This implies an even more significant optical re-brightening than the one experienced by AT2019aalc.
-The second flares of partial TDEs are expected to follow an extremely steep power-law when declining.
+This implies an even more significant optical re-brightening than the one detected in AT2019aalc.
+The second flare of partial TDEs is expected to follow a steep declining with a power-law of .
 In the case of AT2020vdq, the second flare declined much slower (Somalwar et al.
-2023) than this but still clearly faster than the second flare of AT2019aalc which decayed with a power-law index of .
-Other candidate repeating TDEs have been found with periodic flares at different wavelengths; ASASSN 14ko with more than observed optical/UV/X-ray flares (Payne et al.
-2021), eRASSt J045650.3-20375 with X-ray/UV flares (Liu et al.
-2023), AT2022dbl with two optical flares (Lin et al.
-2024), while AT2018fyk (Wevers et al.
-2019a, 2023) and RX J133157.6-324319.7 (Hampel et al.
-2022) showed a one-time X-ray/UV and X-ray re-brightening episodes, respectively.
-In these cases, due to its grazing orbit, the star orbiting in the vicinity of the SMBH is thought to be only partially disrupted.
-After another pericenter passage, the surviving core thus feeds a new accretion flare, resulting in X-ray/UV re-brightening.
-For the Bowen-line and coronal-line emitter transient AT2019avd with two distinct optical flares, Chen et al.
-(2022) discusses a partial TDE scenario.
+2023) than this but still clearly faster than the second flare of AT2019aalc which declined with a power-law index of .
 Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps.
@@
 14 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc.
-The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly seen in the binned light curve published in Trakhtenbrot et al.
+The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot et al.
 (2019).
-One of the two other BFFs originally studied by (Trakhtenbrot et al.
-2019), F01004-2237, experienced a significant optical re-brightening around years after its initial flare detected in .
-Originally, Catalina Sky Survey (CSS) light-curve data revealed that this ULIRG had undergone a luminous continuum flare (Tadhunter et al.
-2021) and later Trakhtenbrot et al.
-(2019) classified it as a BFF similar to AT2017bgt.
-The binned light curve of this source published in Makrygianni et al.
+Originally, F01004-2237 had undergone a luminous flare (Tadhunter et al.
+2021) in 2010 and later Trakhtenbrot et al.
+(2019) classified it as a BFF.
+The binned light curve of F01004-2237 published in Makrygianni et al.
 (2023) shows a bump around a year after its peak.
-We investigated the recent activity of the source and found that it significantly re-brightened in based on its ATLAS monitoring.
-The recent flare was more luminous than the first one and decayed faster.
-In addition to these two BFFs, AT2021loi shows a bump during its decaying, roughly a year after the first flare reached its peak.
+We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring.
+The recent flare was more luminous than the first one and declined faster.
+In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak.
 Makrygianni et al.
@@
 Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date.
-The first flares of the studied BFFs decay on yearly timescales, i.e.
-much slower than TDE flares but more similar to AT2019aalc.
+The first flare of the studied BFFs decay on yearly timescales, much slower than regular TDE flares but more similar to AT2019aalc.
 However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower.
-The BFFs classified so far tend to have remarkable variances in their optical light curve evolution in addition to the similarities we see.
-These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses have been defined based on their light curve evolution (e.g., Charalampopoulos et al.
+The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities we see.
+These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., Charalampopoulos et al.
 2023).
-5.3 Possible explanations of the long-term radio flare
+5.3 Radio characteristics
+5.3.1 Possible explanations of the long-term radio flare
 In this subsection we discuss the long-term radio variability of AT2019aalc.
@@
 Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997).
-Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that the advection-dominated accretion flows cannot explain the observed radio emission in our case.
+Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
 Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc.
-Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG; Green et al.
+Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG: Green et al.
 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993).
@@
 (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies.
-Magnetized coronal winds were proposed to explain the observed radio emission of the Seyfert galaxy Mrk 590 by Koay et al.
-(2016), however, the authors noted that a jet contribution cannot be ruled out.
-These two types of outflows might even be equivalent and cannot be distinguished in Seyfert galaxies with unresolved compact core features seen in their VLBI maps (e.g., Kharb et al.
-2015), like in the case of AT2019aalc.
 For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc.
-Moreover, coronal emission should be rapidly variable at radio wavelengths and should evolve similarly to the X-ray flux, which is clearly not the case.
-Although, a contribution from the mechanisms discussed so far to the total radio flux remains a possibility, we need to look for another possible origin to explain the main origin.
-Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets which can be either on or off-axis relative to the observer.
-These features are often visible on VLBI radio maps of radio-detected Seyfert galaxies (e.g., Kozák et al.
-2024) and a few TDEs (e.g., Paragi et al.
+Moreover, coronal emission should be highly variable at radio wavelengths.
+Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., Paragi et al.
 2017; Mohan et al.
-2022).
-In our case, the high brightness temperature of the EVN-detected radio feature together with its position (in good agreement with the optical Gaia position) indicate its radio core identification, however, no other components were detected.
-Nevertheless, the steep () radio spectrum revealed by our ATCA observations indicates the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum.
-The steep radio spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s).
-The ATCA radio spectrum indicates a high-frequency turnover above GHz which is a possible hint of a newly ejected radio-emitting component.
-We further discuss the radio spectrum of AT2019aalc in the following.
+2022; Kozák et al.
+2024).
+The high brightness temperature of the EVN-detected radio feature indicates its radio core identification.
+The steep () radio spectrum indicate the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum.
+The spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s).
+The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
+5.3.2 The radio spectrum
 The turnover suggests the multiple component nature of the spectrum.
@@
 The high-frequency peak might be explained with the kinematic age of a newly ejected component.
-Since older sources are larger than young ones, the expanding relativistic jet or non-relativistic outflow loses energy as the density decreases and the spectral peak moves to lower energies over time.
-With other words, as the projected linear size of the radio source increases, the turnover frequency decreases.
 These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al.
 2020).
-The low-frequency peak is, however, more likely to be explained with star formation activity (as NLSy1 galaxies have a significantly higher star formation rate as other AGNs) instead of AGN emission.
-The latter is assumed to be synchrotron self-absorbed (SSA) to explain the faintness or the non-detection of these galaxies at lower frequencies.
-As another possibility to explain the high-frequency excess of, Berton et al.
-(2020) discusses a scenario in which the AGN emission is absorbed below GHz due to free-free absorption (FFA).
-In that case, the circumnuclear gas might be ionized both by the AGN and by hot stars and might act as a screen for the jet at low frequencies or the passage of the jet in the interstellar medium causes the formation of a cocoon of ionized gas and this way being responsible for the FFA.
-Finally, coronal emission could also be responsible for the inverted spectrum.
-In the case of AT2019aalc, the coronal emission scenario can be safely excluded as we should see a rapidly variable radio light curve which should evolve similarly to the X-ray emission.
-On the other hand, the other two scenarios might also be challenging.
-The large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA.
-Moreover, the presence of a relatively strong [O iii] line in the SDSS host spectrum implies earlier AGN activity as being responsible for the radio emission before the long-term flare.
+The low-frequency peak is, however, more likely to be explained with star formation activity instead of AGN emission since the latter is absorbed due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
+However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA.
 Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
-Nevertheless, the contribution of a newly ejected component to the overall radio spectrum remains a possibility.
-In the cases of the earlier discussed NLSy1 galaxies, the high frequency peak is thought to be suggesting a newly born radio source with no signs of earlier jet activity.
-Similarly to gigahertz-peaked sources (GPS), the high-frequency peak is thought to be shifting towards lower energies over time.
-This way, NLSy1 galaxies might be extremely young GPS which was first suggested by Oshlack et al.
-(2001).
-In our case, a newly ejected component can be responsible for the high-frequency excess whilst the low-frequency part is naturally explained with optically thin synchrotron emission of the AGN indicating that we do not see a newly born radio-AGN but an AGN with recently enhanced, intermittent activity.
-The result is a two-component radio spectrum with a relatively steep spectrum at lower frequencies and a high-frequency excess.
-The steepness of the spectrum between MHz and GHz might be explained with cooling of electrons originating from past AGN activity, and also a contribution from star formation can play a role here.
+A newly ejected component can explain the high-frequency excess whilst the low-frequency part is originating from optically thin synchrotron emission of the AGN.
+Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior.
 Järvelä et al.
@@
 (2020).
-The broad-line AGN classification of the host galaxy of AT2019aalc is more consistent with a restarted AGN picture as this type of AGN are thought to be older than NLSy1s making a restarted activity scenario over a newly born radio source one more probable.
-The restarted activity results in a peaked spectrum at higher frequencies, superimposed on a spectrum from a period (or periods) of earlier activity peaking at lower frequencies.
-Intermittent activity is further supported by theoretical models suggesting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al.
+Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al.
 2009).
-This is compatible with the accretion-induced lines seen in the post-flare-1 optical spectrum of AT2019aalc (see Subsubsect.
-4.5.2 and van Velzen et al.
-2024).
-We note that the fitted ellipsoid model fit component to the EVN visibility data is aligned to the northwest-southeast direction (clearly differs from the position angle of the restoring beam indicating the angle-dependent resolution of the VLBI array) which potentially indicates a newly ejected radio-emitting component.
-Higher frequency observations with improved angular resolution are required in order to resolve the radio structure we detected at GHz.
-We observed AT2019aalc with the EVN and the Korean VLBI Network (KVN) jointly at GHz in March and expect to reach sub-mas-scale angular resolution, roughly an order of magnitude better than with the currently discussed observation.
-These results will be presented in a follow-up publication
 5.4 X-ray properties
@@
 Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess.
-This soft excess can be well modeled by a blackbody model with a best-fit temperature in the range keV.
+It can be well modeled by a blackbody model with a best-fit temperature in the range keV.
 The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al.
 2012).
-The stacked Swift/XRT X-ray spectrum of AT2019aalc can be adequately described with an absorbed blackbody+power-law model.
-We note that in the cases of strongly beamed jetted sources e.g.
-blazars, the X-ray spectrum can be well modeled with a simple or broken power-law without adding any additional model component (e.g., Kammoun et al.
-2015; Acciari et al.
-2022).
-Although AT2019aalc shows enhanced radio activity, its brightness temperature suggests a non-beamed nature of its radio emission in agreement with the X-ray spectrum.
-The X-ray spectrum is dominated by a power-law component with a soft power-law index of while the thermal component is extremely soft with a blackbody temperature of eV derived from the Swift/XRT observations.
-Most of the BLSy1s typically have a significantly harder X-ray spectrum with a power-law index (e.g., Pons & Watson 2014).
-Our value is more typical for sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al.
+The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV.
+Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., Pons & Watson 2014).
+Our value is more commonly seen in sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al.
 1996; Saxton et al.
 2020).
-The soft X-ray excess appears in the stacked spectrum and becomes more significant over time as we added a blackbody component in order to better describe the X-ray spectrum.
-Furthermore, the data of the peak of the second X-ray flare is clearly better-fit with a single blackbody model than with a power-law one (Subsect.
-4.3) implying not just the need of the blackbody component to adequately fit the data but even its dominance over the power-law component at this epoch.
-A softening of the power-law index over time can be seen as well.
-These results suggest significantly enhanced accretion or disk instability.
-Delayed accretion flares have been observed in TDEs (Saxton et al.
-2020).
-In our case, the post-flare-1 spectrum shows clear signs of accretion and the second flare is followed by enhanced accretion activity as well.
 The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x].
@@
 2009).
-For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability we consider only observations taken on the same day).
-This implies that the expected soft X-ray luminosity is consistent with the observed values.
-The soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as it was suggested for ECLEs by Wang et al.
+For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day).
+Thus, the soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as suggested for ECLEs by Wang et al.
 (2012).
-The photoionization or collisional excitation of the dust grains is rather due to a current event.
-This further supports that the reoccurring optical/UV outbursts seen during the decaying of the flares are driven by the soft X-ray flares which lead these outbursts only with a couple of weeks.
-These suggest an accretion-driven origin of the reoccurring outbursts in the optical/UV during the decaying of the flare.
+The reoccurring optical/UV bumps seen during the decaying of the main optical flares are likely driven by the soft X-ray flares which lead these outbursts only with a couple of weeks.
+The soft X-ray excess appears to be more dominant during these flares.
+Interestingly, the soft X-ray flares show hints of quasi periodicity of days.
+A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit with soft X-ray flares on time-scales of hours to days while partial TDEs on timescales of months to years (e.g., Chakraborty et al.
+2021; Quintin et al.
+2023; Miniutti et al.
+2023; Nicholl et al.
+2024).
+We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring.
+This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: Evans et al.
+2023) as LSXPS J031335.6020905.
+The potential connection between BFFs and QPEs should be explored in future studies.
 5.5 SED fitting
@@
 These fitted curves to the optical/UV (’blue’ blackbody) and IR datapoints (’red’ blackbody) are shown in Fig.
-21, together with a combined double blackbody fit.
+15, together with a combined double blackbody fit.
 The red blackbody fit was introduced to account for the IR dust echo.
 These curves fit the SED around the optical peak of the investigated flare.
-The continuum emission of TDEs are well described by a thermal blackbody model (Gezari 2021).
+The continuum emission of TDEs is well described by a thermal blackbody model (Gezari 2021).
 The case of AT2019aalc appears to be more complex.
-An excess in the UW2 band can be seen regardless the consideration of a simple or the double blackbody fit.
-The UV emission have a contribution from the Bowen Fluorescence mechanism and potentially also from Comptonization explaining the poor fit towards the blue.
-The red wing of the optical and the IR part cannot be well fitted with a simple blackbody model suggesting a complex, multi-component dusty reprocessing region.
-The excess in the -band might be explained with the early appearance of some Fe lines in this range (such as [Fe x] and [Fe vii]) indicating the release of this element previously locked up by the dust grains.
-Since all of this seem to be started already around the optical peak, at least a fraction of the dust is located on subpc scales to the SMBH.
-This is further consistent with the extreme line widths of the coronal lines (see Subsect.
-5.1.3).
-Therefore, not only the dusty tori is responsible for the IR dust echo.
-It is possible that the significant amount of dust in the nucleus region is the consequence of enhanced star formation sometime in the past.
-The Baldwin, Phillips& Terlevich (BPT, Baldwin et al.
-1981) emission-line ratio diagnostics of the host SDSS spectrum are log([O iii] and log([N ii] .
-We plotted the BPT line-ratio of SDSS spectra with indicating the ratios of the host galaxy of AT2019aalc and the BFFs AT2021loi, AT2017bgt and F01004-2237 in Fig.
-22 using astroml (Vanderplas et al.
-2012; Ivezić et al.
-2014).
-All of the investigated galaxies are in the composite region, consequently, in addition to AGN activity, star formation activity might also have a remarkable contribution for producing the narrow lines seen in the spectra.
-Moreover, these ratios place these two galaxies solidly in the region of the diagram where the post-starburst sample of French et al.
-(2015) is located.
-We note that to derive the ratios for AT2019aalc and F01004-2237, we used pre-flare spectra.
-While as no host spectra are available for AT2017bgt and AT2021loi, we considered the spectra taken as early as possible in these cases.
+The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization explaining the poor fit towards the blue.
+The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations as the IR light curve is peaked well after the optical one.
+The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range but such a significant excess cannot be fully explained with the coronal lines only.
+The inferred blackbody temperature and radius values estimated for the ’blue’ blackbody fit are K and cm, respectively.
+The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by van Velzen et al.
+(2020) and Thomsen et al.
+(2022).
+On the other hand, the radius is smaller than calculated for optical selected TDEs but is more consistent with the X-ray detected ones.
+The ’red’ blackbody fit resulted in K and cm.
 5.6 What type of transient is AT2019aalc?
@@
 Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case.
-The most remarkable difference is, however, the decay of the optical emission, which is significantly slower than for any TDE identified so far.
-The color evolution of the flares is also atypical for TDEs.
+The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far.
+The reddening in color when the main optical flares are declining is also atypical for TDEs.
 The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc.
-Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of strong high-ionization coronal lines.
-The mechanism that powers the long-term accretion enhancement in BFFs is still unclear.
+Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
+The mechanism that powers the BFFs is still unclear.
 TDEs in AGN might play a role.
-In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al.
-2017) or the AGN’s radiation field.
-The bumps in optical and UV during the decay of the flares of AT2019aalc led by soft X-ray flares indicate the flare’s accretion-driven origin consistently within this picture.
-Moreover, the flaring episodes in X-rays seem to be happen periodically which is expected from interaction between a circularizing stellar stream and the accretion disk.
-The origin through reprocessing is further consistent with the ionization of the BLR and so on the increase in the Bowen and the Balmer lines seen in the spectra during the second flare.
-The dusty region (expected from a source with IR dust echo) and the liberation of Fe and S from these dust grains due to photoionization by the soft X-ray flares provide an explanation for the high-ionization coronal lines.
-The increase of the W1W2 color implies significant cooling expected from cooling dust after the onset of the dust echo.
-This scenario provides explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares.
-The second flare is even more luminous than the first one, similarly to the case of the partial TDE AT2020vdq (Somalwar et al.
-2023).
-A partial TDE scenario might explain the second optical flare.
+In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al.
+2017, and ASASSN-14ko, Huang et al.
+2023) or the AGN’s radiation field.
+According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (Chan et al.
+2019).
+The debris-disk interaction results in enhanced accretion and multiple shocks (Chan et al.
+2019), notably leading to X-ray emission (Ryu et al.
+2024).
+Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk.
+These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare.
+The soft X-ray emission is likely reprocessed in optical/UV by the dense environment explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines.
+The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares.
+TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (Chan et al.
+2019; Ryu et al.
+2024).
+Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (Somalwar et al.
+2023) and the BFF F01004-2237.
+Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (Zhong 2024; Liu et al.
+2024).
+Liu et al.
+(2024) claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq.
+A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2017bgt and AT2021loi) with double-peaked optical flares as well.
 Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well.
-The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understood their unique optical light curves.
+The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understand their unique optical light curves.
 Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs.
 This would explain the TDE-like properties of these sources.
-The most remarkable differences i.e.
-the slowly decaying optical emission and the continuum emission not being well described by blackbody models might appear due to a more complex environment the TDE takes place with comparing to quiescent galaxies.
+The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”).
+TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (Chan et al.
+2019).
 Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al.
-2012) and the BFFs is consistent with this picture.
-A known potential partial TDE in an AGN is the case of the active galaxy IC 3599.
+2012) and the BFFs (see Fig.13) is consistent with this picture.
+A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599.
 This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al.
@@
 Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig.
-17).
-Interestingly, IC 3599 is the only known AGN with fading coronal lines (Frederick et al.
+13).
+In addition, IC 3599 is the only known AGN with fading coronal lines (Frederick et al.
 2019).
-Alternatively, a double TDE i.e.
-the tidal disruptions of a pair of binary stars might explain the double-peaked optical light curve.
-Notably, double TDEs may represent nearly of all stellar tidal disruptions (Mandel & Levin 2015).
-The strength of the high-ionozation coronal lines that regular AGN activity cannot explain the unusual behaviour of this transient.
-It is not clear why we do not see as strong high-ionozation coronal lines in normal AGNs as in the transients with significant dust echo emission i.e.
-TDEs and BFFs.
-It might be explained with dust sublimation due to the UV radiation of the AGN (the sublimation radius is typically in the subpc range, van Velzen et al.
-2021b) which clears the dust from the nucleus region.
-AGN-driven outflows play also a role here.
-The limited variability of the host galaxy of AT2019aalc prior the first flare and its location in the BPT diagram suggest weak AGN activity before the transient occurred.
-Therefore, the dust possibly originating from star formation activity in the nucleus could survive.
-Additionally, the first flare could not clear the nucleus region which suggests that it must have been way less luminous in UV than the second flare.
-Unfortunately, due to lack of UV data around the first flare this hypothesis cannot be further tested.
-Nevertheless, the detection of some Fe lines in the post-flare-1 spectrum implies that a certain fraction of the dust was destroyed following the first flare.
-Interestingly, the host AGN of other BFFs also showed weaker pre-transient activity than persistently accreting AGN (Trakhtenbrot et al.
-2019; Makrygianni et al.
-2023).
-The TDE-like features suggest that TDEs might restarted these AGN.
-Beyond the TDE-related scenarios, it is possible that the restarted AGN activity is rather explained with the dusty environment itself.
-As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs available.
+5.7 Radio properties of BFFs
+As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs.
 Based on the radio observations explained in Makrygianni et al.
-(2023) AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz which can be compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc.
+(2023), AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz.
 Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability.
 We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz.
-The host galaxy of AT2017bgt was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy.
-In addition, the BFF candidate VT J1548 was also found to exhibit a radio spectrum peaking around GHz (in contrast with non-jetted AGN that peak below GHz, e.g., Padovani et al.
-2017) which was explained with an ejection of a synchrotron-emitting outflow (Somalwar et al.
-2022).
-These findings allow us to conclude that newly ejected radio outflow/jet components play a crucial role in BFFs.
-These outflows ejected due to enhanced accretion might be partly responsible for the liberation of Fe and S from the dust grains located relatively close to the SMBH.
-The collisional ionization in the case of AT2019aalc is supported by the extreme broadness of the coronal lines especially [S xii.
+The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy.
+The BFF AT2020afhd was detected with the VLA at GHz with mJy (Christy et al.
+2024).
+If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations published in Chan et al.
+(2019).
+Radio observations of any newly discovered BFF will be essential to better understand these transients.
 5.8 BFFs and high-energy neutrinos
-AT2019aalc has been associated with the high-energy neutrino event IC-191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al.
+AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al.
 2024).
 Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect.
-5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1; Abbasi et al.
+5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: Abbasi et al.
 2023).
-We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC-230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023).
+We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023).
 In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour.
-We present the source positions together with the neutrino best-fit positions and both contours in Figs.
-24 and 24.
+We present the source positions together with the neutrino best-fit positions and both contours in Fig 16.
 We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A.
 This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
-The neutrino associated with AT2021loi is days delayed with respect to its first optical peak and days with respect to its second optical peak.
-Note, that the nuclear transients (including AT2019aalc) are associated with neutrinos that were delayed by days, which might be explained with delayed mass accretion, delayed jet formation or outflow ejection (van Velzen et al.
+The neutrino associated with AT2021loi was detected days and days after the first and second optical peaks, respectively.
+Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks.
+This might be explained with delayed outflow activity (van Velzen et al.
 2024).
 In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al.
-2023) is compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc.
-However, from the current data, we cannot tell if the radio-emitting region of AT2019aalc has been ejected before or after the neutrino.
-Although we do not see a clear increase in the MHz RACS data months after the neutrino arrival.
-However, if the radio flare is indeed related to a newly ejected, synchrotron self-absorbed component, it is not surprising that it does not result in any increasing at low frequencies at early times.
+2023) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc.
 Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al.
@@
 These findings indicate that outflows play an important role in the neutrino production.
-Moreover, Winter & Lunardini (2023) estimate the highest neutrino fluence for AT2019aalc among the studied nuclear transients associated with neutrinos.
-A similar case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
-If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources.
+A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
+If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources.
 Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al.
-2024; Hinkle 2024; see also Subsect.
+2024; Hinkle 2024: see also Subsect.
 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al.
-2023; see also Subsect.
+2023: see also Subsect.
 5.1).
 6 Conclusions
-We conducted a multi-wavelength campaign during the second optical flare of the nuclear transient AT2019aalc associated with a high-energy neutrino event.
-We found that the second optical flare is even more luminous and decays even slower than the first one.
+We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc associated with a high-energy neutrino event.
+We found that the second optical flare is more luminous and decays slower than the first one.
 Both flares evolve rapidly and with constant color.
-The optical flare is accompanied by UV emission which peaks at extremely high luminosities (at erg/s).
-The IR emission increases, similar to the first flare when it peaked delayed with respect to the optical and was explained with a dust echo flare.
-We detected a very soft X-ray flare that peaked before the optical/UV emission peaks.
-Later, the Swift/XRT detected an almost times more luminous X-ray flare which is clearly dominated by a blackbody component, suggesting enhanced accretion or disk instability.
-Our VLBI observation provides a hint of extension of the radio-emitting source and clearly favors an AGN-like origin over a star-formation scenario.
-The radio monitoring revealed a long-term flare and an inverted spectrum above GHz, which is most probably explained by a newly ejected, self synchrotron-absorbed outflow or jet component.
-As we do not see obvious signs of variability well before the first optical flare, we suggest that the transient event in renewed the activity of the host AGN.
-Restarted AGN tend to have inverted radio spectrum.
-This also naturally explains the long-term flare.
-The results from optical spectroscopy help us to better understand the event responsible for restarting the AGN.
-In the optical spectra, the presence of persistent Bowen lines are consistent with intensified accretion over a year after the second flare.
-The UV bright nature of the transient is consistent with the Bowen mechanism which pumps lines mostly in this range.
-The line intensities of the Bowen lines and the temperature of the BLR both peak around the peak of the optical flare indicating that the optical flare is driven by the ionization of the clouds of most probably the BLR.
-As we find the X-ray flaring episodes prior to the rise in optical/UV, the origin of the latter is most probably reprocessed X-ray photons which are also required for the extreme ionization implied by the Bowen lines.
+The UV emission peaks at extremely high luminosities (at erg/s).
+The re-brightening is accompanied by an IR flare, similarly to the first flare when it was explained with dust echo emission.
+An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature.
 The X-ray flares are very soft which explains the unusually strong He ii lines.
-These findings altogether suggest the accretion-driven origin of the flaring episode we monitored.
-Note, that reprocessed soft X-ray emission i.e.
-by an outflow surrounding the disk is a possible explanation for the origin of the optical/UV emission of TDEs (Bu et al.
-2022).
-We detected several persistent or even late-time increasing high-ionization coronal lines in the spectra of AT2019aalc.
-The detection of [S xii requires an extremely high ionization potential.
-To our knowledge, AT2019aalc is only the third extragalactic transient event detected with this line and only Seyfert galaxies (including the neutrino emitter NGC 1068) and ECLEs were reported with a line with such a high ionization potential.
-The Fe/X-ray ratios disfavor a past TDE interpretation responsible for the coronal lines, which was proposed to explain these lines seen in most of the ECLEs studied earlier.
-The slowly decaying optical emission might instead be explained with continuous photoionization of the dust grains (located at distances consistent with the BLR instead of the NLR as seen for ECLEs before).
-This is consistent with the early observation of the dust echo emission with respect to the optical peak which suggests that the dusty region is located very close to the SMBH.
-Based on the narrow-line ratios, we found that enhanced star formation activity is a favorable explanation for the dusty environment.
-Moreover, the gas-rich environment might explain the restarted activity of the AGN.
-The high [Fe x / [O iii luminosity ratios of the BFFs, coronal line-detected TDEs and ECLEs suggest a possible connection between these source types.
-These sources tend to have large dust echo-like flares, which implies that the liberation of Fe from the dust grains can be explained via their TDE(-like) flaring episodes.
-The significant photoionization is supported by the detection of strongly ionized neon lines in the EELR.
-A delayed outflow/jet component and collisional excitation provides an alternative explanation for the coronal lines and their late-time increasing or appearance.
-The BFF and coronal-line emitter AT2021loi has an inverted radio spectrum and based on survey data we found a flaring radio source at the position of AT2017bgt.
-These findings suggest that newly ejected radio-emitting components play an important role not only in BFFs, but might be connected to the coronal line emission.
-A detailed radio study of other BFFs and coronal line emitters is crucial to better understand the appearance of these highly ionized lines.
-Altogether, our results are consistent with the Bowen Fluorescence Flare classification of AT2019aalc.
+In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare.
+With VLBI we clearly detect a non-thermal radio source which favors an AGN-like origin.
+The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component.
+Altogether, our results suggest a Bowen Fluorescence Flare classification of AT2019aalc.
 We also classify AT2019aalc as an Extreme Coronal Line Emitter.
-A connection between the ECLEs, some TDEs and BFFs suggested by our spectroscopy results makes a non-standard TDE-related scenario more probable to explain the optical flares and i.e.
-the restarted activity of the host AGN.
-Moreover, due to the pre-existing BLR and therefore the high densities, TDE-AGN are more capable for the Bowen Flurosence mechanism and the high-ionization coronal lines.
-Only a few TDEs were detected with coronal lines; these are also known as TDEs with extreme dust echo emission.
-Continuous interactions of a pair of binary star and the SMBH favored by Trakhtenbrot et al.
-(2019) might explain the re-brightened and slowly decaying optical emission of BFFs.
-An alternative scenario for the two distinct optical flares is a partial disruption of a single star in an AGN.
-Here, the slowly decaying reprocessed optical/UV emission originates from soft X-ray flares powered by continuous tidal interaction between the stellar debris and the pre-existing accretion disk.
-The double-peaked optical light curve is explained with a second disruption due to the surviving core after the first passage.
-However, more BFFs are needed to study the mechanism responsible for these transients.
-TDE-related scenarios nevertheless provide an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the ionized gas.
-Future observations will help us to better understand the nature of the transient; a binary TDE would not produce a third flare, a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
+The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment.
+The stellar debris-disk interaction in such a system could result in complex light curve variations e.g., the plateau observed in optical/UV (Chan et al.
+2020).
+A TDE-related scenario provides an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the disk gas which was notably suggested by TDE-AGN simulations (Ryu et al.
+2024).
+Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
 Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs.
-We found that AT2021loi is coincident with the high-energy neutrino event IC-230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al.
+We found that AT2021loi is coincident with the high-energy neutrino event IC230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al.
 (2024).
-Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionozation coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode/bumps in optical.
-In both cases the host galaxies are classified as broad-line AGN with remarkable star formation activity.
-Acknowledgements
+Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode in optical.
+In both cases the AGN are classified as broad-line AGN in-prior the transients occurred.
+Acknowledgements.
 We thank Benny Trakhtenbrot for useful discussions.
 PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal.
-EH acknowledges support by NASA under award number 80GSFC21M0002.
-The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes.
+EH acknowledges support by NASA under award number 80GSFC21M0002.The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes.
 Scientific results from data presented in this publication are derived from the following EVN project code: EV027.
@@
 This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy.
-The Australia Telescope Compact Array is part of the Australia Telescope National Facility https://ror.org/05qajvd42 which is funded by the Australian Government for operation as a National Facility managed by CSIRO.
+The Australia Telescope Compact Array is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO.
 We acknowledge the Gomeroi people as the Traditional Owners of the Observatory site.
@@
 Operations are conducted by COO, IPAC, and UW.
+Appendix A Spetroscopic details
+Date | Telescope | Instrument | Exp. time | Slit width | Seeing FWHM |
+(s) | (arcsec) | (arcsec) | |||
+Post-flare-1 | |||||
+2021-07-06 | Keck-I | LRIS | 300 | 1.0 | 1.1 |
+Over-flare-2 | |||||
+2023-06-15 | Keck-I | LRIS | 300 | 1.0 | 1.4 |
+2023-06-22 | LDT | DeVeny | 800 | 1.5 | 1.2 |
+2023-07-11 | Palomar Hale m | DBSP | 180 | 1.5 | 1.5 |
+2023-08-17 | Palomar Hale m | DBSP | 900 | 1.5 | 1.1 |
+2024-05-21 | Palomar Hale m | DBSP | 1200 | 1.5 | 1.3 |
+Date | Instrument | Resolving power | |||
+(approx.) | (days) | ||||
+2008-04-05 | SDSS | 1500-2500 | -11 years | 2a,b | |
+2021-06-07 | LRIS | 1000 | -737 | 0.21 0.01 | 4a,b,c,d |
+2023-06-15 | LRIS | 1000 | -28 | 0.64 0.02 | 4a,b,c,d |
+2023-06-22 | Deveny | 750 | -21 | 0.50 0.01 | 4a,b,c,d |
+2023-07-11 | DBSP | 1000 | -2 | 0.83 0.04 | 4a,b,c,d |
+2023-08-17 | DBSP | 1000 | +35 | 0.25 0.01 | 5a,b,c,d,e |
+2024-05-21 | DBSP | 1000 | +315 | 0.30 0.02 | 5a,b,c,d,e |

### v3

@@
 Interestingly, the source has re-brightened 4 years after its discovery.
-Aims. We aim to classify the transient and explain the mechanism responsible for its second optical flare.
+Aims. Our aim is to constrain the physical mechanism responsible for the second optical flare, which may also provide clues to the origin of the initial event.
 Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023/2024.
 Results.
-The observations revealed a uniquely bright UV counterpart and multiple X-ray flares during the second optical flaring episode of the transient.
+The observations revealed multiple X-ray flares during the second optical flaring episode of the transient and a uniquely bright UV counterpart.
 The second flare, similarly to the first one, is also accompanied by IR dust echo emission.
@@
 We found that, in addition to AT2019aalc, another BFF AT2021loi is spatially coincident with a high-energy neutrino event.
-The multi-wavelength properties of these transients suggest a possible connection between ECLEs, BFFs and TDEs in AGN.
+We propose a repeating TDE scenario within an AGN framework to explain the multi-wavelength properties of AT2019aalc and suggest a possible connection among ECLEs, BFFs, and TDEs occurring in AGNs.
 Key Words.:
 galaxies: active – galaxies: Seyfert – quasars: emission lines – individual: AT2019aalc – Neutrinos1 Introduction
-Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988).
+Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (1988Natur.333..523R).
 About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum.
-TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying emission.
-Several TDEs were detected in X-rays with a very soft spectrum typically explained by late-time accretion disk formation (Saxton et al.
-2020).
+TDE optical light curves typically exhibit a rapid, colorless rise followed by a fast decay on monthly timescales.
+Several TDEs were detected in X-rays with soft spectra typically explained by late-time accretion disk formation (2020SSRv..216...85S).
 In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate.
-non-relativistic outflows (Horesh et al.
-2021; Cendes et al.
-2023), or only in instances by newly launched relativistic on-axis jets (Berger et al.
-2012; Brown et al.
-2015; Cenko et al.
-2012; De Colle & Lu 2020; Yao et al.
-2023a).
+non-relativistic outflows (delay1; delay2), or only in four instances by newly launched relativistic on-axis jets (2012ApJ...748...36B; 2015MNRAS.452.4297B; 2012ApJ...753...77C; jetted_tdes; 2022cmc).
 Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H.
-In a few cases, Bowen Fluorescence (BF: Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al.
-2019; Charalampopoulos et al.
-2022a).
-TDEs generally result in one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare.
+In a few cases, Bowen Fluorescence (BF: 1934PASP...46..146B; 1935ApJ....81....1B) lines e.g., N iii at appear (e.g., 2019ApJ...873...92B; spectra2).
+TDEs generally produce one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare.
 These might be explained in the following way.
-Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH.
+A star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH.
 The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening.
+This scenario is naturally explained by the Hills capture mechanism (1988Natur.331..687H), where a binary star system is disrupted by the SMBH, capturing one star on a highly eccentric orbit prone to repeated partial disruptions.
+In this process, the captured star’s orbit brings it close enough to the SMBH multiple times, causing episodic stripping of stellar material and producing recurring flares over timescales of months to years.
 Only a few partial TDEs or candidates are known to date.
-These are AT2020vdq (Somalwar et al.
-2023), ASASSN 14ko (Payne et al.
-2021), eRASSt J045650.3-203750 (Liu et al.
-2023), RX J133157.6-324319.7 (Hampel et al.
-2022), AT2018fyk (Wevers et al.
-2019, 2023), AT2019avd (Chen et al.
-2022), AT2022dbl (Lin et al.
-2024) and AT2021aeuk (Sun et al.
-2025).
-Based on the observed properties of TDE host galaxies, French et al.
-(2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al.
-2023b).
-Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al.
-2019; Ryu et al.
-2024).
-These transients are challenging to distinguish from regular AGN variability.
-Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (Blanchard et al.
-2017) or the multiple soft X-ray flares in IC 3599 (Campana et al.
-2015) and GSN 069 (Shu et al.
-2018)).
+These are AT2020vdq (2020vdq), ASASSN14ko (2021ApJ...910..125P; 14ko2; 14ko3; 2023ApJ...956L..46H), eRASSt J045650.3203750 (2023A&A...669A..75L; 2024A&A...683L..13L), RX J133157.6-324319.7 (2022RAA....22e5004H), AT2018fyk (fyk1; fyk2; 2024ApJ...970..116W; 2024ApJ...971L..31P), AT2019avd (avd1; 2021ApJ...920...56F; avd2; avd3; avd4), AT2019azh (2021MNRAS.500.1673H), AT2022dbl (2024ApJ...971L..26L; 2024arXiv241215326H; 2025ApJ...987L..20M) and AT2021aeuk (at2021aeuk_jingbo).
+Based on the observed properties of TDE host galaxies, 2014ApJ...792...53V estimated a TDE rate of a few per year per galaxy, which is well borne out by recent data (2023ApJ...955L...6Y) and earlier observational estimates (1999MNRAS.306...35S).
+Importantly, a similar fraction of TDEs should occur in AGN hosts (e.g., tdes_in_agn; 2024MNRAS.527.8103R).
+These transients are challenging to distinguish from regular AGN variability and only a few candidate TDEs occurring in AGN (in the following: TDE-AGN) have been studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (2017ApJ...843..106B) or the multiple soft X-ray flares in IC 3599 (2015A&A...581A..17C) and GSN 069 (2018ApJ...857L..16S)).
 In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
-The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al.
-2021).
-In addition to TDEs, flaring NLSy1s, changing-look AGN (CLAGN) and different types of SNe were classified as typical nuclear transients.
-Recently, Trakhtenbrot et al.
-(2019) identified a new class of spectroscopically unique flares from accreting SMBHs.
-In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features.
-Later, the peculiar X-ray transient AT2019avd (Trakhtenbrot et al.
-2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear.
-A more promising case is AT2021loi (Makrygianni et al.
-2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al.
-(2019).
-These events show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission.
-These AGN are characterized by significantly and longer-term intensified accretion.
-Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al.
-(2023).
-Using the Zwicky Transient Facility (ZTF, Masci et al.
-2019) Public Survey data, Dgany et al.
-(2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare.
-Two other candidates, namely AT2022fpx, (Koljonen et al.
-2024) and AT2020afhd (Arcavi et al.
-2024) were found later.
-In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al.
-2023).
-The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Stein et al.
-2021) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Reusch et al.
-2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos.
-Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (van Velzen et al.
-2024) led to the identification of the transient event AT2019aalc (a.k.a.
+It is worth noting that transients classified as Ambiguous Nuclear Transients (ANTs; e.g., 2025MNRAS.537.2024W), as well as several of the events studied by 2021ApJ...920...56F, could also be TDEs in AGN hosts, potentially increasing their overall number.
+Recently, Trakhtenbrot identified a new class of spectroscopically unique flares from accreting SMBHs.
+In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj were identified as the first members of the class based on their unique optical spectroscopic features.
+Later, the peculiar X-ray transient AT2019avd (avd1; 2021ApJ...920...56F; avd2; avd3; avd4) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear.
+A more promising case is AT2021loi (AT2021loi), a UV-bright transient event with very similar multi-wavelength properties to the three flares studied by Trakhtenbrot.
+These events show unobscured AGN-like optical spectra with significant and persistent BF lines, high UV-to-optical ratio and optical emission that decays on yearly timescales.
+These AGN are characterized by accretion that intensifies significantly over long timescales.
+Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following AT2021loi.
+Using the Zwicky Transient Facility (ZTF, 2019PASP..131a8003M) Public Survey data, 2023ApJ...957...57D found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare.
+Two other candidates, namely AT2022fpx, (2024MNRAS.532..112K) and AT2020afhd (2024TNSAN..53....1A) were found later.
+In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (2023MNRAS.521.5046S).
+The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Bran) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Tywin) points to TDEs as a potential new class of sources of the extragalactic neutrinos.
+Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (Lancel) led to the identification of the transient event AT2019aalc (a.k.a.
 ZTF19aaejtoy) with the neutrino alert IC191119A.
-In addition, two candidate obscured TDEs (Jiang et al.
-2023) and the extremely energetic candidate TDE AT2021lwx (Yuan et al.
-2024) were associated with high-energy neutrinos.
-A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources play a crucial rule in producing neutrinos (Li et al.
-2024).
-In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc especially on its significant optical re-brightening that has started years after its initial flare (Veres et al.
-2023).
+In addition, two candidate obscured TDEs (hidden) and the extremely energetic candidate TDE AT2021lwx (lwx) were associated with high-energy neutrinos.
+A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources plays a crucial rule in producing neutrinos (xray_neu).
+In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc, especially on the significant optical re-brightening that started four years after its initial flare (re-br).
 The paper is organized as follows.
-We explain the discovery of the transient event and give an insight about its host galaxy in Sect.
+We discuss the discovery of the transient event and give an insight about its host galaxy in Sect.
 2, summarize our observing campaign and the data reduction in Sect.
 3, present the results of the observations in Sect.
-4 and discuss them in Sect.
+4, discuss them in Sect.
 5 and finally give a summary and conclude in Sect.
@@
 Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and .
-In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al.
-2012).
+In this model (2006PASP..118.1711W), mas angular size corresponds to pc projected linear size at the source redshift (2012ApJS..203...21A).
 This redshift corresponds to a luminosity distance of Mpc.
-All magnitudes are given in AB system (Oke 1974).
+All magnitudes are given in AB system (1974ApJS...27...21O) unless stated otherwise.
 2 The transient event: AT2019aalc
-An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal i.e.
-the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Jiang et al.
-2016; Dou et al.
-2016; Lu et al.
-2016; van Velzen et al.
-2021b; Jiang et al.
-2021; van Velzen et al.
-2024).
-In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al.
-2021; Reusch et al.
-2022).
-This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al.
-2024).
-This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
-The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al.
-2013).
-Based on its Sloan Digital Sky Survey (SDSS: Abazajian et al.
-2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al.
-2013; Liu et al.
-2019).
-The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: Condon et al.
-1998).
+An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal, particularly an the infrared dust echo, stemming from reprocessed emission of optical/UV photons to the infrared by interaction with hot dust around the SMBH, at a distance of pc (2016ApJ...828L..14J; 2016ApJ...832..188D; 2016MNRAS.458..575L; 2021SSRv..217...63V; 2021ApJ...911...31J; Lancel).
+In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Bran; Tywin).
+This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (Lancel).
+This archival search revealed the nuclear transient AT2019aalc, with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (2019GCN.26258....1I), detected by the IceCube Observatory.
+The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (2013AJ....146..151O).
+Based on its Sloan Digital Sky Survey (SDSS: sdssdr7) spectrum it was further classified as a broad-line Seyfert-1 galaxy (2013ApJ...763..133T; 2019ApJS..243...21L).
+The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: 1998AJ....115.1693C).
 Fig.
-1 shows the Pan-STARRS (Chambers et al.
-2016) i-band optical image of the galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in .
+1 shows the Pan-STARRS (2016arXiv161205560C) -band optical image of the galaxy, together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by the Swift/XRT in .
 The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
-In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (van Velzen et al.
-2024).
+In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (Lancel).
 Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event.
-Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, van Velzen et al.
-2024) and low redshift.
-Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention.
+Notably, 2023ApJ...948...42W estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, Lancel) and low redshift.
+Since the proposed association with the high-energy neutrino event was announced roughly two years after the discovery of the transient event, the source was not at the focus of attention.
 Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event.
-Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al.
-2022; van Velzen et al.
-2024) suggesting a very unusual transient.
-Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al.
-2023), illustrated by the long-term ZTF difference light curve shown in Fig.
+Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Tywin; Lancel), suggesting a very unusual transient.
+Interestingly, roughly four years after the first flare, a significant optical re-brightening started in mid-May 2023 (re-br), illustrated by the long-term ZTF difference light curve shown in Fig.
 2.
@@
 3.1.1 ZTF
-ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al.
-(2023).
-We corrected for extinction by adopting a galactic extinction of (Schlafly & Finkbeiner 2011).
+ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in 2019PASP..131a8003M.
+We corrected for extinction by adopting a galactic extinction of (2011ApJ...737..103S).
 As shown in Fig.
@@
 This implies that during these epochs, the flux decreased below the mean flux of the reference images.
-For the -band these were obtained between March and May, 2018.
+For the -band, these were obtained between March 2018 and May 2018.
 While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
 3.1.2 UVOT
-The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks.
-The Ultraviolet and Optical Telescope (UVOT, Roming et al.
-2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ).
-We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package.
-We measured the flux using uvotsource, applying a circular aperature of arcseconds (which captures nearly all the flux of the host galaxy).
+The Neil Gehrels Swift Observatory (2004ApJ...611.1005G) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks.
+The Ultraviolet and Optical Telescope (UVOT, uvot) onboard the satellite observed the target field with the following four filters in each epoch: UVW2 (central wavelength, 1928 ), UVM2 (2246 ), UVW1 (2600 ) and U (3465 ).
+We analyzed the UVOT images using the Swift/UVOT tools included in the heasoft 6.30 software package.
+We measured the flux using uvotsource, applying a circular aperture of arcseconds (which captures nearly all the flux of the host galaxy).
 The background was measured using four 6 arcsecond regions in four quadrants away from the target source.
-To estimate the UV flux before the flare, we use GALEX (Martin et al.
-2005) observations obtained in 2007 and 2011.
-We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al.
-2016) to extract the flux.
-Over the 4 year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag.
-Following van Velzen et al.
-(2021a), we combine the mean GALEX flux with the SDSS (York et al.
-2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al.
-2009) that describes these data.
+To estimate the UV flux before the flare, we use GALEX (Martin05) observations obtained in 2007 and 2011.
+We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million16) to extract the flux.
+Over the four year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag.
+Following vanVelzen20, we combine the mean GALEX flux with the SDSS (york02) magnitudes and apply Prospector (Johnson17) to find the best-fit stellar population synthesis model (Conroy09) that describes these data.
 From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux.
@@
 3.2 Infrared
-The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al.
-2003) observed AT2019aalc on 2023-07-05.
+The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, 2003SPIE.4841..451W) observed AT2019aalc on 2023-07-05.
 The imaging was obtained in the J-, H,- and Ks-bands.
-The data was processed with the Swarp software (Bertin et al.
-2002).
+The data was processed with the Swarp software (2002ASPC..281..228B).
 We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels.
@@
 The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
-We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al.
-2010), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al.
-2011), at (W2).
-It visits any point in the sky roughly every six months for about ten times.
-Using timewise (Necker & Mechbal 2024; Necker et al.
-2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements.
+We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, 2010AJ....140.1868W), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (2011ApJ...731...53M), at (W2).
+NEOWISE visits any point in the sky roughly every six months, with each visit consisting of about ten exposures.
+Using timewise (neckerJannisNeTimewiseV02024; 2025A&A...695A.228N), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements.
 All selected single-exposure photometry data points are within arcsecond of the nucleus of the host.
@@
 3.3 X-ray
-The X-ray Telescope (XRT, Burrows et al.
-2005) observations were obtained simultaneously with the UVOT observations.
+The X-ray Telescope (XRT, 2005SSRv..120..165B) observations, in the energy range of keV, were obtained simultaneously with the UVOT observations.
 The observations were performed in photon counting mode.
-We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way.
-First, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks.
+We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way: first, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks.
 The ximage tool allowed us to determine the source detection level and its position.
@@
 We binned each spectrum to have a minimum of counts per spectral bin.
-We fit the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b).
-First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al.
-2016).
-The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976).
-To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023).
+We fit the individual X-ray spectra using the XSPEC fitting environment (swift).
+First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (2016A&A...594A.116H).
+The spectrum was then fitted using Cash-statistic (C-stat, 1976A&A....52..307C).
+To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model, as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (2023ATel16118....1P).
 When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary.
 The resultant fit quality is C-Stat for degrees of freedom.
-We than added a blackbody model component, to take into account a possible soft excess.
-The blackbody temperature and the normalization of the new component were also free to vary and we fit again which resulted in a better fit with C-Stat for degrees of freedom.
+We then added a blackbody model component, to take into account a possible soft excess.
+The blackbody temperature and the normalization of the new component were also free to vary.
+We fit again, which resulted in a better fit with C-Stat for degrees of freedom.
 We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component.
-After running simulations, the test resulted only in smaller likelihood differences than the observed one clearly implying that the difference is unlikely to be due to random fluctuations.
+After running simulations, the test resulted only in smaller likelihood differences than the observed one, clearly implying that the difference is unlikely to be due to random fluctuations.
 Therefore, accounting for the soft X-ray excess results in a significantly better fit.
@@
 3.4.1 VLBI observations
-We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al.
-2004).
+We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, 2004SPIE.5489..332G).
 These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres).
@@
 The data were recorded in left and right circular polarizations with a data rate of Gbit s-1.
-The correlation was done at the EVN Data Processor (Keimpema et al.
-2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s.
+The correlation was done at the EVN Data Processor (JIVEcorrelator) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s.
 The Compact Symmetric Object (CSO) 15210430 (a.k.a.
@@
 The data were calibrated with the U.S.
-National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources.
-We imaged the phase-calibrator with the Difmap software (Shepherd et al.
-1994), using the hybrid mapping technique (Högbom 1974).
+National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, aips) software package.
+The standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources were followed.
+We imaged the phase-calibrator with the Difmap software (difmap), using the hybrid mapping technique (hybridmapping).
 Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure.
@@
 At the end of the procedure, several steps of amplitude and phase self-calibration were performed.
-A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data.
+A simple elliptical Gaussian modelfit component (1995ASPC...82..267P) was fitted directly to the visibility data.
 Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
@@
 The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz.
-The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator.
+The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 was used as phase calibrator.
 The observations are summarized in Table LABEL:table:atca.
-The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: Sault et al.
-1995).
-The data reduction procedure was conducted in accordance with standard protocol.
+The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: miriad).
+The data reduction procedure was conducted in accordance with standard protocols.
 In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised.
@@
 Any corrupted data identified during calibration was manually flagged.
-Due to the limited ()-coverage and the low elevation of the source resulting in a very elongated beamshape, we were not able to perform self-calibration and cleaning.
-Fortunately, AT2019aalc was always dominating the flux within the resolution elements of our short observations.
+Due to the limited ()-coverage and the low elevation of the source, which resulted in a very elongated beamshape, we were not able to perform self-calibration and cleaning.
+Fortunately, AT2019aalc always dominated the flux within the resolution elements of our short observations.
 Therefore, we imaged the region around our source for each observation using an oversampling of the synthesised beam by a factor of at least 10.
 Under the assumption that AT2019aalc is always a point-source, we can take the highest pixel value in the image as our measured flux.
-Due to the uncertainties this method introduced we used a plain % flux error.
+Due to the uncertainties this method introduced, we used a plain % flux error.
 | Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz |
-28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
-30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
-08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 |
-19.07.2023 | 210 | 6D | 2.1 | 1 |
-09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 |
-19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 |
-09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 |
-29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 |
+| 28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
+| 30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
+| 08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 |
+| 19.07.2023 | 210 | 6D | 2.1 | 1 |
+| 09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 |
+| 19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 |
+| 09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 |
+| 29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 |
 3.5 Spectroscopic observations
-The Low Resolution Imaging Spectrograph (LRIS: Oke et al.
-1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni).
-Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal).
-The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al.
-2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al.
-2022) and PypeIt (Prochaska et al.
-2020) for the DBSP.
+We observed the spectrum of AT2019aalc on 2021-07-06 using the Low Resolution Imaging Spectrograph (LRIS: 1995PASP..107..375O) mounted on the Keck-I telescope at Maunakea (PI: Kulkarni).
+Later, during the second optical flare, we obtained five more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, 1982PASP...94..586O) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal).
+The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (2019PASP..131h4503P), the software package PypeIt (2020JOSS....5.2308P) for DeVeny, and the software packages DBSP DRT (2022JOSS....7.3612M) and PypeIt (2020JOSS....5.2308P) for the DBSP.
 Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig.
 17.
-To determine the type of AGN prior the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates (see Fig.
+To determine the type of AGN prior to the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: 2004PASP..116..138C; 2017MNRAS.466..798C; 2023MNRAS.526.3273C) and the newest GALAXEV (2003MNRAS.344.1000B) models (CB2019), assuming a 2001MNRAS.322..231K initial mass function (IMF) and an upper mass limit of , as templates (see Fig.
 18).
-Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al.
-(1989) extinction curve.
+Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by 2011ApJ...737..103S and the 1989ApJ...345..245C extinction curve.
 Furthermore, it was redshift corrected and logarithmically rebinned.
-We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to .
-Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al.
-(1989): a first one for stellar, and a second one for nebular emission.
-We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF.
+We included several emission lines in the fit and assumed four kinematic components: one for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to .
+Additionally, we included two separate dust components, each one with the extinction curve of 1989ApJ...345..245C: a first one for stellar, and a second one for nebular emission.
+We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets in pPXF.
 However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line.
@@
 In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
-The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (Guo et al.
-2018; Shen et al.
-2019; Ren et al.
-2024).
+The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (qso1; qso2; qso3).
 The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex.
@@
 AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days.
-The second flare started mid-May , approximately years after the initial flare (Veres et al.
-2023).
+The second flare started mid-May , approximately four years after the initial flare (re-br).
 The second flare evolved very similarly to the original one.
@@
 During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3).
-Both optical flares brightened with constant color () and showed reddening when decaying.
+Both main optical flares brightened with constant color () and showed reddening when decaying.
 Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys.
-To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: Shappee et al.
-2014; Kochanek et al.
-2017) and the Catalina Real-Time Transient Survey (CRTS: Djorgovski et al.
-2011).
-In addition, Palomar Transient Factory (PTF Law et al.
-2009) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu.
-No significant optical variability is visible between April and the discovery of the transient in January by ZTF (see Fig.
+To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: 2014ApJ...788...48S; 2017PASP..129j4502K) and the Catalina Real-Time Transient Survey (CRTS: 2011arXiv1102.5004D).
+In addition, Palomar Transient Factory (PTF 2009PASP..121.1395L) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu.
+No obvious signs of flaring activity is visible between April and the discovery of the transient in January by ZTF (see Fig.
 4).
 The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare.
-The U-band peak is coincident with the optical peak while the shorter wavelength bands peaked weeks later.
+The U-band peak is coincident with the optical peak, while the shorter wavelength bands peaked weeks later.
+Furthermore, the U- and UVW2-band peaks occur closer in time than the UVW1 and UVM2 peaks, suggesting a complex SED evolution near the maximum.
 We note that right after the -band peak, a short plateau phase started in optical and UV.
 The transient is most luminous at the shortest wavelength ranges covered by UVOT.
-The UV emission peaks around erg/s in the filter UW2.
-After the peak, the UW2U color shows reddening.
-Towards the end of the UVOT monitoring, the measured UM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
-Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (Bianchi et al.
-2011) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux.
+The UV emission peaks around erg/s in the filter UVW1.
+After the peak, the UVW2U color shows reddening.
+Towards the end of the UVOT monitoring, the measured UVM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
+Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (2011Ap&SS.335..161B) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux.
 The amplitude of this flare is higher than typically seen for AGN.
-Only out of the AGN presented in the GALEX Time Domain Survey (Gezari et al.
-2013) have larger amplitude NUV variability.
+Only out of the AGN presented in the GALEX Time Domain Survey (Gezari13) have larger amplitude NUV variability.
 4.2 Infrared
@@
 3.
-We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al.
-2003).
-The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
+We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (2003tmc..book.....C).
+The observed magnitudes in April are higher with mag in J-, and Ks-bands and mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
 The top panel of Fig.
 3 shows the light curve starting just before the first optical flare.
-The distinct first dust echo flare that peaks around six months after the first optical peak.
-The rise of the dust echo of the second flare becomes also visible in the latest WISE data releases.
-The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al.
-2004).
+The distinct first dust echo flare peaks around six months after the first optical peak.
+The rise of the dust echo of the second flare also becomes visible in the latest WISE data releases.
+The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (berkEnsemblePhotometricVariability2004).
 In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare.
 Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying.
-Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al.
-2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with .
-Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al.
-2012).
-However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al.
-2012), the color of AT2019aalc suggests a reliability of only %.
+Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (2025A&A...695A.228N), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with .
+Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (2012ApJ...753...30S).
+However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (2012ApJ...753...30S), the color of AT2019aalc suggests a reliability of only %.
 With the onset of the dust echo, the IR emission becomes dominated by the heated dust.
 The change from bluer to redder colors comes from cooling of the dust.
-The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue shown in Fig.5.
-Notably, AT2019aalc is not part of the Flaires sample (Necker et al.
-2024), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs.
+The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue, shown in Fig.
+5.
+Notably, AT2019aalc is not part of the Flaires sample (2025A&A...695A.228N), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs.
 This is because the broad application of the Flaires pipeline necessarily included a strict cut on extraneous variability that AT2019aalc did not pass.
 The authors did note however that a desirable relaxation of this criterion would include AT2019aalc.
-The source is, however, part of the dust echo sample of van Velzen et al.
-(2024) and, notably has the highest dust echo flux of all ZTF transients.
+The source is, however, part of the dust echo sample of Lancel and, notably has the highest dust echo flux of all ZTF transients.
 4.3 X-ray
-The X-ray luminosity evolution is shown in the bottom panel of Fig.
+The keV X-ray luminosity evolution is shown in the bottom panel of Fig.
 3 while the count rate variability is plotted in the top panel of Fig.
 6.
-The first flaring episode peaked at in the end of June (see Fig.
-3.) This flare peaked weeks before the optical peak.
+The first flaring episode peaked at a luminosity of in the end of June (see Fig.
+3.) This flare peaked two weeks before the optical peak.
 The count rate dropped by a factor of within a timescale of days.
-Later, the X-ray emission shows limited variability for several months.
-A hint of increased emission was detected in mid-February which was also followed by a bump in the optical and UV light curves around weeks later.
-In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode.
+Later, the X-ray emission showed limited variability for several months.
+A hint of increased emission was detected in mid-February , which was also followed by a bump in the optical and UV light curves around two weeks later.
+In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s.
+which is almost a magnitude larger than the peak of the first X-ray flaring episode.
 Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later.
@@
 6.
-AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al.
-2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence.
+AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (2021A&A...656A.132S) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence.
 The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV.
-The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al.
-2024).
-One observation with XMM-Newton Observatory (Jansen et al.
-2001) was performed on 2021-02-21 during a period of optical quiescence.
-The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023).
+The source had a soft thermal spectrum described with a blackbody temperature of eV (Lancel).
+One observation with XMM-Newton Observatory (2001A&A...365L...1J) was performed on 2021-02-21 during a period of optical quiescence.
+The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (2023ATel16118....1P).
 4.4 Radio
 4.4.1 Temporal evolution
-The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, White et al.
-1997) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability.
-These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz which is typical for Seyfert galaxies (Meurs & Wilson 1984).
+The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, 1997ApJ...475..479W) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability.
+These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz, which is typical for Seyfert galaxies (1984A&A...136..206M).
 Later, the transient’s sky location was observed in the Karl G.
-Jansky Very Large Array Sky Survey (VLASS, Lacy et al.
-2020).
+Jansky Very Large Array Sky Survey (VLASS, 2020PASP..132c5001L).
 The VLASS observes the entire sky visible to the VLA at GHz wide band three times.
 Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively.
-This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al.
-2024).
-In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: McConnell et al.
-2020).
+This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, Lancel).
+In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: racs1).
 The source was observed in two different fields and, consequently, at two different epochs.
-The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
+The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare observed in the VLASS data started at least days after the peak of the first optical flare, sometime between April and November .
 The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021.
-Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau.
+Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increase resulted in a long-lasting plateau.
 We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig.
@@
 We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities.
-The spectral index we derived between GHz and GHz this way is .
+The spectral index we derived this way is .
 Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz.
@@
 We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of .
-We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999).
-The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al.
-(2005).
-This yields mas implying that the fitted Gaussian component is resolved.
-We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Condon et al. 1982; Ulvestad et al. 2005):
+We calculated the uncertainties of the flux density and size parameters following the formulas of fomalont.
+The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following 2005AJ....130.2473K.
+This yields mas, implying that the fitted Gaussian component is resolved.
+We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Tb; Tb2):
 where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz.
 The obtained brightness temperature is K at GHz.
-The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015).
+The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., 2015A&A...574A..84L).
 4.5 Spectroscopic results
 We present the archival and the newly obtained optical spectra of AT2019aalc in Fig.
-17 with marking the notable emission lines.
-In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and five spectra that were taken during the second flare (over-flare-2 spectra).
+17, marking the notable emission lines.
+In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and the five spectra taken during the second flare (over-flare-2 spectra).
 4.5.1 Pre-flare spectrum
-A spectrum of the pre-flare system was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
+A spectrum of the pre-flare system was observed in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
 We identify strong Balmer emission lines.
 We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN.
-The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al.
-2019) for our source are consistent with this classification.
+The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (2019ApJS..243...21L) for our source are consistent with this classification.
 Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected.
 These lines are produced in the narrow-line region (NLR) in AGN.
-We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., Santoro et al.
-2020).
-The pre-flare spectrum also shows the detections of two high-ionization coronal lines ([Fe vii] and [Fe xiv]) which are further discussed below.
+We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., 2020A&A...644A..54S).
+The pre-flare spectrum also indicates the detections of two high-ionization coronal lines ([Fe x] and [Fe xiv]) which are further discussed below.
 4.5.2 Post-flare-1 spectrum
-The first LRIS spectrum was taken almost years after the first optical flare and years before the second one.
+The first LRIS spectrum was observed almost two years after the first optical flare and two years before the second one.
 In this spectrum, prominent Balmer lines were detected, similarly to the SDSS spectrum. The spectrum exhibits the aforementioned forbidden lines detected in the archival spectrum.
-We identify the Bowen Fluorescence line O iii and He ii .
+We identify the Bowen Fluorescence lines O iii and He ii .
 The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected.
-We identify the [Ne V] emission lines at and in the spectrum.
+Furthermore, we identify the [Ne V] emission lines at and in the spectrum.
 Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here.
-These lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks and are sometimes considered as coronal lines (e.g., Cann et al.
-2020).
+These neon lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks, and are sometimes considered as coronal lines (e.g., 2020ApJ...895..147C).
 4.5.3 Over-flare-2 spectra
-We performed optical spectroscopic observations of AT2019aalc during the second flare.
-These are characterized by a blue continuum with several emission lines appearing in comparison to the spectra discussed above.
-We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017) ,in Fig.
+We performed five optical spectroscopic observations of AT2019aalc during the second flare.
+These are characterized by a blue continuum with several more emission lines than the spectra discussed above.
+We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in 2017MNRAS.465...95P, in Fig.
 10.
-This comparison shows similarities but also significant differences discussed below.
-Over the second optical flare, strong Balmer lines were detected. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
-Over the second optical flare of AT2019aalc, we detected the Bowen Fluorescence line complex N iii +He ii .
-This complex was not detected in the earlier spectra.
+This comparison shows similarities but also remarkable differences discussed below.
+Over the second optical flare, strong Balmer lines were detected, with two of them (H and H) appearing only in the over-flare-2 spectra.
+Over the second optical flare of AT2019aalc, we identified the Bowen fluorescence line complex N iii +He ii .
+The complex was not detected in the earlier spectra.
 The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13).
 As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales.
-We also find hints of O iii which is also known as a Bowen line (Selvelli et al.
-2007), however, due to a possible blending from Fe vii, its detection is questionable.
-Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which supports the hint of detection.
-O iii was detected in some metal-rich TDEs before (Leloudas et al.
-2019) such as AT2019dsg (Cannizzaro et al.
-2021).
-We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum.
-Furthermore, a new coronal line appears: (Fe xi (ionization potential: eV).
+We also find hints of O iii, which is also known as a Bowen line (2007A&A...464..715S), however, due to a possible blending from Fe vii, its detection is questionable.
+Nevertheless, this Bowen line is created through the same channel (O1) as the above-discussed O iii , which supports the hint of detection.
+O iii was detected in some metal-rich TDEs before (2019ApJ...887..218L), such as AT2019dsg (2021MNRAS.504..792C).
+We detect the high-ionization coronal lines idenfitied in the post-flare-1 LRIS spectrum over the second flare in each spectrum.
+Furthermore, a new coronal line appears: [Fe xi] , which has an ionization potential of (e.g., 2022ApJ...936..140R).
 We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
@@
 In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum.
-The Balmer series lines here exhibit a redward wing which, however, in some cases, might result from blending with other lines, such as H and the BF line N,iii .
-Interestingly, the post-flare-I spectrum shows similar structure of the Balmer line series.
+The Balmer series lines here exhibit a redward wing which however, in some cases, might result from blending with other lines, such as H and the BF line N iii .
+Interestingly, the post-flare-I spectrum shows a similar structure to the Balmer line series.
 5.1.2 Bowen fluorescence lines
 Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation.
-The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (Netzer et al.
-1985; Netzer 1990).
+The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV), as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (1985ApJ...299..752N; 1990agn..conf...57N).
 This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc.
-The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al.
-2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al.
-1990).
+The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot) and AT2021loi and only in a few Seyfert galaxies before (e.g., 1986ApJ...310..679M; 1990ApJ...362...74S).
 However, we note that due to the atmospheric cutoff at these lines are difficult to detect.
@@
 10).
-The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al.
-2001), however, it has been observed in several optical TDEs (Charalampopoulos et al.
-2022b) and other SMBH-related transients: BFFs (Trakhtenbrot et al.
-2019; Makrygianni et al.
-2023), the peculiar transient event AT2019avd (Malyali et al.
-2021) and a few flaring NLSy1s (Frederick et al.
-2021).
-The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot et al.
-(2019).
+The BF line N iii is not generally seen in AGN (e.g., 2001AJ....122..549V), however, it has been observed in several optical TDEs (spectra2) and BFFs (Trakhtenbrot; AT2021loi).
+The detection of a BF line almost two years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot.
 In the case of AT2017bgt, this line was still detected days after the discovery of the transient.
-The Bowen features in F01004-2237 were detected even years after the initial flare (Tadhunter et al.
-2021).
+The Bowen features in F01004-2237 were detected even years after the initial flare (f01).
 We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig.
 11 around the N iii He ii complex.
-We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig.
-12 zooming into the O iii He ii region.
-The He ii ionized helium line is presented in the spectrum taken between the optical flares..
-This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g.
-Wang & Kron 2020).
-This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al.
-2019; Cannizzaro et al.
-2021).
+We also compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig.
+12, zooming into the O iii He ii region.
+The He ii emission line is commonly detected in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g.
+2020MNRAS.498.4550W).
+This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (1986Natur.322..511P; 2019A&A...622L..10S; 2021MNRAS.504..792C).
 In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism.
-Furthermore, we calculated a maximum line intensity ratio of which significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al.
-(2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ).
-Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot et al.
-2019).
+Furthermore, we calculated a maximum line intensity ratio of , which significantly exceeds the typical ratios estimated for AGN (earlier 2001AJ....122..549V estimated ratios of while 2012MNRAS.421.1043S gives ).
+Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot).
 5.1.3 Coronal lines
-Based on a study of SDSS spectra, Wang et al.
-(2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines.
-Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to exhibit coronal lines (Onori et al.
-2022; Newsome et al.
-2022; Short et al.
-2023).
-The BFF AT2021loi (Makrygianni et al.
-2023) and the spectacular transient event AT2019avd (Malyali et al.
-2021) also exhibit strong coronal lines.
-The BFF F01004-2237 showed variable coronal lines on a timescale of years.
-Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
-Wang et al.
-(2012) state that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later confirmed by Short et al.
-(2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN.
+Based on a study of SDSS spectra, 2012ApJ...749..115W reported the discovery of seven Extreme Coronal Line Emitter (ECLE) galaxies, which show extremely strong coronal lines.
+Numerous TDEs were also found to exhibit coronal lines.
+The BFF AT2021loi (AT2021loi) and the spectacular transient event AT2019avd (avd1) also exhibit strong coronal lines.
+The BFF F01004-2237 showed variable coronal lines on a timescale of years (f01).
+Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we detect enhanced coronal line emission in this BFF, too.
+The most plausible origin of the ECLEs is that they represent echoes of a past TDE outburst (2012ApJ...749..115W; 2024MNRAS.528.4775H; 2024MNRAS.532..112K; 2024ApJ...977..258N).
+Notably, 2023MNRAS.525.1568S revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN.
 The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario.
-Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al.
-2012).
-Later, Callow et al.
-(2024) discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations.
-They found that the ECLE rates are below the TDE rates, however, can still be consistent with assuming that only to per cent of all TDEs produce variable coronal lines.
-According to Wang et al.
-(2012), the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the TDE flare.
-A follow-up spectroscopic analysis of the original ECLE sample of Wang et al.
-(2012) revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (Clark et al.
-2024).
+Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (2012ApJ...749..115W).
+Later, ecle_rate_callow discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations.
+They found that the ECLE rates are below the TDE rates, however, can still be consistent with the assumption that only to percent of all TDEs produce variable coronal lines.
+According to 2012ApJ...749..115W, the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the echo of past TDE flares.
+A follow-up spectroscopic analysis of the original ECLE sample of 2012ApJ...749..115W revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (ecle_mw).
 The large dust echo of AT2019aalc is consistent with this picture.
-Furthermore, the TDEs with coronal line detection have more luminous IR dust echo compared to other optical TDEs.
-AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al.
-(2024) due to its strong dust echo.
-Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors.
+Furthermore, the TDEs with coronal line detection have more luminous IR dust echoes compared to other optical TDEs.
+AT2019qiz is one of the few TDEs included in the accretion flare sample of Lancel, due to its strong dust echo.
+Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of 2024MNRAS.531.2603H, which contains ANTs with high dust covering factors.
 The extreme dust echo seems to characterize the BFFs as well.
-The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al.
-2024).
-The Bowen line transient AT2019pev shows strong high-ionization coronal lines (Frederick et al.
-2021; Yu et al.
-2022) and is also part of the MIR flare sample of Hinkle (2024).
+The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (2024MNRAS.531.2603H) and two of them are also part of the Flaires sample of dust-echo-like IR flares (2025A&A...695A.228N).
+The BFF F01004-2237 exhibited a MIR flare (2017ApJ...841L...8D).
+The Bowen line transient AT2019pev shows strong high-ionization coronal lines (2021ApJ...920...56F; 2022MNRAS.515.5198Y) and is also part of the MIR flare sample of 2024MNRAS.531.2603H.
 These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs.
-We performed modeling of the publicly available LRIS spectrum of AT2017bgt taken on 2017-02-23 using PyQSOFit, following the procedure described in Subsect.
+Fig.
+13 shows the [Fe x] versus the [O iii luminosities of the ECLEs, the coronal-line detected TDEs and BFFs and AT2019aalc, together with the coronal-line detected SDSS Seyfert sample of 2009MNRAS.397..172G.
+For this plot, we performed modeling of the publicly available spectra888We retrieved the LRIS spectra from the Transient Name Server (TNS; https://www.wis-tns.org) and the ESO Faint Object Spectrograph and Camera version 2 (EFOSC2) spectra from the ESO Science Portal (archive.eso.org/scienceportal) of several TDEs and BFFs using PyQSOFit, following the procedure described in Subsect.
 3.5.
+These are the TDEs: AT2017gge (2017-09-14, EFOSC2); AT2018dyk (2018-08-08, LRIS); AT2018bcb (2018-05-06, EFOSC2); AT2021qth (2022-05-26, LRIS); AT2022upj (2022-10-26, DBSP); and the BFFs: AT2017bgt (2017-02-23, LRIS) and F01004-2237 (2018-08-13, EFOSC2).
+Absolute fluxes, and hence derived luminosities, are subject to systematic uncertainties from slit losses and atmospheric extinction differences.
+We adopt a conservative % uncertainty on the flux scale to account for these effects, which is combined in quadrature with the statistical uncertainties from the PyQSOFit spectral fits.
+This method is applied only to cases where the slit position angle differed at least five degrees from the parallactic angle.
 Fig.
-13 shows the [Fe x] versus the [O iii luminosities of the ECLEs studied by Wang et al.
-(2012), AT2019qiz (Short et al.
-2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al.
-(2009).
-This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
+13 clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
+We note that the only object in the sample of 2009MNRAS.397..172G classified as a galaxy rather than an AGN is, in fact, the same source as one of the variable ECLEs (which appear as nearly overlapping points in our plot).
+No publicly available spectra were found of the TDEs AT2021dms, AT2024mvz and AT2021acak while in the case of AT2017gge, it was not possible to safely distinguish [Fe x] from H.
+We note that AT2021acak exhibited two distinct optical flares (2023RAA....23b5012L), with the second being more luminous, and showed a He ii N iii complex similar to that of AT2019aalc.
 5.2 Optical re-brightening
 The long-term ZTF light curve of AT2019aalc shows two well-separated flares.
-Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales.
-Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption.
-AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began years after the initial flare.
-The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al.
-2023).
+Both flares rose quickly and reached their peaks within two months, followed by a slow decay on yearly timescales.
+Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (1988Natur.333..523R), and without any prominent bumps in the case of full disruption.
+AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began four years after the initial flare.
+The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (2020vdq).
 This implies an even more significant optical re-brightening than the one detected in AT2019aalc.
+In contrast, the candidate pTDE AT2022dbl exhibits with a brighter first flare.
 The second flare of partial TDEs is expected to follow a steep declining with a power-law of .
-In the case of AT2020vdq, the second flare declined much slower (Somalwar et al.
-2023) than this but still clearly faster than the second flare of AT2019aalc which declined with a power-law index of .
+In the case of AT2020vdq and AT2022dbl, the second flares declined slower than this (2020vdq; 2025ApJ...987L..20M), but still faster than the second flare of AT2019aalc, which declined with a power-law index of .
+Notably, AT2020vdq and AT2022dbl both exhibit slower-declining second flares compared to their first flares, similar to AT2019aalc.
 Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps.
 Fig.
-14 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc.
-The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot et al.
-(2019).
-Originally, F01004-2237 had undergone a luminous flare (Tadhunter et al.
-2021) in 2010 and later Trakhtenbrot et al.
-(2019) classified it as a BFF.
-The binned light curve of F01004-2237 published in Makrygianni et al.
-(2023) shows a bump around a year after its peak.
+14 shows the optical evolution of the three BFFs with detected re-brightening episodes, together with AT2019aalc.
+The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot.
+Originally, F01004-2237 had undergone a luminous flare (f01) in 2010 and later Trakhtenbrot classified it as a BFF.
+The binned light curve of F01004-2237 published in AT2021loi shows a bump around a year after its peak.
 We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring.
@@
 In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak.
-Makrygianni et al.
-(2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs.
-Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date.
-The first flare of the studied BFFs decay on yearly timescales, much slower than regular TDE flares but more similar to AT2019aalc.
-However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower.
-The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities we see.
-These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., Charalampopoulos et al.
-2023).
+AT2021loi suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs.
+OGLE17aaj, has not shown a clear optical re-brightening, however a bump likely appears roughly a year after the peak, similar to AT2017bgt and AT2021loi.
+The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities.
+These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., 2023A&A...673A..95C).
 5.3 Radio characteristics
@@
 In this subsection we discuss the long-term radio variability of AT2019aalc.
-The source experienced a brightening of its flux density by a factor of on a timescale of less than years followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring.
-Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., Mundell et al.
-2009; Koay et al.
-2016) and therefore require further investigation.
+The radio source experienced a brightening of its flux density by a factor of two on a timescale of less than years, followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring.
+Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., seyf_radio_var; 2016MNRAS.460..304K) and therefore require further investigation.
 The unusual radio spectrum with a high frequency turnover is also discussed below.
-The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (Condon 1992).
-Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see Bontempi et al.
-2012 as an example how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies).
-This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, Gallimore et al.
-1997) as the origin of the detected radio emission.
-Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997).
-Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
+The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (sfrTb).
+Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see 2012MNRAS.426..588B as an example of how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies).
+This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, 1997Natur.388..852G) as the origin of the detected radio emission.
+Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, 1997ApJ...477..585M).
+Assuming that accretion flows dominate the radio emission of AT2019aalc, following 1998ApJ...499..198Y we estimate an expected GHz radio luminosity of erg s-1, orders of magnitude lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
 Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc.
-Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG: Green et al.
-1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993).
-Later, Behar et al.
-(2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies.
-For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc.
+2008MNRAS.390..847L found a correlation of studying the Palomar–Green (PG: 1986ApJS...61..305G) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by 1993ApJ...405L..63G).
+Later, 2015MNRAS.451..517B came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies.
+For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of , suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc.
 Moreover, coronal emission should be highly variable at radio wavelengths.
-Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., Paragi et al.
-2017; Mohan et al.
-2022; Kozák et al.
-2024).
+Out-flowing synchrotron-emitting material induced by enhanced accretion is commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., 2017IAUS..324..119P; 2022ApJ...927...74M; 2024Galax..12....8K).
 The high brightness temperature of the EVN-detected radio feature indicates its radio core identification.
-The steep () radio spectrum indicate the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum.
-The spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s).
+The steep () radio spectrum indicates the presence of another, steep-spectrum synchrotron-emitting component(s) in the system, as radio cores are characterized by a flat radio spectrum.
+The spectrum may be the composition of the flat-spectrum radio core and the additional, steep-spectrum component(s).
 The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
 5.3.2 The radio spectrum
-The turnover suggests the multiple component nature of the spectrum.
-Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al.
-2020).
+The turnover suggests the multi-component nature of the spectrum.
+Radio spectra with high-frequency excesses have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (2020A&A...636A..64B).
 These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN.
 The high-frequency peak might be explained with the kinematic age of a newly ejected component.
-These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al.
-2020).
-The low-frequency peak is, however, more likely to be explained with star formation activity instead of AGN emission since the latter is absorbed due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
-However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA.
+These extremely young radio sources therefore peak at very high frequencies (e.g., 2020A&A...636A..64B).
+The low-frequency peak is more likely originating from star formation activity instead of AGN emission, since the latter is absorbed at these frequencies due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
+However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at lower frequencies.
 Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
-A newly ejected component can explain the high-frequency excess whilst the low-frequency part is originating from optically thin synchrotron emission of the AGN.
+A newly ejected component can explain the high-frequency excess, whilst the low-frequency part originates from optically thin synchrotron emission of the AGN.
 Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior.
-Järvelä et al.
-(2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al.
-(2020).
-Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al.
-2009).
+2021FrASS...8..147J states that restarted AGN activity can explain at least some of the radio spectra studied by 2020A&A...636A..64B.
+Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (2009ApJ...698..840C).
 5.4 X-ray properties
 The X-ray spectra of AGN at energies above keV have a power-law-like shape.
-These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al.
-2015).
-Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess.
-It can be well modeled by a blackbody model with a best-fit temperature in the range keV.
-The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al.
-2012).
+These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., 2015A&A...582A..40K).
+Below keV some AGN show an excess in their spectrum referred to as the soft X-ray excess.
+It can be modeled well by a blackbody model with a best-fit temperature in the range keV.
+The origin of the excess might be explained via Comptonized disk emission.
+The reflection of hard X-ray photons from the surface of the disk can also account for the excess (2012MNRAS.420.1848D).
 The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV.
-Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., Pons & Watson 2014).
-Our value is more commonly seen in sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al.
-1996; Saxton et al.
-2020).
-The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x].
-The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al.
-2009).
+Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., 2014A&A...568A.108P).
+AT2019aalc exhibits properties that are more common in sources characterized by enhanced accretion, e.g., NLSy1s or TDEs which both have typically soft X-ray spectra with (1996A&A...305...53B; 2020SSRv..216...85S).
+The soft X-ray flux of Seyfert galaxies was found to be correlated well with the observed fluxes in the coronal line [Fe x] .
+The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (2009MNRAS.397..172G).
 For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day).
-Thus, the soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as suggested for ECLEs by Wang et al.
-(2012).
-The reoccurring optical/UV bumps seen during the decaying of the main optical flares are likely driven by the soft X-ray flares which lead these outbursts only with a couple of weeks.
+This suggests that this coronal line is excited by the X-ray flares discussed here instead of past X-ray flares that are thought to power the ECLEs.
+The reoccurring optical/UV bumps appearing during the decay of the main optical flares are likely driven by the soft X-ray flares, which lead these outbursts only by a couple of weeks.
 The soft X-ray excess appears to be more dominant during these flares.
-Interestingly, the soft X-ray flares show hints of quasi periodicity of days.
-A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit with soft X-ray flares on time-scales of hours to days while partial TDEs on timescales of months to years (e.g., Chakraborty et al.
-2021; Quintin et al.
-2023; Miniutti et al.
-2023; Nicholl et al.
-2024).
+Interestingly, the soft X-ray flares show hints of quasi-periodicity of days.
+A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit soft X-ray flares on time-scales of hours to days, while partial TDEs exhibit such flares on timescales of months to years (e.g., qpe2; qpe1; qpe3; qizqpe).
 We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring.
-This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: Evans et al.
-2023) as LSXPS J031335.6020905.
+This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: 2023MNRAS.518..174E) as LSXPS J031335.6020905.
 The potential connection between BFFs and QPEs should be explored in future studies.
 5.5 SED fitting
-Following Reusch et al.
-(2022), we fitted two blackbody models in order to describe the SED of the transient using lmfit.
-These fitted curves to the optical/UV (’blue’ blackbody) and IR datapoints (’red’ blackbody) are shown in Fig.
+Following Tywin, we fitted two blackbody models in order to describe the SED of the transient using lmfit.
+These curves fitted to the optical/UV (blue blackbody) and IR datapoints (red blackbody) are shown in Fig.
 15, together with a combined double blackbody fit.
-The red blackbody fit was introduced to account for the IR dust echo.
+The red blackbody fit was introduced in order to account for the IR dust echo.
 These curves fit the SED around the optical peak of the investigated flare.
-The continuum emission of TDEs is well described by a thermal blackbody model (Gezari 2021).
+The continuum emission of TDEs is described well by a thermal blackbody model (2021ARA&A..59...21G).
 The case of AT2019aalc appears to be more complex.
-The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization explaining the poor fit towards the blue.
-The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations as the IR light curve is peaked well after the optical one.
-The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range but such a significant excess cannot be fully explained with the coronal lines only.
-The inferred blackbody temperature and radius values estimated for the ’blue’ blackbody fit are K and cm, respectively.
-The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by van Velzen et al.
-(2020) and Thomsen et al.
-(2022).
-On the other hand, the radius is smaller than calculated for optical selected TDEs but is more consistent with the X-ray detected ones.
-The ’red’ blackbody fit resulted in K and cm.
+The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization, explaining the poor fit towards the blue.
+The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations, as the IR light curve is peaked well after the optical one.
+The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range, but such a significant excess cannot be fully explained with the coronal lines only.
+Instead, a variable H component, given its much larger equivalent width, or reprocessing, is more likely to significantly contribute to the observed excess and warrants consideration.
+The inferred blackbody temperature and radius values estimated for the blue blackbody fit are K and cm, respectively.
+The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by 2020SSRv..216..124V and 2022ApJ...937L..28T.
+On the other hand, the radius is smaller than those calculated for optically selected TDEs, however is more consistent with the X-ray detected ones.
+The red blackbody fit resulted in K and cm.
 5.6 What type of transient is AT2019aalc?
-AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color.
-The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021).
-Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case.
+AT2019aalc shares similarities with TDEs including the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color.
+The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (2021ARA&A..59...21G).
+Bowen lines appear in the spectra of several TDEs even if these are typically weaker than in our case.
 The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far.
 The reddening in color when the main optical flares are declining is also atypical for TDEs.
-The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc.
+However, in a more complex AGN environment, compared to the simpler case of standard TDEs in quiescent galaxies, reprocessing may already be significant at early times, implying that the observed reddening does not necessarily reflect genuine cooling.
+The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, optical emission decaying on yearly timescales, and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc.
 Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
@@
 TDEs in AGN might play a role.
-In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al.
-2017, and ASASSN-14ko, Huang et al.
-2023) or the AGN’s radiation field.
-According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (Chan et al.
-2019).
-The debris-disk interaction results in enhanced accretion and multiple shocks (Chan et al.
-2019), notably leading to X-ray emission (Ryu et al.
-2024).
+In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, 2017ApJ...843..106B, and ASASSN-14ko, 2023ApJ...956L..46H) or the AGN’s radiation field.
+According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (tdes_in_agn).
+The debris-disk interaction results in enhanced accretion and multiple shocks (tdes_in_agn), notably leading to X-ray emission (2024MNRAS.527.8103R).
 Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk.
 These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare.
-The soft X-ray emission is likely reprocessed in optical/UV by the dense environment explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines.
+The soft X-ray emission is likely reprocessed to optical/UV by the dense environment, explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines.
 The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares.
-TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (Chan et al.
-2019; Ryu et al.
-2024).
-Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (Somalwar et al.
-2023) and the BFF F01004-2237.
-Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (Zhong 2024; Liu et al.
-2024).
-Liu et al.
-(2024) claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq.
-A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2017bgt and AT2021loi) with double-peaked optical flares as well.
-Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well.
-The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understand their unique optical light curves.
-Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs.
-This would explain the TDE-like properties of these sources.
+TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (tdes_in_agn; 2024MNRAS.527.8103R).
+Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (2020vdq) and the BFF F01004-2237.
+Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (vdq2; partial_models).
+vdq2 claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq.
+A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2021loi and F01004-223) with double-peaked optical flares as well.
+A TDE-in-AGN scenario would explain the TDE-like properties of these sources.
 The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”).
-TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (Chan et al.
-2019).
-Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al.
-2012) and the BFFs (see Fig.13) is consistent with this picture.
+TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (tdes_in_agn).
+Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, 2012ApJ...749..115W) and the BFFs (see Fig.
+13) is consistent with this picture.
 A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599.
-This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al.
-2015).
+This low-luminosity AGN exhibits periodic soft X-ray flares interpreted as being due to interaction with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (2015A&A...581A..17C).
 Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig.
 13).
-In addition, IC 3599 is the only known AGN with fading coronal lines (Frederick et al.
-2019).
+In addition, IC 3599 is the only known AGN with fading coronal lines (2019ApJ...883...31F) a behavior also observed in the TDE AT2019qiz (2023MNRAS.525.1568S), the BFF F01004-2237 (f01) and most of the ECLEs (2013ApJ...774...46Y).
+Another feature AT2019aalc shares with other candidate pTDEs is the Heii + Niii complex, observed in the optical spectra of all five optically selected pTDE candidates (see Fig.
+13.
+in 2024arXiv241215326H).
+Interestingly, other BFFs also showed weaker pre-transient AGN activity than persistently accreting AGN (Trakhtenbrot; AT2021loi).
+It suggests that TDEs might have reactivated these AGN.
+However, between the two main optical flares, AT2019aalc exhibited negative fluxes, i.e., it was fainter than the baseline (see Fig.
+2).
+This behavior suggests a retrograde orbit of the disrupted star relative to the accretion disk, as such encounters gain angular momentum and lead to a subsequent lower AGN state (2022MNRAS.514.4102M).
+In contrast, a prograde orbit adds angular momentum to the disk, resulting in a higher AGN state, which may explain the light curve of OGLE17aaj, where the source persisted in a high and stable state for years after the peak (2019A&A...622L...2G), unlike AT2019aalc.
 5.7 Radio properties of BFFs
 As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs.
-Based on the radio observations explained in Makrygianni et al.
-(2023), AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz.
+Based on the radio observations explained in AT2021loi, AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz.
 Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability.
@@
 The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy.
-The BFF AT2020afhd was detected with the VLA at GHz with mJy (Christy et al.
-2024).
-If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations published in Chan et al.
-(2019).
+F01004-2237 was detected with VLBI at Ghz, however, remained undetected at GHz which indicates an inverted spectrum (2024ApJ...970....5H).
+The BFF AT2020afhd was detected with the VLA at GHz with mJy (2024TNSAN..56....1C).
+If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations (tdes_in_agn).
 Radio observations of any newly discovered BFF will be essential to better understand these transients.
 5.8 BFFs and high-energy neutrinos
-AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al.
-2024).
+AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (2019GCN.26258....1I; Lancel).
 Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect.
-5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: Abbasi et al.
-2023).
-We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023).
+5.6), we searched for spatial coincidences between the earlier-classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: 2023ApJS..269...25A).
+We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (2023GCN.33773....1I).
 In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour.
 We present the source positions together with the neutrino best-fit positions and both contours in Fig 16.
-We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A.
+We also report a coincidence between the prototype BFF AT2017bgt and the high-energy neutrino event IC-200410A.
 This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
@@
 Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks.
-This might be explained with delayed outflow activity (van Velzen et al.
-2024).
-In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al.
-2023) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc.
-Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al.
-2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients.
+This might be explained with delayed outflow activity (Lancel).
+In the case of AT2021loi, the inverted radio spectrum (AT2021loi) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc.
+Interactions between the out-flowing material and clouds in the surrounding of the SMBH (2022MNRAS.514.4406W) or between ultra-high-energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (2023ApJ...948...42W) were proposed to explain the observed neutrino emission of the three nuclear transients.
 These findings indicate that outflows play an important role in the neutrino production.
-A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
-If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources.
-Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al.
-2024; Hinkle 2024: see also Subsect.
-4.2) and unusually strong high-ionization coronal lines (Makrygianni et al.
-2023: see also Subsect.
+A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.echos
+If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echoes and therefore more candidate neutrino sources.
+Notably, both AT2019aalc and AT2021loi are characterized by significant dust echoes (Lancel; 2024MNRAS.531.2603H: see also Subsect.
+4.2) and unusually strong high-ionization coronal lines (AT2021loi: see also Subsect.
 5.1).
 6 Conclusions
-We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc associated with a high-energy neutrino event.
-We found that the second optical flare is more luminous and decays slower than the first one.
-Both flares evolve rapidly and with constant color.
-The UV emission peaks at extremely high luminosities (at erg/s).
-The re-brightening is accompanied by an IR flare, similarly to the first flare when it was explained with dust echo emission.
+We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc, which have been associated with a high-energy neutrino event.
+We found that the second optical flare was more luminous and decayed slower than the first one.
+Both flares rose rapidly and with constant color.
+The UV emission peaked at extremely high luminosities (at erg/s).
+The re-brightening was accompanied by an IR flare, similarly to the first flare, when it was explained with a dust echo emission.
 An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature.
-The X-ray flares are very soft which explains the unusually strong He ii lines.
+The X-ray flares are very soft, which explains the unusually strong He ii lines.
 In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare.
-With VLBI we clearly detect a non-thermal radio source which favors an AGN-like origin.
+With VLBI, we clearly detect a non-thermal radio source, which favors an AGN-like origin.
 The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component.
@@
 The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment.
-The stellar debris-disk interaction in such a system could result in complex light curve variations e.g., the plateau observed in optical/UV (Chan et al.
-2020).
-A TDE-related scenario provides an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the disk gas which was notably suggested by TDE-AGN simulations (Ryu et al.
-2024).
+The stellar debris-disk interaction in such a system could result in complex light curve variations, e.g., the plateau observed in optical/UV (tdes_in_agn_2).
+A TDE-related scenario provides an explanation for the suddenly enhanced activity of the nucleus.
 Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
-Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs.
-We found that AT2021loi is coincident with the high-energy neutrino event IC230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al.
-(2024).
-Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode in optical.
-In both cases the AGN are classified as broad-line AGN in-prior the transients occurred.
+Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the four known BFFs.
+We found that AT2021loi is coincident with the high-energy neutrino event IC230511, confirming the flaring AGN-neutrino connection studied by Lancel.
+Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and a re-brightening episode in optical.
+In both cases the AGN are classified as broad-line AGN prior to when the transients occurred.
 Acknowledgements.
+We sincerely thank the anonymous referees for their detailed comments and insightful suggestions, which have greatly enhanced the clarity and quality of this paper.
 We thank Benny Trakhtenbrot for useful discussions.
+We thank Elisa Pueschel for proofreading.
 PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal.
-EH acknowledges support by NASA under award number 80GSFC21M0002.The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes.
+EH acknowledges support by NASA under award number 80GSFC21M0002.
+We thank the Swift team for approving and conducting our ToO requests.The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes.
 Scientific results from data presented in this publication are derived from the following EVN project code: EV027.